### v3

# Entropy and determinants for unitary representations

Abstract
Several notions of entropy have been studied widely in the ergodic theory of probability-preserving actions of countable groups. These include the generalization of Kolmogorov–Sinai entropy to amenable groups based on Følner sequences, percolative entropy for general countable groups, and Bowen’s sofic entropy for sofic groups.
In this work we pursue these notions across an analogy between ergodic theory and representation theory. We arrive at new quantities associated to unitary representations of groups and representations of other C*-algebras. Our main results show that these new quantities can often be evaluated as Fuglede–Kadison determinants. The resulting determinantal formulas offer various non-commutative generalizations of Szegő’s limit theorem for Toeplitz determinants. They make contact with Arveson’s theory of subdiagonal subalgebras, and also with some exact entropy calculations in the ergodic theory of actions by automorphisms of compact Abelian groups.
[…]
0.4pt0cm
1 Introduction
1.1 A Szegő limit theorem over amenable groups
If $\varphi$ is a positive definite function on $\mathbf{Z}$, then Bochner’s theorem identifies it as the Fourier–Stieltjes transform of a finite Borel measure $\mu$ on the circle group $\mathbf{T}$. Let $m_{\mathbf{T}}$ be the Lebesgue probability measure on $\mathbf{T}$. In this context, Szegő’s limit theorem describes the asymptotic behaviour of finite-dimensional Toeplitz determinants obtained from $\varphi$:
Theorem S.
Let $D_{n}$ be the determinant of the Toeplitz matrix $[\varphi(i-j)]_{i,j=1}^{n}$ for each $n$, and let $\mu_{\mathrm{ac}}$ be the absolutely continuous part of $\mu$. Then
$D_{n}^{1/n}\to\exp\int_{\mathbf{T}}\log\frac{d\mu_{\mathrm{ac}}}{dm_{\mathbf{T}}}\,dm_{\mathbf{T}}\qquad\hbox{as}\ n\to\infty,$ (1.1)
taking the right-hand side to be $\exp(-\infty)=0$ if necessary. ∎
A thorough account of Theorem S can be found in standard texts such as [136, 137], which I generally follow for terminology in the sequel. See [136, Chapter 1] for a broader overview of this area and a history for some of the key contributions. In fact, Szegő’s early works contain (1.1) only when $\mu\ll m$, with the extension to general measures $\mu$ accomplished first by Verblunsky. Theorem S also has a generalization in which $\mu$ and $\varphi$ take values among positive definite $k$-by-$k$ matrices [136, Section 2.13].
The right-hand side of (1.1) admits the following ‘non-commutative’ point of view. Let $f:=d\mu_{\mathrm{ac}}/dm_{\mathbf{T}}$, so this is an element of $L^{1}(\mathbf{T})$. We can regard $L^{\infty}(\mathbf{T})$ as a von Neumann algebra of multiplication operators on $L^{2}(\mathbf{T})$, and then integration with respect to $m_{\mathbf{T}}$ defines a tracial state on $L^{\infty}(\mathbf{T})$. If $f$ is essentially bounded, then the multiplication operator $M_{f}$ belongs to this von Neumann algebra; in general, $M_{f}$ can be defined as an operator affiliated to that algebra. Then the right-hand side of (1.1) is the logarithm of the Fuglede–Kadison determinant of $M_{f}$ with respect to $m_{\mathbf{T}}$. We review these concepts more carefully in Section 3 below. In the context of Szegő’s theorem, this observation goes back at least to [9], where Arveson proposed a non-commutative generalization to ‘subdiagonal subalgebras’, which are discussed further below.
Under various additional hypotheses, generalizations of Theorem S to positive definite functions on $\mathbf{Z}^{d}$ were obtained in [85, 105, 51]. More recently, further generalizations appeared in works by Deninger [43, 44, 45], Li [103] and Li and Thom [104]. They handled more general amenable groups, but still with additional hypotheses on the positive definite functions.
For a countable group $\Gamma$, one of the many equivalent conditions for amenability is the existence of a sequence $(F_{n})_{n\geq 1}$ of finite subsets satisfying the right Følner condition:
$|F_{n}\triangle F_{n}g|=o(|F_{n}|)\qquad\hbox{as $n\to\infty$ for every $g\in\Gamma$}.$ (1.2)
This condition is discussed in [94, Section II.5], or the analogous left Følner condition is explained in [75, Section 3.6]. The inversion map on $\Gamma$ converts one condition to the other, so both are characterizations of amenability. We refer to a sequence of subsets that satisfies (1.2) as a right Følner sequence.
Our first main result is a version of Theorem S for arbitrary matrix-valued positive definite maps on a countable amenable group $\Gamma$ (see Section 5 for the definition of these maps). Let $\varphi:\Gamma\to\mathrm{M}_{k}$ be positive definite, and for each finite subset $F$ of $\Gamma$ consider the $F$-by-$F$ block matrix
$\varphi[F]=[\varphi(g^{-1}h):\ g,h\in F].$
If $\varphi$ is associated to a unitary representation $\pi$ by the vectors $v_{1}$, $v_{2}$, …, $v_{k}$, then $\varphi[F]$ is the Gram matrix of the tuple of vectors
$[\pi(g)v_{i}:\ g\in F,\ i=1,\dots,k],$
so it is positive semidefinite.
Let $\lambda$ be the left regular representation. Classic results from functional analysis give a unique decomposition of $\varphi$ as $\varphi_{\mathrm{sing}}+\varphi_{\mathrm{ac}}$, where the minimal dilation of $\varphi_{\mathrm{sing}}$ is disjoint from $\lambda$ and the minimal dilation of $\varphi_{\mathrm{ac}}$ is a subrepresentation of $\lambda$ (see Propositions 3.3 and 4.9 below). In addition, $\varphi_{\mathrm{ac}}$ can be represented in terms of a self-adjoint operator affiliated to $\lambda^{\oplus k}(\Gamma)^{\prime}$ (see Propositions 3.5 and 4.10), and using this we can define the Fuglede–Kadison determinant $\Delta\varphi_{\mathrm{ac}}$ (see Definition 4.12).
Theorem A.
Any right Følner sequence $(F_{n})_{n\geq 1}$ satisfies
$(\det\varphi[F_{n}])^{1/|F_{n}|}\to\Delta\varphi_{\mathrm{ac}}\qquad\hbox{as}\ n\to\infty.$ (1.3)
Notice that the contribution of the singular part $\varphi_{\mathrm{sing}}$ to the left-hand side disappears in the limit, just as in Szegő’s original theorem.
The literature contains a number of precedents lying between Szegő’s theorem and Theorem A. For example, [43, Theorem 3.2] implies the case of Theorem A in which $\varphi_{\mathrm{sing}}=0$ and $\varphi$ can be expressed in terms of a positive invertible element of the von Neumann algebra $\mathrm{M}_{k}(\lambda(\Gamma)^{\prime\prime})$. We discuss the previous literature more fully in Subsection 6.3.
1.2 Analogies with entropy and ergodic theory
For a $k$-tuple of vectors in a Hilbert space, their Gram matrix specifies their lengths and relative positions. In doing so, it is roughly analogous to the joint distribution of $k$ discrete random variables. If $Q$ is this Gram matrix, then the analog of the joint Shannon entropy of those random variables is the quantity $\log\det Q$.
More formally, log-determinants arise as values of differential rather than discrete entropy. If $(X_{1},\dots,X_{k})$ is a multivariate Gaussian random vector with covariance matrix $Q$, then its differential Shannon entropy equals $\log\det Q$ up to a normalization [40, Theorem 8.4.1]. From this point of view, the analogy sketched above is really between discrete and differential entropy. However, the present paper concerns a purely ‘linear’ setting, so we generally discuss this analogy without the extraneous construct of a Gaussian random vector. Either way, a few key differences must be kept in mind when comparing with discrete entropies, particularly these: (i) log-determinants (differential entropies) depend on the ambient dimension in which a (random) vector takes it values; (ii) log-determinants (differential entropies) can have either sign, or even equal $\infty$- ∞.
Now let $\Gamma$ be a countable group, and consider actions of two different kinds: measure-preserving actions on probability spaces, and unitary actions on Hilbert spaces. These two settings enjoy a $\Gamma$-equivariant version of the analogy between joint distributions and Gram matrices. Observables (equivalently, finite measurable partitions) on a probability space are analogous to vectors in a Hilbert space, or more generally to finite tuples of vectors. If $(X,\mu,T)$ is a measure-preserving $\Gamma$-system and $\alpha:X\to A$ is an observable, then it generates the shift-system $(A^{\Gamma},\alpha^{\Gamma}_{\ast}\mu,S)$, where $\alpha^{\Gamma}_{\ast}\mu$ is the law of the $\Gamma$-indexed stochastic process $(\alpha\circ T^{g}:\ g\in\Gamma)$. Analogously to this, if $\pi$ is a unitary representation of $\Gamma$ and $v\in H_{\pi}$, then these define the positive definite function $\langle\pi(\cdot)v,v\rangle$ on $\Gamma$, and this construction can be extended to finite tuples of vectors by considering matrix-valued positive definite functions.
This analogy has long stimulated research in both areas. For example, it underlies Kechris’ adaptation of the relation of weak containment to measure-preserving systems [93, 31]. Within this analogy, the logarithm of the limit on the left-hand side in Theorem A is the analog of the Kolmogorov–Sinai entropy of a stationary process over an amenable group [96].
The analogy between ergodic theory and representation theory crystallizes into at least two different formal relationships. On the one hand, any measure-preserving system gives rise to its Koopman representation [93, Section 10]. On the other hand, any orthogonal real representation of $\Gamma$ can be used to construct a measure-preserving action on a Gaussian Hilbert space, and this construction can be adapted to start with a unitary representation instead [93, Appendices C–E]. However, the first of these relationships does not correctly connect the notions of entropy that we study in this work, and the second introduces unnecessary complications.
On the one hand, the limit in (1.3) generally bears no relation to Kolmogorov–Sinai entropy when $\varphi$ is associated to a Koopman representation; indeed, the quantity in (1.3) does not even define an invariant of unitary equivalence in general.
On the other hand, if we start with a unitary representation and construct the associated Gaussian system, then there are cases in which a suitable limit of log-determinants for the former should equal a ‘differential’ analog of Kolmogorov–Sinai entropy for the latter. For single transformations or actions of $\mathbf{Z}^{d}$, differential Kolmogorov–Sinai entropy for stationary $\mathbf{R}$-valued processes was studied in [78], prompted by earlier work in information theory such as [119]. Section 6 of [78] includes some exact calculations for Gaussian systems and linear transformations between them which boil down to applications of Theorem S itself. However, this differential version of Kolmogorov–Sinai entropy starts to behave quite wildly beyond those Gaussian examples, and for those examples alone we might as well stay within the setting of representation theory. Overall, an analogy at the level of intuition seems more revealing than either the Koopman or the Gaussian construction for our work in this paper.
Beyond Theorem A, we consider two other notions of entropy from ergodic theory, and develop their analogs for unitary representations of groups or representations of other C*-algebras. The first of these notions is Kieffer’s definition of entropy using a ‘random past’ from [96, Theorem 3], which leads to Theorem B. The second is Bowen’s ‘sofic entropy’ from [25], which leads to Theorem C. In both cases, we show that these notions of entropy are given by $\log\Delta\varphi_{\mathrm{ac}}$ for some positive functional $\varphi$, except in Theorem C a certain degeneracy may occur and then the entropy equals $\infty$- ∞ (also reflecting a known feature of sofic entropy).
Each of these theorems may be seen as a different ‘non-commutative’ version of Theorem S. In ergodic theory, various predecessors of our results are already discussed this way in the literature, for example in [110, 104, 80].
1.3 Subdiagonal subalgebras and random orders
Theorem S has many proofs. Most of the classical ones begin with the following reformulation, which depends on the total ordering of $\mathbf{Z}$.
Consider again the ingredients in Theorem S. The symmetries of those Toeplitz matrices and the Schur determinant formula give the relation
$\frac{D_{n+1}}{D_{n}}=\|1_{\mathbf{T}}-P_{n}(1_{\mathbf{T}})\|^{2},$
where $P_{n}$ is the orthogonal projection from $L^{2}(\mu)$ to $\mathrm{span}\{z,\dots,z^{n}\}$ [136, Theorem 1.5.11]. It follows that
$\lim_{n\to\infty}D_{n}^{1/n}=\lim_{n\to\infty}\frac{D_{n+1}}{D_{n}}=\|1_{\mathbf{T}}-P(1_{\mathbf{T}})\|^{2},$
where $P$ is the orthogonal projection from $L^{2}(\mu)$ to
$N:=\overline{\mathrm{span}}\{z,z^{2},\dots\}$
(taking the closure in $L^{2}(\mu)$).
This orthogonal projection is the closest point of $N$ to the function $1_{\mathbf{T}}$, so Theorem S is equivalent to
[…]
This is the form in which Szegő’s theorem most often appears in the literature on analytic functions, such as in [67, Section V.8] or [68, Theorem IV.3.1]. Those references give essentially Szegő’s own proof, which he published first for the special case $\mu\ll m$, and then much later for the general case by incorporating arguments of Kolmogorov and Krein to handle $\mu_{\mathrm{sing}}$. Two further variations on this proof are recounted in [136, Sections 2.4 and 2.5]. Meanwhile, Verblunsky gave the first proof of the general case of Theorem S in [152, 151] (see also [136, Section 2.3]).
Viewed from ergodic theory, the reformulation (1.4) is the analog of the formula for the entropy rate of a stationary finite-valued stochastic process $(\xi_{n})_{n=-\infty}^{\infty}$ in terms of its ‘past’:
$\mathrm{h}(\xi)=\mathrm{H}(\xi_{0}\mid\xi_{-1},\xi_{-2},\dots).$ (1.5)
See, for instance, [157, Theorem 4.14].
To generalize Szegő’s theorem to positive definite functions on $\mathbf{Z}^{d}$, one can use the ‘past’ defined by a lexicographic ordering. This approach was developed by Helson and Lowdenslager in [85, 86] (with some later refinements in [109]). See also [105, 51] for alternative proofs under some additional assumptions on the positive definite functions. In ergodic theory, the same use of the lexicographic ordering appeared in some early works on the entropy of measure-preserving $\mathbf{Z}^{d}$-actions such as [39, 91].
This idea has a further natural generalization to any countable group $\Gamma$ that admits a left-invariant total order. But some countable groups do not, including all groups that are not torsion-free. To remove the need for this assumption, one can instead couple to a stationary random ordering of the group. All countable groups admit at least one of these: the ‘Bernoulli random order’, described in Example 7.5 below.
In ergodic theory, conditioning on the ‘past’ of a stationary random order can sometimes serve as a replacement for the formula (1.5). If $\Gamma$ is amenable, this idea gives rise to another classic formula for the entropy rate of a process due to Kieffer [96]. Our next main theorem develops the analog of this idea for positive definite functions. We find once again that the resulting quantity always agrees with the expected Fuglede–Kadison determinant, even if $\Gamma$ is not amenable.
Before formulating the theorem precisely, let us motivate it further via a related finite-dimensional calculation. Let $n$ and $k$ be positive integers. Consider $nk$ vectors in some Hilbert space $H$ indexed as an $n$-by-$k$ array, say
$V=[x_{m,i}:\ m=1,\dots,n,\ i=1,\dots,k].$
Let $V_{m}$ be the $k$-tuple $[x_{m,1},\dots,x_{m,k}]$ for each $m$. Then the Gram matrix of $V$ has this $n$-by-$n$ block structure:
$V^{\ast}V=\left[\begin{array}[]{cccc}V_{1}^{\ast}V_{1}&V_{1}^{\ast}V_{2}&\dots&V_{1}^{\ast}V_{k}\\ V_{2}^{\ast}V_{1}&V_{2}^{\ast}V_{2}&\dots&V_{2}^{\ast}V_{k}\\ \vdots&\vdots&&\vdots\\ V_{k}^{\ast}V_{1}&V_{k}^{\ast}V_{2}&\dots&V_{k}^{\ast}V_{k}\end{array}\right].$
Let
[…]
and let $R_{m}$ be the orthogonal projection onto $N_{m}$ and $R_{m}^{\perp}:=I-R_{m}$. We can express the determinant of $V^{\ast}V$ in terms of these subspaces by an iterated appeal to Schur’s determinantal formula (recalled as (2.3) below). Taking logarithms and normalizing, the result is
$\frac{1}{n}\log\det(V^{\ast}V)=\frac{1}{n}\sum_{m=1}^{n}\log\det((R_{m}^{\perp}V_{m})^{\ast}(R_{m}^{\perp}V_{m})),$ (1.6)
allowing both sides to equal $\infty$- ∞ if necessary. Within the analogy between Gram matrices and joint distributions from Subsection 1.2, this is the analog of the chain rule for discrete Shannon entropy.
Formula (1.6) remains valid under any re-ordering of the tuples $V_{1}$, …, $V_{n}$, and so we can now take an expectation over a uniform random order on the right-hand side. Then the symmetry of the random order lets us replace the average of $n$ terms with a single average over orders. Using $\omega$ to denote a permutation of $1$, …, $n$, and now writing $R_{\omega}$ for the orthogonal projection onto the random subspace
[…]
we arrive at
$\frac{1}{n}\log\det(V^{\ast}V)=\frac{1}{n!}\sum_{\omega\in S_{n}}\log\det((R_{\omega}^{\perp}V_{1})^{\ast}(R_{\omega}^{\perp}V_{1})).$
Theorem B is an infinite-dimensional, equivariant generalization of this formula. Let $\Gamma$ be a countable group, let $(\Omega,\mu)$ be a standard probability space with a measure-preserving $\Gamma$-action, and let ${(<_{\omega}:\ \omega\in\Omega)}$ be an equivariant random total order of $\Gamma$. These notions are introduced carefully in Subsections 7.2 and 7.3. Let $\varphi$ be an $\mathrm{M}_{k}$-valued positive definite function on $\Gamma$, and suppose it is associated to the representation $\pi$ by the $k$-tuple $x_{1}$, …, $x_{k}$ in $H_{\pi}$. Finally, for each $\omega$, let $R_{\omega}$ be the orthogonal projection from $H_{\pi}$ to the closed subspace
$\overline{\mathrm{span}}\{\pi(g)x_{i}:\ g<_{\omega}e,\ i=1,\dots,k\}.$
Theorem B.
In the situation above, we have
$\int\log\det[\langle R_{\omega}^{\perp}x_{j},R_{\omega}^{\perp}x_{i}\rangle]\ d\mu(\omega)=\log\Delta\varphi_{\mathrm{ac}}.$ (1.7)
For example, when $k=1$ this simplifies to
$2\int\log\|x-R_{\omega}x\|\ d\mu(\omega)=\log\Delta\varphi_{\mathrm{ac}}.$
From the viewpoint of ergodic theory, Theorem B is somewhat surprising, because for a non-amenable group $\Gamma$ it gives much better behaviour than the original notion of random-order entropy for a finite-valued stationary processes. Random-order entropy does agree with Rokhlin or sofic entropy (discussed below) for special examples of processes, such as those in [12, 4]. But in general it is only an upper bound for those quantities, and it need not be invariant under isomorphism [133, Section 7]. However, for a positive definite function $\varphi$ on a group $\Gamma$ for which both Theorems B above and C below can be applied, they always give the same value $\log\Delta\varphi_{\mathrm{ac}}$ for their respective notions of entropy. Since Theorem B concerns the analog of random-order entropy, and Theorem C concerns the analog of sofic entropy, this suggests that the worst pathologies from ergodic theory do not appear for unitary representations.
It turns out that most of the work towards Theorem B can already be found in the literature. It takes place in the context of Arveson’s theory of ‘subdiagonal subalgebras’ of finite von Neumann algebras [9]. These are a noncommutative generalization of the inclusion of the Hardy space $H^{\infty}(\mathbf{T})$ in $L^{\infty}(\mathbf{T})$. As such, they provide an abstract notion of a ‘past’ in a general finite von Neumann algebra. We recall their definition precisely in Subsection 7.5.
Arveson conjectured a generalization of Szegő’s theorem to this noncommutative setting in [9, item 4.4($\gamma$)]. As far as I know, this was the first reference to suggest that the log-integral should be replaced with a Fuglede–Kadison determinant. Arveson proved this generalization in a few of his examples in [9, Section 5], including some that recover Helson and Lowdenslager’s results for $\mathbf{Z}^{d}$. He also showed how to construct subdiagonal subalgebras for any countable group with an invariant total order, but did not prove his Szegő formula in this generality.
Arveson’s conjectured Szegő theorem was finally proved in general by Labuschagne in [100], in the equivalent form of an abstract ‘Jensen formula’. We recall this theorem as Theorem 7.3 below.
We prove Theorem B in Section 7. We first show that the singular part of $\varphi$ can be discarded without changing the left-hand side of (1.7). Then we form the von Neumann crossed product of $\Gamma$ with its action on $L^{\infty}(\mu)$, and construct a maximal subdiagonal subalgebra of this crossed product using the random total order. This is a generalization of Arveson’s construction for ordered groups, which can be regarded as the case in which $(\Omega,\mu)$ is trivial. Finally, applying Theorem 7.3 to this subdiagonal subalgebra leads to Theorem B.
1.4 Almost periodic entropy
Since its introduction in [25], Bowen’s notion of sofic entropy has taken a central place in the ergodic theory of actions of non-amenable groups. This is the final notion of entropy that we pursue across the analogy between ergodic theory and representation theory in this paper.
Consider again a countable group $\Gamma$. It is sofic if it admits a certain kind of approximation using ‘near actions’ by permutations of finite sets. This property was introduced by Gromov in [76], and first called ‘sofic’ by Weiss in [159]. Given a finite set $A$ and a stationary $A$-valued process over $\Gamma$, Bowen’s idea was to fix such a sequence of approximating near actions on finite sets $V_{1}$, $V_{2}$, …, and then count the $A$-colourings of each $V_{n}$ whose ‘local statistics’ under those near actions are close to the law of the process. Very roughly, the lim-sup exponential growth rate of these counts defines the sofic entropy of the process along that sequence of near actions. The surveys [28, 26] offer a thorough account.
Our next subject is an analogous construction for a unitary representation $\pi$ of $\Gamma$ and a tuple $v_{1},\dots,v_{k}\in H_{\pi}$. This time we approximate by a sequence of finite-dimensional representations, and for each of these we consider the vectors that give roughly the same positive definite function on $\Gamma$ as $v_{1}$, …, $v_{k}$ do in $\pi$.
In pursuing this idea, we quickly find that it is easier to allow greater generality still. We fix a separable, unital C*-algebra $\mathfrak{A}$ and consider all its unital representations on separable Hilbert spaces. The case of a group $\Gamma$ is recovering by taking $\mathfrak{A}$ equal to $C^{\ast}\Gamma$. However, even in that case certain helpful auxiliary constructions lead us to consider other C*-algebras as well.
Given a representation $\pi$ of $\mathfrak{A}$ and vectors $v_{1}$, …, $v_{k}\in H_{\pi}$, define their type to be the $\mathrm{M}_{k}$-valued map
$\Phi^{\pi}_{v_{1},\dots,v_{k}}(a):=[\langle\pi(a)v_{j},v_{i}\rangle]_{i,j=1}^{k}\qquad(a\in\mathfrak{A}).$
This is a completely positive map on $\mathfrak{A}$, and any $\mathrm{M}_{k}$-valued completely positive map arises this way by Stinespring’s theorem. Next, if $O$ is any set of such completely positive maps for a fixed value of $k$, and $\pi$ is any representation, then we define
$\mathcal{X}(\pi,O):=\{(v_{1},\dots,v_{k})\in H_{\pi}^{k}:\ \Phi^{\pi}_{v_{1},\dots,v_{k}}\in O\}.$
Imagining that $O$ is a small neighbourhood of a particular map $\varphi$, this is the analog of a set of ‘good models’ for a given shift-invariant measure in sofic entropy theory.
Finally, consider a sequence $\boldsymbol{\pi}=(\pi_{n})_{n\geq 1}$ of representations of $\mathfrak{A}$ whose dimensions $d_{n}$ are finite but diverge. We refer to it as an almost periodic sequence for $\mathfrak{A}$. See Subsection 8.4, which also discusses possible modes of convergence for such a sequence.
For a completely positive map $\varphi$, we define the almost periodic entropy of $\varphi$ along $\boldsymbol{\pi}$ to be
$\mathrm{h}_{\boldsymbol{\pi}}(\varphi):=\inf_{O}\limsup_{i\to\infty}\frac{1}{d_{i}}\log\frac{\mathrm{vol}_{2kd_{i}}\mathcal{X}(\pi_{i},O)}{v(d_{i})^{k}},$
where $O$ ranges over neighbourhoods of $\varphi$, $\mathrm{vol}_{2kd_{i}}$ refers to Lebesgue measure in $2kd_{i}$ real dimensions, and $v(d_{i})$ is the volume of the unit ball in $\mathbf{C}^{d_{i}}$. See Definition 9.4 for a full explanation. The formula above is a direct analog of the usual definition of sofic entropy for a finite-valued stationary process over a countable sofic group.
We introduce this new notion of entropy carefully in Section 9, and develop its basic properties. Many of these resemble properties of sofic entropy, but some differences emerge. For example, almost periodic entropy is not an invariant of unitary equivalence of representations, but it does satisfy a transformation formula if one changes cyclic vector within a fixed representation.
Our final main result is a formula for $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ as a Fuglede–Kadison determinant. It holds whenever $\varphi$ is ‘asymptotically associated’ to $\boldsymbol{\pi}$ and the pulled-back traces $d_{i}^{-1}\mathrm{tr}_{d_{i}}\circ\pi_{i}$ converge to a limiting tracial state $\tau$ of $\mathfrak{A}$. ‘Asymptotic association’ means that, for every neighbourhood $O$ of $\varphi$, the set $\mathcal{X}(\pi_{i},O)$ is nonempty for infinitely many $i$; if this fails then $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ is simply forced to be $\infty$- ∞.
Theorem C.
Suppose that $d_{n}^{-1}\mathrm{tr}_{d_{n}}\circ\pi_{n}\to\tau$ and that $\varphi$ is asymptotically associated to $\boldsymbol{\pi}$. Let $\Delta$ be the Fuglede–Kadison determinant defined from $\tau$, and let $\varphi_{\mathrm{ac}}+\varphi_{\mathrm{sing}}$ be the Lebesgue decomposition of $\varphi$ relative to $\tau$. Then
$\mathrm{h}_{\boldsymbol{\pi}}(\varphi)=\log\Delta\varphi_{\mathrm{ac}}.$
Let us emphasize two features of Theorem C that are substantially different from Theorems A and B:
  • Any tracial positive functional $\tau$ on $\mathfrak{A}$ may appear in Theorem C, provided it can arise as a suitable limit of normalized finite-dimensional traces. By contrast, Theorems A and B refer specifically to the regular character on a group $\Gamma$ and its associated tracial functional on $C^{\ast}\Gamma$.
  • Theorem C does not make any assumption on $\mathfrak{A}$ itself that corresponds to soficity of a group. However, such an assumption is implicit in the hypothesis that $\tau$ is a limit of normalized finite-dimensional traces. By applying this theorem to $C^{\ast}\Gamma$ when $\Gamma$ is a free group and $\tau$ is lifted from a quotient group of $\Gamma$, one recovers a theorem for precisely Rădulescu’s class of hyperlinear groups from [125]: see Subsection 8.4.
In comparison with Theorem S, Theorem C has the interesting new feature that $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ may equal $\infty$- ∞ if $\varphi$ is not asymptotically associated to $\boldsymbol{\pi}$, even though $\log\Delta\varphi_{\mathrm{ac}}$ may still be finite in that case. If $\mathfrak{A}=C^{\ast}\Gamma$ and $\tau$ is the state given by the regular character of $\Gamma$, then this is possible only if $\Gamma$ is non-amenable, and reflects basic features of the representation theory of non-amenable groups.
After proving Theorem C, Subsection 9.6 includes various consequences, for example concerning different possible modes of convergence for the sequence $\boldsymbol{\pi}$ itself.
As far as I know, the nearest precursors to Theorem C in the literature are formulas for the sofic entropy of certain special measure-preserving systems in [110, 111] and especially [80, 84]. We compare our work with these in Subsection 9.7.
Acknowledgements
This work depended on insightful conversations and correspondence with many people. In this regard I am particularly grateful to Nir Avni, Uri Bader, David Blecher, Lewis Bowen, Peter Burton, Amir Dembo, Louis Labuschagne, Michael Magee, Magdalena Musat, Narutaka Ozawa, Sorin Popa, Mikael Rørdam, Brandon Seward and Dima Shlyakhtenko.
For the purpose of open access, the author has applied a Creative Commons Attribution (CC-BY) licence to any Author Accepted Manuscript version arising from this submission.
2 Notation, conventions, and basic facts
This section sets some notation that we use throughout, some of which is slightly non-standard. Readers may prefer to skip it and then refer back to it as necessary.
2.1 Linear algebra
We assume standard ideas from linear algebra and matrix analysis. For definiteness I use [88] as a reference wherever possible
Throughout this work, our focus is restricted to linear algebra and functional analysis over $\mathbf{C}$ rather than $\mathbf{R}$. This is the appropriate choice for studying unitary representations and C*-algebras.
At some points below it is convenient to distinguish between row and column vectors. In those cases we regard the elements of $\mathbf{C}^{k}$ as row vectors by default, and write $\mathbf{C}^{\oplus k}$ for the space of height-$k$ column vectors when these are needed specifically. More generally, if $H$ is a vector space, then we write $H^{\oplus k}$ for the $k$-fold inflation of $H$, which is the vector space of height-$k$ column vectors with entries in $H$. If $S$ is a set, possibly infinite, and $H$ is a Hilbert space, then we extend this notation further by writing $H^{\oplus S}$ for the Hilbert-space direct sum of an $S$-indexed family of copies of $H$, still regarded as a space of column vectors. This insistence on column vectors is slightly unusual in functional analysis, but for finite $k$ it enables us to use matrix-vector notation from linear algebra in places where it simplifies the exposition.
We write $\mathrm{M}_{n,k}$ for the space of $n$-by-$k$ matrices over the complex numbers, and identify these with linear maps from $\mathbf{C}^{\oplus k}$ to $\mathbf{C}^{\oplus n}$ using matrix-vector multiplication. By writing such a matrix as $[v_{1},\dots,v_{k}]$, where $v_{1}$, …, $v_{k}$ are its columns, we can identify it with a $k$-tuple of vectors in $\mathbf{C}^{\oplus n}$. We generalize this notation further by allowing columns from any vector space $H$, so a linear map from $\mathbf{C}^{\oplus k}$ to $H$ may still be written in the form $[v_{1},\dots,v_{k}]$. Moreover, we sometimes abuse notation by calling the linear map $V$ itself a ‘$k$-tuple of vectors in $H$’. If $H$ is an inner product space, then the adjoint $V^{\ast}$ is the map from $H$ to $\mathbf{C}^{\oplus k}$ whose output coordinates are given by the inner products with the vectors $v_{i}$.
We abbreviate $\mathrm{M}_{k,k}$ to $\mathrm{M}_{k}$ and regard it as a $\ast$-algebra over $\mathbf{C}$ in the usual way. We write $I_{k}$ for the $k$-by-$k$ identity matrix. We write $\mathrm{tr}_{k}$ and $\mathrm{det}$ for the usual trace and determinant on any such algebra, and we set
$\overline{\mathrm{tr}}_{k}M:=k^{-1}\mathrm{tr}_{k}\,M\qquad(M\in\mathrm{M}_{k}).$
We write $\mathrm{M}_{k+}$ for the closed cone of positive semidefinite members of $\mathrm{M}_{k}$. It defines the positive definite ordering on self-adjoint matrices (also called the ‘Loewner ordering’). If $Q\in\mathrm{M}_{k+}$, then its determinant and trace are related by the inequality
$(\det Q)^{1/k}\leq\overline{\mathrm{tr}}_{k}Q.$ (2.1)
This is simply the inequality of arithmetic and geometric means applied to the eigenvalues of $Q$.
[…]
If $P$ is an orthogonal projection in a Hilbert space $H$, then we use $P^{\perp}$ as a shorthand for $I-P$.
2.2 Gram matrices and Schur complements
Let $V=[v_{1},\dots,v_{k}]$ be a $k$-tuple in a Hilbert space $H$, interpreted as a linear map from $\mathbf{C}^{\oplus k}$ to $H$. Its Gram matrix is this $k$-by-$k$ matrix of inner products:
$[\langle v_{j},v_{i}\rangle]_{i,j=1}^{k}=V^{\ast}V.$ (2.2)
Now assume in addition that $M$ is a closed subspace of $H$, and let $P$ be the orthogonal projection onto $M^{\perp}$ (notice the orthogonal complement). Suppose that we are allowed to perturb the entries of $V$ inside $M$. Any such perturbation may be written in the form
$Z=PV+Y$
for some $k$-tuple $Y$ in $M$. The Gram matrix of this perturbation is
$Z^{\ast}Z=(PV)^{\ast}(PV)+V^{\ast}PY+Y^{\ast}PV+Y^{\ast}Y=(PV)^{\ast}(PV)+Y^{\ast}Y,$
where the second equality holds because $PY$ and its adjoint $Y^{\ast}P$ are both zero.
In particular, this shows that $PV$ uniquely minimizes the Gram matrix $Z^{\ast}Z$ among all such perturbation in the positive definite ordering. It therefore also has minimal determinant among the Gram matrices of these perturbations [88, Corollary 7.7.4(e)]. The Gram matrix of $PV$ is an abstract form of the Schur complement of $V^{\ast}V$ relative to $M$. If $M$ is the span of another tuple $U=[u_{1},\dots,u_{\ell}]$, and we write $W$ for the combined tuple $[u_{1},\dots,u_{\ell},v_{1},\dots,v_{k}]$, then the Schur complement participates in Schur’s determinantal formula:
$\det(W^{\ast}W)=\det(U^{\ast}U)\cdot\det((PV)^{\ast}(PV)).$ (2.3)
See [88, Subsection 0.8.5 and Exercises 7.1.P28 and 7.3.P8].
2.3 Landau’s asymptotic notation
If $g:\mathbf{N}\to(0,\infty)$, then we write $o(g(n))$ as a placeholder for any other function $f:\mathbf{N}\to\mathbf{R}$ that satisfies $f(n)/g(n)\to 0$.
2.4 Measure and integration in high dimensions
For any dimension $d$, we write $\mathrm{vol}_{d}$ for Lebesgue measure on $\mathbf{R}^{d}$. For any positive integers $d$ and $k$, we also write $\mathrm{vol}_{2kd}$ for the measure on $\mathrm{M}_{k,d}$ obtained by identifying this space with $\mathbf{R}^{2kd}$. We write $\mathrm{S}^{d-1}$ for the unit sphere in $\mathbf{R}^{d}$ or in any other vector space that has a standing identification with $\mathbf{R}^{d}$. We write $\sigma_{d-1}$ for the surface-area measure on $\mathrm{S}^{d-1}$ normalized to have total mass $1$, and we refer to an integral with respect to $\sigma_{d-1}$ as a ‘spherical average’.
3 Operator algebras, C*-algebras and representations
Theorem C is formulated in the generality of abstract C*-algebras, but Theorems A and B concern the special case of matrix-valued positive definite functions on groups. This section and Section 4 make preparations for these results in terms of general C*-algebras. Then Section 5 describes that special case of group algebras more fully before we begin work on proving Theorem A.
Although C*-algebras are the natural generality for much of our work below, I have tried to avoid assuming very advanced parts of that theory. This section recalls most of the background we need. Where possible I cite the standard texts [37, 124] on functional analysis and [38, 48, 49] on C*-algebras and operator algebras. I cite a few specific results from original papers.
Readers familiar with these topics may wish to skip this section and refer back to it as necessary.
3.1 C*-algebras, von Neumann algebras, and affiliated operators
Throughout this paper, $\mathfrak{A}$ is a separable, unital C*-algebra and we study separable representations, meaning that they act on separable complex Hilbert spaces. We usually denote the unit of $\mathfrak{A}$ by $1$. We usually denote a representation by a single letter such as $\pi$, and then write its Hilbert space as $H_{\pi}$ when necessary.
Our guiding example is the group C*-algebra $C^{\ast}\Gamma$ of a countable group $\Gamma$. This is the maximal C*-completion of the group algebra $\mathbf{C}[\Gamma]$, so it is indeed separable and unital. Our results about almost periodic entropy in Section 9 could also apply to the unital augmentation of the group C*-algebra of a locally compact, second countable group.
Once a particular representation $\pi$ of $\mathfrak{A}$ is being considered, operator theory provides many auxiliary constructions of operators in $\mathfrak{B}(H_{\pi})$, for instance via the Borel functional calculus or symmetry considerations. These often fall outside the operator-norm closure of $\pi(\mathfrak{A})$, which is a C*-algebra, but within its bi-commutant $\pi(\mathfrak{A})^{\prime\prime}$, which agrees with the weak-operator closure of $\pi(\mathfrak{A})$ and is a von Neumann algebra. In this paper we adhere rather strictly to the following convention:
  • C-algebras may exist in the abstract;
  • a von Neumann algebra is a weak-operator closed $\ast$-subalgebra of $\mathfrak{B}(H)$ for some particular Hilbert space $H$.
In particular, we may be somewhat casual about identifying C-algebras that are isomorphic, but two ‘copies’ of a von Neumann algebra acting on different Hilbert spaces are regarded as different von Neumann algebras.
Some of the constructions we need later involve operators on a Hilbert space that are closed and densely defined but unbounded. The basic theory of these can be found in [37, Chapter X], including the unbounded version of the spectral theorem [37, Section X.4] which leads to an unbounded extension of the polar decomposition [37, Exercise X.4.6] (or see [124, Section VIII.9]).
A closed and densely defined operator $T$ is affiliated to the von Neumann algebra $\mathfrak{N}$ if it commutes with every unitary element of $\mathfrak{N}^{\prime}$: see [49, Exercise I.1.10]. This includes the assertion that those unitary elements preserve $\mathrm{dom}\,T$. The collection of operators affiliated to $\mathfrak{N}$ is closed under forming adjoints, and if such an operator $T$ has polar decomposition $U|T|$, then the uniqueness of that decomposition implies that $U\in\mathfrak{N}$ and that $|T|$ is also affiliated to $\mathfrak{N}$. If $\mathfrak{N}$ is a finite von Neumann algebra, then sums and products of affiliated operators are also still affiliated operators, provided each of these constructions is followed by taking graph closures [49, Exercise III.1.13.c].
3.2 Representations and positive functionals
If $\pi$ is a representation of $\mathfrak{A}$ and $M$ is a closed $\pi$-invariant subspace of $H_{\pi}$, then we write $\pi^{M}$ for the associated subrepresentation, meaning that $\pi^{M}(a)=\pi(a)|M$ for each $a\in\mathfrak{A}$. A representation $\pi$ is irreducible if its only closed invariant subspaces are $\{0\}$ and $H_{\pi}$.
We use $\oplus$ to denote direct sums in the categories of Hilbert spaces or representations [48, Subsection 2.2.3], and we use $\otimes$ for tensor products of Hilbert spaces or of operators or von Neumann algebras on them [49, Sections I.2.3–4]. We do not need the more involved theory of tensor products of abstract C*-algebras. For a representation $\pi$ and any positive integer $k$, we write either $\pi^{\oplus k}$ or $\pi\otimes I_{k}$ for the direct sum of $k$ copies of $\pi$. We extend this notation to allow $k=\infty$ by taking $I_{\infty}$ to be the identity on some fixed choice of infinite-dimensional separable complex Hilbert space. We refer to $\pi^{\oplus k}$ as the $k$-fold inflation of $\pi$ (this construction is called ‘ampliation’ in [48]).
If $\pi$ is a representation of $\mathfrak{A}$, then a subset $S$ of $H_{\pi}$ is cyclic for $\pi$ if it is not contained in any proper closed invariant subspace of $H_{\pi}$, or equivalently if
$\overline{\sum_{v\in S}\pi(\mathfrak{A})v}=H_{\pi}.$
If $\pi$ and $\rho$ are two representations of $\mathfrak{A}$, then:
  • $\pi$ is equivalent to $\rho$, written $\pi\simeq\rho$, if there is a unitary operator from $H_{\pi}$ to $H_{\rho}$ that intertwines $\pi$ with $\rho$;
  • $\pi$ is contained in $\rho$, written $\pi\lesssim\rho$, if $\pi$ is equivalent to a subrepresentation of $\rho$;
  • $\pi$ and $\rho$ are disjoint, written $\pi\spoon\rho$, if no nontrivial subrepresentations of them are equivalent.
Another possible relation is that $\pi$ is contained in an inflation of $\rho$, which we can always write as $\pi\lesssim\rho^{\oplus\infty}$ if both representations are separable.
The space of all equivalence classes of irreducible representations is called the spectrum of $\mathfrak{A}$, and is denoted by $\widehat{\mathfrak{A}}$.
Now let $\kappa:=\pi\oplus\rho$. Regard $H_{\pi}$ and $H_{\rho}$ as invariant subspaces of $H_{\kappa}$, and let $P$ be the orthogonal projection from $H_{\kappa}$ onto $H_{\pi}$. The next result is [48, Proposition 5.2.4].
Lemma 3.1.
We have $\pi\spoon\rho$ if and only if $P$ lies in the centre of $\kappa(\mathfrak{A})^{\prime\prime}$. ∎
If $\pi$ is a representation of $\mathfrak{A}$ and $v\in H_{\pi}$, then the formula
$\Phi^{\pi}_{v}(a):=\langle\pi(a)v,v\rangle$ (3.1)
defines a positive linear functional on $\mathfrak{A}$. We write that the functional in (3.1) is associated to $\pi$ by $v$. Alternatively, if the representation $\pi$ is understood, we adapt a term from information theory by calling $\Phi^{\pi}_{v}$ the type of $v$ (compare [40, Section 11.1], for example).
All positive linear functionals on $\mathfrak{A}$ arise this way because of the GNS construction [48, Section 2.4]. We denote the GNS representation of $\varphi$ by $\pi_{\varphi}$. By the uniqueness of the GNS construction, $\varphi$ is associated to $\pi$ if and only if $\pi_{\varphi}$ is contained in $\pi$.
We write $\mathfrak{A}^{\ast}_{+}$ for the space of positive linear functionals on $\mathfrak{A}$, and always endow it with the weak topology inherited from $\mathfrak{A}$. We write $\Sigma(\mathfrak{A})$ for the further subset of positive functionals $\varphi$ that satisfy $\varphi(1)=1$. Those functionals are called states, and $\Sigma(\mathfrak{A})$ is the state space of $\mathfrak{A}$. These standard notions and their relation to representations can be found in [48, Chapter 2].
Given any two continuous linear functionals $\varphi$ and $\psi$ on $\mathfrak{A}$, we write $\varphi\leq\psi$ if the functional $\psi-\varphi$ is positive. This defines a partial order on $\mathfrak{A}^{\ast}$ in which $\mathfrak{A}^{\ast}_{+}$ is the non-negative cone. For positive functionals, this relation implies containment of GNS representations by the following basic result (see [48, Proposition 2.5.1]).
Proposition 3.2.
Assume that $\psi$ is associated to $\pi_{\psi}$ by $v$. Let $c>0$. Then $\varphi\leq c\psi$ if and only if there exists $T\in\pi_{\psi}(\mathfrak{A})^{\prime}$ such that (i) $0\leq T\leq c$ and (ii) $\varphi$ is associated to $\pi_{\psi}$ by $Tv$. If such a $T$ exists, then it is unique.
In particular, if $\varphi\leq c\psi$ for some $c>0$, then $\varphi$ is associated to $\pi_{\psi}$. ∎
We also need two other relations involving positive functionals and representations:
  • If $\pi$ is a representation and $\varphi$ is a positive functional, then $\varphi$ is $\pi$-normal if $\varphi$ is associated to $\pi^{\oplus\infty}$. This is equivalent to $\varphi$ having the form $\widetilde{\varphi}\circ\pi$ for some normal positive functional $\widetilde{\varphi}$ on $\pi(\mathfrak{A})^{\prime\prime}$, by the general characterization of normal positive functionals on a von Neumann algebra [38, Theorem 46.4].
  • On the other hand, $\varphi$ is $\pi$-singular if $\pi_{\varphi}\spoon\pi$. If $\pi=\pi_{\psi}$, then we call $\varphi$ and $\psi$ themselves disjoint.
The next proposition describes general functionals in terms of normal and singular ones.
Proposition 3.3.
Let $\pi$ be a representation and $\varphi$ a positive functional. Then there is a unique decomposition
$\varphi=\varphi_{\mathrm{ac}}+\varphi_{\mathrm{sing}}$ (3.2)
into positive summands such that $\varphi_{\mathrm{ac}}$ is $\pi$-normal and $\varphi_{\mathrm{sing}}$ is $\pi$-singular. Each of $\varphi_{\mathrm{ac}}$ and $\varphi_{\mathrm{sing}}$ is linear as a function of $\varphi$. ∎
Proposition 3.3 is from [141]. The heart of the proof is a corresponding decomposition of $\pi_{\varphi}$ relative to $\pi$. That decomposition can also be found as [112, Theorem 1.11] (phrased there in the special case of group representations, but with a proof that works in general).
If $X$ is a compact metrizable space and $\mathfrak{A}=C(X)$, then the Riesz–Markov theorem identifies positive linear functionals on $\mathfrak{A}$ with finite Borel measures on $X$. Through this identification, routine steps reduce Proposition 3.3 to the Lebesgue decomposition from measure theory. With this in mind, we call (3.2) the Lebesgue decomposition of $\varphi$ relative to $\pi$. This also motivates the choice of subscripts for the two components in (3.2).
Starting from that example, the study of functionals on a C*-algebra or von Neumann algebra has long been regarded as a ‘non-commutative’ extension of the theory of measure and integration. See [132, 131] and [47] for early developments of this point of view.
To accompany the Lebesgue decomposition, there are also non-commutative extensions of the Radon–Nikodym theorem. These describe the structure of $\pi$-normal functionals in terms of vectors or operators related to $\pi$. There are actually several such theorems that allow for the non-commutativity of $\mathfrak{A}$ in different ways. Proposition 3.2 is sometimes classed as one of these theorems, although the assumption $\varphi\leq c\psi$ is typically much stronger than $\varphi$ being $\pi_{\psi}$-normal. In this work we also need another such theorem, but it depends on the extra structure of a tracial vector: see Proposition 3.5 below.
Further references.
[…]
3.3 Tracial functionals and vectors
A linear functional $\tau$ on $\mathfrak{A}$ is tracial if
$\tau(ab)=\tau(ba)\qquad(a,b\in\mathfrak{A}).$ (3.3)
The study of these is motivated by two fundamental sources of examples. First, if $\pi$ is a $d$-dimensional representation with $d$ finite, then the normalized trace on $\mathfrak{B}(\mathbf{C}^{d})$ pulls back to the tracial state $\overline{\mathrm{tr}}_{d}\circ\pi$ on $\mathfrak{A}$. Second, for group C*-algebras, tracial states correspond to characters of the group (beware that some references, including [48], use these terms slightly differently).
Suppose that a tracial positive functional $\tau$ is associated to the representation $\lambda$ by the cyclic vector $\xi$, and let $H:=H_{\lambda}$. Then (3.3) becomes an identity for $\xi$:
$\langle\lambda(b)\xi,\lambda(a^{\ast})\xi\rangle=\langle\lambda(a)\xi,\lambda(b^{\ast})\xi\rangle\qquad(a,b\in\mathfrak{A}).$ (3.4)
In any representation, a vector satisfying (3.4) is also called tracial.
The tracial property of $\xi$ gives rise to a rich structure around the von Neumann algebra $\mathfrak{N}:=\lambda(\mathfrak{A})^{\prime\prime}$ and its commutant $\mathfrak{N}^{\prime}$. First, we can take weak-operator limits in $a$ and then separately in $b$ in (3.4) to conclude that the extended positive functional
$\widetilde{\tau}(A):=\langle A\xi,\xi\rangle\qquad(A\in\mathfrak{N})$ (3.5)
is actually tracial on the whole of $\mathfrak{N}$. The right-hand side of (3.5) actually defines a normal positive functional on the whole of $\mathfrak{B}(H)$, and we denote this by $\widetilde{\tau}$ as well. Clearly $\xi$ is still cyclic for $\mathfrak{N}$. It follows that $\xi$ is also cyclic for $\mathfrak{N}^{\prime}$ [49, Corollary I.6.1], hence also separating for both $\mathfrak{N}$ and $\mathfrak{N}^{\prime}$ [49, Proposition I.1.5], and that $\mathfrak{N}$ and $\mathfrak{N}^{\prime}$ are both finite von Neumann algebra [49, Proposition I.6.9(ii)].
Next, if $\xi$ is cyclic and tracial for the von Neumann algebra $\mathfrak{N}$, then the formula
$J(A\xi):=A^{\ast}\xi\qquad(A\in\mathfrak{N})$ (3.6)
gives a well-defined involution on the subspace $\mathfrak{N}\xi$. The map $J$ is anti-linear, and it converts inner products to their conjugates as a consequence of the trace identity. In particular, $J$ is an isometry for the restriction of the norm of $H$ to $\mathfrak{N}\xi$. It therefore extends by continuity to an involution of the whole of $H$ that has the same properties. We still denote this by $J$, and call it the canonical involution associated to $\mathfrak{N}$ and $\xi$.
These constructions go back to Murray and von Neumann. The theory is recounted with more complete references in [49, Section I.5.1–2] in the alternative framework of Hilbert algebras, which are shown to be equivalent to algebras with tracial cyclic vectors in [49, Section I.6.2]. The following facts can all be found in those sections.
Lemma 3.4.
The canonical involution has the following properties:
  • i.
    $J\xi=\xi$;
  • ii.
    $JA\xi=A^{\ast}\xi$ for every $A\in\mathfrak{N}$;
  • iii.
    the map $A\mapsto JAJ$ is an involutive $\ast$-anti-automorphism of $\mathfrak{B}(H)$;
  • iv.
    the map from (iii) preserves the functional $\widetilde{\tau}$ from (3.5) and exchanges the von Neumann subalgebras $\mathfrak{N}$ and $\mathfrak{N}^{\prime}$ of $\mathfrak{B}(H)$. ∎
Here is the Radon–Nikodym theorem that compares a positive functional to a trace:
Proposition 3.5.
In the setting above, let $\varphi$ be a $\lambda$-normal positive functional. Then there is a unique non-negative operator $T$ affiliated to $\mathfrak{N}^{\prime}$ such that $\xi\in\mathrm{dom}\,T$ and $\varphi$ is associated to $\lambda$ by $T\xi$. ∎
As far as I know, a version of Proposition 3.5 first appeared as a special case of [52, Theorem 4]; see also [132, Theorem 14] and the discussion that surrounds it, or [49, Exercise III.1.14]. Proposition 3.5 really has two substantial parts, followed by a final simplification. Some references present these parts separately, so let us quickly sketch how they are put together:
  • The first part shows that, if $\varphi$ is $\lambda$-normal, then it is actually associated to $\lambda$ itself as opposed to $\lambda^{\oplus\infty}$. This holds because $\xi$ is separating, as discussed above: see [49, Theorem III.1.4]. This consequence of a separating vector is derived from the fact that, if two von Neumann algebras both have cyclic and separating vectors, then any normal isomorphism between them is spatial [49, Theorem III.1.3].
  • The second part shows that any vector in $H$ has the form $T\xi$ for a unique operator $T$ affiliated to $\mathfrak{N}^{\prime}$; see [49, Exercise III.1.13].
  • To finish the proof, one replaces $T$ with $|T|$. This does not change the associated functional $\varphi$ because the unitary part of the polar decomposition of $T$ commutes with $\lambda$. After this replacement, we may assume that $T$ is non-negative, and having done so $\varphi$ determines $T$ uniquely. This is because $\varphi$ determines its own GNS representation and associating vector uniquely up to unitary equivalence, and such a unitary equivalence then collides with the uniqueness of the polar decomposition of $T$.
If $T$ is the operator from Proposition 3.5 and $E$ is its spectral resolution on $[0,\infty)$, then the condition that $\xi\in\mathrm{dom}\,T$ is equivalent to
$\int_{[0,\infty)}t^{2}\ \widetilde{\tau}E(dt)=\int_{[0,\infty)}t^{2}\ \langle E(dt)\xi,\xi\rangle<\infty.$ (3.7)
This says that $T$ is ‘square-integrable’ as defined in [132, Subsection 3.4].
We sometimes need a variant of $T$ affiliated to $\mathfrak{N}$ rather than $\mathfrak{N}^{\prime}$. To obtain this, set $S:=JTJ$, where $J$ is the canonical involution. By Lemma 3.4, this satisfies $\xi\in\mathrm{dom}\,S$ and
$S\xi=JTJ\xi=JT\xi=T^{\ast}\xi=T\xi.$ (3.8)
Further references.
[…]
3.4 Fuglede–Kadison determinants
Let $\mathfrak{A}$ be a C-algebra and $\tau$ a tracial positive functional on it. For any positive invertible $a\in\mathfrak{A}$, its Fuglede–Kadison determinant with respect to $\tau$ is
$\Delta a:=\exp(\tau(\log a)).$
This definition is sometimes extended to other invertible elements $a$ by applying it to $|a|$, but we do not need this. See [63, 64] or [49, Section I.6.11].
If $\mathfrak{N}$ is a von Neumann algebra, $\widetilde{\tau}$ is a normal tracial positive functional on it, and $A\in\mathfrak{N}$ is positive and invertible, then $\Delta A$ can be expressed in terms of the spectral resolution of $A$. Using this expression, we can extend the domains of both $\widetilde{\tau}$ and the resulting determinant $\Delta$ to include suitable classes of affiliated operators. If $T$ is a non-negative operator affiliated to $\mathfrak{N}$, and $E$ is its spectral resolution on $[0,\infty)$, then we call $T$ integrable with respect to $\widetilde{\tau}$ if
$\int_{[0,\infty)}t\ \widetilde{\tau}E(dt)<\infty.$ (3.9)
In this case we define $\widetilde{\tau}(T)$ to be the left-hand side of (3.9). Similarly, we call $T$ log-integrable with respect to $\widetilde{\tau}$ if
$\int_{[0,\infty)}\log^{+}t\ \widetilde{\tau}E(dt)<\infty$
(notice that this controls the singularity of $\log$ at $\infty$ but not at 00). If this holds, then we define the Fuglede–Kadison determinant of $T$ with respect to $\widetilde{\tau}$ to be
$\Delta T:=\exp\int_{[0,\infty)}\log t\ \widetilde{\tau}E(dt).$ (3.10)
The integral here is well-defined by log-integrability, but it may take the value $\infty$- ∞, in which case $\Delta T:=0$. The properties of square-integrability (from (3.7)), integrability, and log-integrability are progressively weaker. This extension of the Fuglede–Kadison determinant is covered carefully in [77, Section 2].
If $\widetilde{\tau}$ is a state, then $\widetilde{\tau}E$ is a probability measure, and we can apply Jensen’s inequality for expectations of concave functions [126, Theorem 3.3] to the integral of the logarithm in (3.10). This yields an infinitary analog of the det-trace inequality (2.1) when $\widetilde{\tau}$ is a state: for any non-negative integrable operator $T$ affiliated to $\mathfrak{N}$, we have
$\Delta T\leq\widetilde{\tau}(T).$ (3.11)
If $\widetilde{\tau}$ is a normal tracial positive functional on $\mathfrak{N}$, and $\Delta$ is the associated Fuglede–Kadison determinant, then the definitions above permit the following approximation. Suppose that $T$ is a log-integrable non-negative operator affiliated to $\mathfrak{N}$, and let $E$ be its spectral resolution on $[0,\infty)$. For any $\delta\in(0,1)$, let
$T_{\delta}:=(T\vee\delta)\wedge\delta^{-1}:=\int_{[0,\infty)}(t\vee\delta)\wedge\delta^{-1}\ E(dt),$
where $\vee$’ stands for ‘$\max$’ and $\wedge$’ stands for ‘$\min$’. Then $T_{\delta}$ is an element of $\mathfrak{N}$ satisfying $\delta\leq T_{\delta}\leq\delta^{-1}$. Let $E_{+}:=E(0,\infty)=1-E\{0\}$.
Lemma 3.6.
As $\delta\downarrow 0$, these operators satisfy
  1. i.
    $T_{\delta}x\to Tx$ and $T_{\delta}^{-1}Tx\to E_{+}x$ for every $x\in\mathrm{dom}\,T$;
  2. ii.
    $\Delta T_{\delta}\to\Delta T$.
Proof.
Fix $x\in\mathrm{dom}\,T$. The spectral theorem gives
$\|Tx-T_{\delta}x\|^{2}=\int_{[0,\infty)}|t-(t\vee\delta)\wedge\delta^{-1}|^{2}\ \langle E(dt)x,x\rangle.$
The expression $\langle E(\cdot)x,x\rangle$ is a finite Borel measure on $\mathbf{R}$, and the function $t^{2}$ is integrable with respect to it because of the assumption that $x\in\mathrm{dom}\,T$. Therefore the dominated convergence theorem gives $\|Tx-T_{\delta}x\|\to 0$ as $\delta\downarrow 0$.
Another calculation from the spectral theorem gives
$\|E_{+}x-T_{\delta}^{-1}Tx\|^{2}=\int_{(0,\infty)}|1-\varphi_{\delta}(t)|^{2}\ \langle E(dt)x,x\rangle,$
where
$\varphi_{\delta}(t)=\left\{\begin{array}[]{ll}t/\delta&\quad t<\delta\\ 1&\quad\delta\leq t\leq\delta^{-1}\\ \delta t&\quad t>\delta^{-1}.\end{array}\right.$
As $\delta\downarrow 0$, the functions $|1-\varphi_{\delta}(t)|^{2}$ are all bounded by $1+t^{2}$ and converge pointwise to 00 on $(0,\infty)$. Therefore the dominated convergence theorem also gives $\|E_{+}x-T_{\delta}^{-1}Tx\|\to 0$ as $\delta\downarrow 0$. This completes the proof of (i).
Finally, since $\widetilde{\tau}$ is normal, the composition $\widetilde{\tau}E$ is a Borel measure on $[0,\infty)$, and by assumption the function $\log^{+}t$ is integrable with respect to this measure. Therefore we also have
$\log\Delta T_{\delta}=\int_{0}^{\infty}\log((\delta\vee t)\wedge\delta^{-1})\ \widetilde{\tau}E(dt)\to\log\Delta T\quad\hbox{as}\ \delta\downarrow 0,$ (3.12)
where we use the log-integrability of $T$ and the dominated convergence theorem to control the integrals over $[1,\infty)$, and the monotone convergence theorem to control the integrals over $[0,1)$. The convergence (3.12) holds even if $\log\Delta T=-\infty$, and its exponential is conclusion (ii). ∎
We now connect these preliminary results with Proposition 3.5. Let $\lambda$ be a representation of $\mathfrak{A}$ with a cyclic tracial vector $\xi$, let $\tau$ be the resulting tracial functional on $\mathfrak{A}$, and let $\widetilde{\tau}$ be its extension to $\lambda(\mathfrak{A})^{\prime\prime}$ as in formula (3.5). Let $\widetilde{\tau}$ also denote the tracial positive functional defined on $\mathfrak{N}:=\lambda(\mathfrak{A})^{\prime}$ by the same formula. Write $\Delta$ for the associated Fuglede–Kadison determinant on log-integrable operators affiliated to either von Neumann algebra. Lastly, abbreviate $\Delta(\lambda(a))$ to $\Delta a$ for $a\in\mathfrak{A}$ when the role of $\lambda$ is clear.
Let $\varphi$ be another positive functional on $\mathfrak{A}$ that is $\lambda$-normal, so it can be represented using the operator $T$ affiliated to $\mathfrak{N}$ from Proposition 3.5. The operator $T$ is an equivariant analog of the representation of a positive semidefinite matrix $Q$ as the square of another such matrix $V$. In that situation we have $\det Q=(\det V)^{2}$, and this motivates the following definition.
Definition 3.7.
The Fuglede–Kadison determinant of $\varphi$ with respect to $\tau$ is
$\Delta_{\tau}\varphi:=(\Delta T)^{2}.$
We often shorten $\Delta_{\tau}\varphi$ to $\Delta\varphi$ when the choice of $\tau$ is clear from the context.
The operator $T$ is affiliated to $\mathfrak{N}:=\lambda(\mathfrak{A})^{\prime}$, but we can switch to $\lambda(\mathfrak{A})^{\prime\prime}$ by setting $S:=JTJ$ as in (3.8). If $E$ is the spectral resolution of $T$, then Lemma 3.4(iii) identifies $JE(\cdot)J$ as the spectral resolution of $S$, and now Lemma 3.4(iv) gives
$\Delta S=\exp\int_{[0,\infty)}\log s\ \widetilde{\tau}(JE(ds)J)=\exp\int_{[0,\infty)}\log t\ \widetilde{\tau}E(dt)=\Delta T.$ (3.13)
Later we need the following consequence of Lemma 3.6 and Kaplansky’s density theorem in this context.
Corollary 3.8.
For any $\varepsilon>0$, there is a positive invertible element $a$ of $\mathfrak{A}$ such that
$\|\lambda(a)\xi-T\xi\|<\varepsilon\quad\hbox{and}\quad|\Delta\lambda(a)-\Delta T|<\varepsilon.$ (3.14)
Proof.
Let $S:=TJT$, so this satisfies (3.8) and (3.13).
By Lemma 3.6, pick $\delta>0$ small enough that
$\|S_{\delta}\xi-S\xi\|<\varepsilon\qquad\hbox{and}\qquad|\Delta S_{\delta}-\Delta S|<\varepsilon.$
Let $\mathfrak{S}_{\delta}$ be the set of self-adjoint elements $R$ of $\mathfrak{N}$ that satisfy $\delta\leq R\leq\delta^{-1}$. The function $R\mapsto\log R$ is strong-operator continuous on the set $\mathfrak{S}_{\delta}$: for instance, this holds by applying [38, Lemma 44.2] to a Lipschitz function that agrees with $\log$ on $[\delta,\delta^{-1}]$. Since $\widetilde{\tau}$ is positive and normal, it is ultraweakly continuous, and this in turn implies strong-operator continuity when restricted to any bounded subset of $\mathfrak{N}$ such as $\mathfrak{S}_{\delta}$: this follows, for instance, from the explicit description of such functionals in [38, Theorem 46.4]. By the continuity of composition, it follows that the expression
$\Delta R=\exp(\widetilde{\tau}(\log R))$
is also strong-operator continuous on $\mathfrak{S}_{\delta}$. Now the existence of $a\in\mathfrak{A}$ satisfying $\pi(a)\in\mathfrak{S}_{\delta}$ and also the approximations (3.14) follows from the Kaplansky density theorem as in [38, Theorem 44.1(c)]. ∎
As far as I know, the first definition of a Fuglede–Kadison determinant for positive functionals was Arveson’s in [9, Definition 4.3.7]. His setting is functionals on a von Neumann algebra, but the definition itself works for a general C*-algebra without change. His definition is different from ours: his expression is a variational principle. For us this becomes the following proposition.
Proposition 3.9.
Let $A$ be any dense $\ast$-subalgebra of $\mathfrak{A}$. Let $\tau$ be a tracial state on $\mathfrak{A}$, and $\Delta$ its associated Fuglede–Kadison determinant. Finally, let $\varphi$ be another positive functional on $A$, and let $\varphi_{\mathrm{sing}}+\varphi_{\mathrm{ac}}$ be the Lebesgue decomposition of $\varphi$ with respect to $\tau$. Then
$\Delta\varphi_{\mathrm{ac}}=\inf\big{\{}\varphi(a):\ a\in A\ \hbox{positive and invertible and}\ \Delta a\geq 1\big{\}}.$ (3.15)
Beware that Proposition 3.9 requires $\tau$ to be normalized.
Before proving Proposition 3.9, let us discuss a few special cases.
Firstly, in case $\mathfrak{A}=\mathrm{M}_{k}$, Proposition 3.9 is a classical variational principle relating traces and determinants: see [88, Exercise 7.8.P4], for example.
Secondly, suppose that $X$ is a compact metrizable space, that $\mathfrak{A}=C(X)$, and that $\tau$ is integration with respect to a Borel probability measure $\mu$ on $X$. Given another finite Borel measure $\varphi$ on $X$, let $\varphi_{\mathrm{ac}}$ be its absolutely continuous part with respect to $\mu$. Then, as discussed in the Introduction, the associated Fuglede–Kadison determinant is
$\Delta\varphi_{\mathrm{ac}}=\exp\int\log\frac{d\varphi_{\mathrm{ac}}}{d\mu}\ d\mu,$
and Proposition 3.9 shows that this equals
[…]
This time, we can recognize $\Delta\varphi_{\mathrm{ac}}$ as the ‘reversed’ relative entropy (also called Kullback–Leibler divergence) $S(\mu\mid\varphi_{\mathrm{ac}})$, and (3.16) is a classical variational principle for this quantity: see, for instance, [136, Lemma 2.3.3]. Alternatively, we can allow $\mathfrak{A}=L^{\infty}(\mu)$ for some standard probability space $(\Omega,\mu)$ and arrive at an analogous formula in case $\varphi\ll\mu$.
Those two classical examples combine into the following, which we need later during the proof of Theorem B.
Example 3.10.
Let $(\Omega,\mu)$ be a standard probability space, let $k$ be a positive integer, and let $\mathfrak{A}$ be $L^{\infty}(\mu;\mathrm{M}_{k})$ (see, for instance, [38, Secton 52] for properties of this algebra). Define a tracial state on $\mathfrak{A}$ by
[…]
Then the associated Fuglede–Kadison determinant is given by
$\Delta b=\exp\frac{1}{k}\int\log\det b(\omega)\ d\mu(\omega)\qquad(b\in\mathfrak{A}_{+}=L^{\infty}(\mu;\mathrm{M}_{k+})),$
with the usual convention that the integral on the right may equal $\infty$- ∞ and then $\exp(-\infty)=0$.
On the other hand, any $q\in L^{1}(\mu;\mathrm{M}_{k+})$ may be used to define a positive functional on $\mathfrak{A}$ by setting
[…]
(in fact this is the general form of a positive functional on $\mathfrak{A}$ that is normal with respect to $\tau$). For this functional, the variational principle (3.15) turns into
$\exp\frac{1}{k}\int\log\det q(\omega)\ d\mu(\omega)$
$:\displaystyle=\inf\Big{\{}\int\overline{\mathrm{tr}}_{k}(b(\omega)q(\omega))\ d\mu(\omega)$ :
$\qquad\qquad\qquad b\in L^{\infty}(\mu;\mathrm{M}_{k+})\ \hbox{invertible and}\ \int\log\det b\ d\mu\geq 0\Big{\}}.$ (3.17)
This instance of (3.15) also has an elementary proof by combining the two classical special cases mentioned previously and a measurable selection argument.
One can also adjust this example by taking $\mathfrak{A}=C(X;\mathrm{M}_{k})$ for a compact metrizable space $X$. Then a general positive functional $\varphi$ is represented by an $\mathrm{M}_{k+}$-valued Borel measure, and we insert $d\varphi_{\mathrm{ac}}/d\mu$ in place of $q$ in (3.10). ∎
The proof of Proposition 3.9 is very standard, but I include it for completeness. It has elements in common with Corollary 3.8, but it does not take place entirely inside the representation $\lambda$.
Proof of Proposition 3.9.
Let $\tau$ be associated to its GNS representation $\lambda$ by the cyclic tracial vector $\xi$, and let $\widetilde{\tau}$ be its normal extension to $\lambda(\mathfrak{A})^{\prime\prime}$ as in (3.5). Associate $\varphi_{\mathrm{ac}}$ to $\lambda$ by a vector of the form $T\xi$ as in Proposition 3.5. Let $S:=JTJ$, so this satisfies (3.8) and (3.13). In addition, let $\varphi_{\mathrm{sing}}$ be associated to its GNS representation $\pi_{\mathrm{sing}}$ by the vector $u$. Overall, $\varphi$ is associated to $\pi:=\pi_{\mathrm{sing}}\oplus\lambda$ by the vector $(u,S\xi)$.
We now prove a pair of inequalities.
Step 1. If $a\in A$ is positive, invertible, and satisfies $\Delta a\geq 1$, then
$(\Delta S)^{2}\leq\Delta S\cdot\Delta a\cdot\Delta S=\Delta(S\lambda(a)S)\leq\widetilde{\tau}(S\lambda(a)S)=\varphi(a).$
The second equality holds by the multiplicativity of $\Delta$ (see [77, Proposition 2.5]), and the second inequality is the infinitary det-trace inequality (3.11) (this is where we need the assumption that $\tau$ is normalized). This proves the inequality $\leq$”.
Step 2. Let $S_{\delta}:=(S\vee\delta)\wedge\delta^{-1}$ as in Lemma 3.6, and define the operator
$R_{\delta}:=\delta\ \oplus\ (1+\delta)(\Delta S_{\delta})S_{\delta}^{-1}\in\mathfrak{B}(H_{\pi}).$
The projection from $H_{\pi}$ to $H_{\lambda}$ lies in $\pi(\mathfrak{A})^{\prime}\cap\pi(\mathfrak{A})^{\prime\prime}$ by Lemma 3.1, and $S_{\delta}$ commutes with $\lambda(\mathfrak{A})^{\prime}$ by Lemma 3.4(iv). It follows that the whole operator $R_{\delta}$ commutes with $\pi(\mathfrak{A})^{\prime}$, and so it lies in $\pi(\mathfrak{A})^{\prime\prime}$.
Lemma 3.6 gives
$\langle R_{\delta}^{2}(u,S\xi),(u,S\xi)\rangle$ $=\delta^{2}\|u\|^{2}+(1+\delta)^{2}(\Delta S_{\delta})^{2}\|S_{\delta}^{-1}S\xi\|^{2}$
$\to(\Delta S)^{2}\|E_{+}\xi\|^{2}\qquad\hbox{as}\ \delta\downarrow 0,$ (3.18)
and this limit is always at most $(\Delta S)^{2}=\Delta\varphi_{\mathrm{ac}}$ (actually, they are equal, because if $E_{+}\neq 1$ then $\Delta S=0$). On the other hand, for every $\delta>0$, the Borel functional calculus gives
$\langle(\log(R^{2}_{\delta}))(0,\xi),(0,\xi)\rangle=2\log(1+\delta)+\underbrace{2\log\Delta S_{\delta}-2\langle(\log S_{\delta})\xi,\xi\rangle}_{=0}.$ (3.19)
Now let $c>\Delta\varphi_{\mathrm{ac}}$, and choose $\delta$ so small that the left-hand side of (3.4) is less than $c$. Having fixed $\delta$, the quantity (3.19) is still strictly positive. Since both squaring and $\log$ are uniformly continuous on the interval $[\delta,\delta^{-1}]$, Kaplansky’s density theorem gives an element $b_{1}$ of $\mathfrak{A}$ which satisfies $\delta\leq b_{1}\leq\delta^{-1}$ and approximates $R_{\delta}$ so well in the strong operator topology that
$\varphi(b_{1}^{2})=\langle\pi(b_{1}^{2})(u,S\xi),(u,S\xi)\rangle<c$ (3.20)
and
$\tau(\log(b_{1}^{2}))=\langle\pi(\log(b_{1}^{2}))(0,\xi),(0,\xi)\rangle>0\qquad\Rightarrow\qquad\Delta b_{1}^{2}>1.$ (3.21)
To finish the proof, choose $b\in A$ that approximates $b_{1}$ sufficiently closely in norm, and let $a:=b^{\ast}b$. ∎
According to Proposition 3.9, the right-hand side of (3.15) does not depend on $\varphi_{\mathrm{sing}}$ at all. The proof explains why. In forming the operator $R_{\delta}$, we use the first summand to suppress the $\lambda$-singular part of $\pi$ as much as we like. Then Kaplansky’s density theorem gives a close enough approximation $a$ to $R_{\delta}$ so that we retain this feature but also pay as little as we like from its Fuglede–Kadison determinant, which is a function of $\lambda(a)$.
Proposition 3.9 expresses $\Delta\varphi_{\mathrm{ac}}$ as an infimum of weak continuous positive linear functionals of $\varphi$. Therefore the general theory of infinite-dimensional convex analysis gives the following standard consequences: see, for example, [135, Theorem I.6.3], or compare [136, Theorem 2.3.4] or [157, Theorem 9.12].
Corollary 3.11.
The quantity $\Delta\varphi_{\mathrm{ac}}$ has the following properties:
  • a.
    if $\varphi,\psi\in\mathfrak{A}^{\ast}_{+}$ and $\varphi\geq\psi$ in the positive definite order, then $\Delta\varphi_{\mathrm{ac}}\geq\Delta\psi_{\mathrm{ac}}$;
  • b.
    if $\varphi,\psi\in\mathfrak{A}^{\ast}_{+}$ and $t\geq 0$, then
    $\Delta(t\varphi)_{\mathrm{ac}}=\Delta(t\varphi_{\mathrm{ac}})=t\Delta\varphi_{\mathrm{ac}}$
    and
    $\Delta(\varphi+\psi)_{\mathrm{ac}}=\Delta(\varphi_{\mathrm{ac}}+\psi_{\mathrm{ac}})\geq\Delta\varphi_{\mathrm{ac}}+\Delta\psi_{\mathrm{ac}};$
  • c.
    it is weak upper semicontinuous as a function of $\varphi\in\mathfrak{A}^{\ast}_{+}$. ∎
Arveson’s result [9, Corollary 4.3.3] is a special case of this corollary.
Since the function $\log$ is monotone and continuous on $(0,\infty)$, conclusions (a) and (c) from Corollary 3.11 carry over to the expression $\log\Delta\varphi_{\mathrm{ac}}$ as well. As written above, conclusion (b) does not hold for this expression, but using also the concavity of $\log$ we can still conclude that this expression is concave:
$\log\Delta(t\varphi+(1-t)\psi)_{\mathrm{ac}}$ $\geq\log(t\Delta\varphi_{\mathrm{ac}}+(1-t)\Delta\psi_{\mathrm{ac}})$
$\geq t\log\Delta\varphi_{\mathrm{ac}}+(1-t)\log\Delta\psi_{\mathrm{ac}}.$ (3.22)
Later in the paper, both Theorems A and C are proved as pairs of inequalities. In both cases, we use the ‘direct’ formula for $\Delta\varphi_{\mathrm{ac}}$ from Definition 3.7 in the proof of the inequality $\geq$”, but find that the variational principle from (3.15) is more convenient for the proof of the inequality $\leq$”.
Further references.
Definition 3.7 is a special case of a construction from quantum statistical mechanics. Consider two faithful normal positive functionals $\omega_{1}$ and $\omega_{2}$ on a von Neumann algebra. Araki’s ‘quantum relative entropy’ ${S(\omega_{1}\mid\omega_{2})}$ is defined by integrating $\log$ against the spectral decomposition of the relative modular operator from $\omega_{1}$ to $\omega_{2}$. His papers [5, 6, 7] studied how it generalizes the quantum relative entropy for states on finite-dimensional matrix algebras. Up to a logarithm and a normalization, this includes Definition 4.12 in case one of those functionals is tracial. This generalizes the fact that the negative log-determinant of a finite-dimensional positive semi-definite matrix $Q$ is equal to the finite-dimensional quantum relative entropy $S(I_{k}\mid Q)$: see, for instance, [16, equation (4.19)]. Similarly, the integral from the right-hand side of Szegő’s limit theorem (1.1) is sometimes written as a classical relative entropy (also called Kullback–Leibler divergence) between two finite measures.
[…]
4 Tuples of vectors and algebras of matrices
4.1 Tuples of vectors and completely positive maps
Let $\mathfrak{A}$ be a separable C*-algebra, $\pi$ be a separable representation of it, and $k$ be a positive integer. Let $v_{1}$, …, $v_{k}\in H_{\pi}$, and regard the tuple $V:=[v_{1},\dots,v_{k}]$ as a linear map from $\mathbf{C}^{\oplus k}$ to $H_{\pi}$. To keep track of how these vectors move together under the action of $\pi$, we can consider the $\mathrm{M}_{k}$-valued function
$\Phi^{\pi}_{V}(a):=V^{\ast}\pi(a)V=[\langle\pi(a)v_{j},v_{i}\rangle]_{i,j=1}^{k}\qquad(a\in\mathfrak{A}).$ (4.1)
Notice that the order of the indices matches the convention for the Gram matrix of a tuple of vectors in (2.2). We sometimes write $\Phi^{\pi}_{v_{1},\dots,v_{k}}$ instead of $\Phi^{\pi}_{V}$.
If $\varphi$ is a continuous linear map from $\mathfrak{A}$ to $\mathrm{M}_{k}$, then a classic theorem of Stinespring asserts that $\varphi$ equals $\Phi^{\pi}_{V}$ for some representation $\pi$ and $k$-tuple $V$ if and only if $\varphi$ is completely positive. See, for example, [139] or [38, Theorem 34.7]. The most substantial part of this result is the construction of $\pi$ and $V$ from an abstract completely positive map $\varphi$. This is an $\mathrm{M}_{k}$-valued generalization of the GNS construction. If we require in addition that $V$ be cyclic for $\pi$, then the resulting pair $(\pi,V)$ is unique up to unitary equivalence. Its underlying representation is called the minimal dilation of $\varphi$ and is denoted by $\pi_{\varphi}$. Many assertions about GNS representations adapt straightforwardly to minimal dilations, and we extend our terminology accordingly. For example, two matrix-valued completely positive maps are disjoint if their minimal dilations are disjoint.
Generalizing our usage in the scalar-valued case, we write that the completely positive map in (4.1) is associated to $\pi$ by $V$, or we call $\Phi^{\pi}_{V}$ the type of $V$ in $\pi$.
In the present paper we do not use the property of complete positivity directly, but we adopt this as the standard term for the class of maps we need to work with. We write $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$ for the space of all continuous linear maps from $\mathfrak{A}$ to $\mathrm{M}_{k}$, and $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ for the subset of all completely positive ones. Complete positivity is defined by a family of closed linear inequalities, so $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ is a weak-closed cone in $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$.
A completely positive map $\varphi:\mathfrak{A}\to\mathrm{M}_{k}$ is normalized if $\overline{\mathrm{tr}}_{k}\varphi(1)=1$, and unital if $\varphi(1)=I_{k}$. A tuple of vectors in $H_{\pi}$ is orthonormal if and only if their type is unital.
Example 4.1.
If $\varphi\in\mathfrak{A}^{\ast}_{+}$ and $Q\in\mathrm{M}_{k+}$, then we may define an element of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ by
$(\varphi\otimes Q)(a):=\varphi(a)\cdot Q\qquad(a\in\mathfrak{A}).$
If $\varphi$ is associated to a representation $\pi$ by a vector $x\in H_{\pi}$, and $Q$ is equal to the Gram matrix $V^{\ast}V$ for some tuple $V=[v_{1},\dots,v_{k}]$ in another Hilbert space $L$, then $\varphi\otimes Q$ is associated to $\pi\otimes I_{L}$ by the tuple $[x\otimes v_{1},\dots,x\otimes v_{k}]$ in $H_{\pi}\otimes L$.
By taking linear combinations of such examples, we can identify the whole of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$ as a vector space with $\mathfrak{A}^{\ast}\otimes\mathrm{M}_{k}$. This is the reason for using tensor-product notation. ∎
The vector space $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$ has a natural topology obtained by applying the weak topology in each matrix entry. Henceforth we simply refer to this as ‘the weak topology’ of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$. As a vector space topology, it also defines a uniform structure on $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$ (see, for instance, [54, Section 8.1] for the basics of uniform structures). We need this uniform structure a few times through the next lemma.
Lemma 4.2.
For any $\pi$ and $k$, the type map
$H_{\pi}^{k}\mapsto\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k}):[v_{1},\dots,v_{k}]\mapsto\Phi^{\pi}_{v_{1},\dots,v_{k}}$
is continuous, and uniformly continuous on any bounded subset of $H_{\pi}^{k}$.
Proof.
These properties are elementary for the inner product map $H_{\pi}\times H_{\pi}\to\mathbf{C}$. They follow for types by arguing pointwise for each $i$, $j\in\{1,\dots,k\}$, and ${a\in\mathfrak{A}}$. ∎
Consider again a representation $\pi$ and a tuple $v_{1}$, …, $v_{k}$ in $H_{\pi}$. If $a=[a_{ij}]$ is an $\ell$-by-$k$ matrix of elements of $\mathfrak{A}$, then we can define a new $\ell$-tuple in $H_{\pi}$ by the formula
$\left[\begin{array}[]{c}y_{1}\\ \vdots\\ y_{\ell}\end{array}\right]:=[\pi(a_{ij})]\cdot\left[\begin{array}[]{c}v_{1}\\ \vdots\\ v_{k}\end{array}\right],$ (4.2)
understood by following the rules of matrix-vector multiplication. For example, if $a=[q_{ij}\cdot 1]$ for some scalar matrix $Q=[q_{ij}]$, then we can identify $[\pi(a_{ij})]$ with $I_{H_{\pi}}\otimes Q$, and (4.2) becomes
$[y_{1},\dots,y_{\ell}]^{\mathrm{T}}:=(I_{H_{\pi}}\otimes Q)[v_{1},\dots,v_{k}]^{\mathrm{T}}.$ (4.3)
If the tuples $v_{1}$, …, $v_{k}$ and $y_{1}$, …, $y_{\ell}$ are related as in (4.2), and $\varphi$ and $\psi$ are their respective types, then $\psi$ may be written using the tuple $v_{1}$, …, $v_{k}$ like this:
[…]
Writing this right-hand side in terms of $\varphi$ itself yields the following.
Lemma 4.3.
In the situation above, we have
[…]
As a result, with $[a_{ij}]$ held fixed, $\psi$ is continuous as a function of $\varphi$ for the weak topologies.
In particular, in the special case of (4.3), we have
$\psi(b):=(Q^{\mathrm{T}})^{\ast}\varphi(b)Q^{\mathrm{T}}\qquad(b\in\mathfrak{A}).$ (4.5)
Sometimes we need to consider all the ways in which two matrix-valued completely positive maps could ‘sit together’ inside a larger one. To describe these, we borrow a term from Furstenberg’s classic work [65] in ergodic theory. Let $k$ and $\ell$ be positive integers, and let
$K\>=\{1,\dots,k\}\qquad\hbox{and}\qquad L:=\{k+1,\dots,k+\ell\}.$
For any $(k+\ell)$-by-$(k+\ell)$ matrix $M$, we write $M[K]$ for its $K$-by-$K$ submatrix, and similarly for $L$.
Definition 4.4.
Let $\varphi:\mathfrak{A}\to\mathrm{M}_{k}$ and $\psi:\mathfrak{A}\to\mathrm{M}_{\ell}$ be completely positive. A joining of them is a completely positive map $\theta:\mathfrak{A}\to\mathrm{M}_{k+\ell}$ such that
$\theta(a)[K]=\varphi(a)\qquad\hbox{and}\qquad\theta(a)[L]=\psi(a)\qquad(a\in\mathfrak{A}).$
In particular, the diagonal joining is defined by
$\mathrm{diag}(\varphi,\psi)(a):=\left[\begin{array}[]{cc}\varphi(a)&0\\ 0&\psi(a)\end{array}\right]\qquad(a\in\mathfrak{A}).$
This terminology is not standard in representation theory, but it is a convenient way to organize various arguments below.
Comparing with ergodic theory, the diagonal joining of two completely positive maps is the analog of the product of two invariant measures.
If $V=[v_{1},\dots,v_{k}]$ and $W=[w_{1},\dots,w_{\ell}]$ are two tuples in a representation $\pi$, then the combined type $\Phi^{\pi}_{[V,W]}$ is a joining of $\Phi^{\pi}_{V}$ and $\Phi^{\pi}_{W}$. These two tuples generate orthogonal subrepresentations of $\pi$ if and only if
$\Phi^{\pi}_{[V,W]}=\mathrm{diag}(\Phi^{\pi}_{V},\Phi^{\pi}_{W}).$
On the other hand, given any joining $\theta$ of $\Phi^{\pi}_{V}$ and $\Phi^{\pi}_{W}$, the minimal dilation $\pi_{\theta}$ contains canonical copies of both tuples. We may therefore characterize disjointness as follows.
Lemma 4.5.
If $\varphi:\mathfrak{A}\to\mathrm{M}_{k}$ and $\psi:\mathfrak{A}\to\mathrm{M}_{\ell}$ are completely positive, then they are disjoint if and only if $\mathrm{diag}(\varphi,\psi)$ is their only joining. ∎
In fact, starting with Furstenberg’s paper [65], the uniqueness of the product joining is taken as the definition of ‘disjointness’ in ergodic theory, where the lack of orthogonal complements makes other definitions impractical or senseless.
4.2 Matrices over algebras
Most of the results about positive functionals in Section 3 have standard generalizations to $\mathrm{M}_{k}$-valued completely positive maps, and the proofs generally run along the same lines. Alternatively, many facts about $\mathrm{M}_{k}$-valued completely positive maps can be reduced directly to the scalar-valued case by regarding tuples of vectors in a representation of $\mathfrak{A}$ as single vectors in a representation of an algebra of matrices over $\mathfrak{A}$.
If $\mathfrak{A}$ is a C-algebra and $k$ is a positive integer, then we write $\mathrm{M}_{k}(\mathfrak{A})$ for the algebra of $k$-by-$k$ matrices with entries from $\mathfrak{A}$. The algebra operations combine those of $\mathfrak{A}$ with the usual rules for matrices, and we define an involution on $\mathrm{M}_{k}(\mathfrak{A})$ by transposing and applying the involution of $\mathfrak{A}$ entry-wise. Then $\mathrm{M}_{k}(\mathfrak{A})$ is a separable C*-algebra in a canonical way. It has a natural identification with $\mathfrak{A}\otimes\mathrm{M}_{k}$, and we pass freely between these two points of view. See, for instance, [38, Section 34].
If $\pi$ is a representation of $\mathfrak{A}$ and $k$ is a positive integer, then we define a representation $\pi^{(k)}$ of $\mathrm{M}_{k}(\mathfrak{A})$ on $H_{\pi}^{\oplus k}$ by setting $\pi^{(k)}([a_{ij}]):=[\pi(a_{ij})]$ and following the rules of matrix-vector multiplication as in (4.2). From another point of view, we can identify $H_{\pi}^{\oplus k}$ with $H_{\pi}\otimes\mathbf{C}^{\oplus k}$, and then $\pi^{(k)}$ is the Kronecker product of $\pi$ with the canonical representation
$\mathrm{canon}_{k}:\mathrm{M}_{k}\to\mathfrak{B}(\mathbf{C}^{\oplus k}).$ (4.6)
The next lemma is standard, but we sketch its proof for completeness.
Lemma 4.6.
The following hold.
  • a.
    Every representation of $\mathrm{M}_{k}(\mathfrak{A})$ is equivalent to $\pi^{(k)}$ for some representation $\pi$ of $\mathfrak{A}$.
  • b.
    Every subrepresentation of $\pi^{(k)}$ is equivalent to $\rho^{(k)}$ for some $\rho\lesssim\pi$.
  • c.
    An element of $\mathfrak{B}(H_{\pi}^{\oplus k})$ commutes with $\pi^{(k)}$ if and only if it has the form $T^{\oplus k}$ for some $T$ that commutes with $\pi$.
  • d.
    We have $\pi^{(k)}\otimes I_{\ell}\simeq(\pi\otimes I_{\ell})^{(k)}$ for every positive integer $k$ and {${}\ell\in\{1,2,\dots,\}\cup\{\infty\}$.
Proof.
Any representation $\kappa$ of $\mathrm{M}_{k}(\mathfrak{A})\cong\mathfrak{A}\otimes\mathrm{M}_{k}$ on a Hilbert space $K$ is generated by the commuting representations $\kappa|\mathrm{M}_{k}$ and $\kappa|\mathfrak{A}$. For $\mathrm{M}_{k}$, every representation is an inflation of $\mathrm{canon}_{k}$: this is a classical result of pure algebra, or a special case of the representation theory of C*-algebras of compact operators [38, Section 16]. We can therefore write $K$ as $H^{\oplus k}$ for some auxiliary Hilbert space $H$ so that $\kappa|\mathrm{M}_{k}$ is identified with $I_{H}\otimes\mathrm{canon}_{k}$. From here, the commutant of $\kappa|\mathrm{M}_{k}$ is identified with $\mathfrak{B}(H)\otimes I_{k}$ [38, Section 50]. Since $\kappa|\mathfrak{A}$ takes values in this commutant, it in turn must have the form $\pi\otimes I_{k}\simeq\pi^{(k)}$ for some representation $\pi$ of $\mathfrak{A}$ on $H$. Finally, if a projection of $K$ commutes with both $\kappa|\mathrm{M}_{k}$ and $\kappa|\mathfrak{A}$, then it must have the form $P\otimes I_{k}$ for some projection $P\in\pi(\mathfrak{A})^{\prime}$ [38, Proposition 50.11]. Putting these ingredients together gives parts (a–c), and part (d) is a direct calculation. ∎
4.3 Functionals on algebras of matrices
Let $\pi$ be a representation of $\mathfrak{A}$ as before. Consider a $k$-tuple $v_{1}$, …, $v_{k}$ in $H_{\pi}$, and form the single vector
$v:=k^{-1/2}[v_{1},\dots,v_{k}]^{\mathrm{T}}\in H_{\pi}^{\oplus k}.$
The factor of $k^{-1/2}$ has the effect that an orthonormal tuple gives rise to a unit vector. The representation $\pi^{(k)}$ of $\mathrm{M}_{k}(\mathfrak{A})$ satisfies
$\langle\pi^{(k)}(a)v,v\rangle=\frac{1}{k}\sum_{ij}\langle\pi(a_{ij})v_{j},v_{i}\rangle=\langle\Phi^{\pi}_{v_{1},\dots,v_{k}},a\rangle\qquad(a=[a_{ij}]\in\mathrm{M}_{k}(\mathfrak{A})),$
where in general we define the pairing of $a\in\mathrm{M}_{k}(\mathfrak{A})$ and $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$ by
$\langle\varphi,a\rangle:=\frac{1}{k}\sum_{ij}\varphi_{ij}(a_{ij}).$ (4.7)
These calculations have the following immediate consequences.
Lemma 4.7.
If $\varphi$ is associated to $\pi$ by the cyclic tuple $v_{1},\dots,v_{k}\in H_{\pi}$, then $\langle\varphi,\cdot\rangle$ is associated to $\pi^{(k)}$ by the cyclic vector $k^{-1/2}[v_{1},\dots,v_{k}]^{\mathrm{T}}$. In particular, $\pi_{\langle\varphi,\cdot\rangle}$ is equivalent to $\pi_{\varphi}^{(k)}$. ∎
Formula (4.7) defines the pairing isomorphism
$\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})\to\mathrm{M}_{k}(\mathfrak{A})^{\ast}.$ (4.8)
With this isomorphism understood, we henceforth regard either space as carrying the dual norm or weak topology.
By Lemma 4.6, any representation of $\mathrm{M}_{k}(\mathfrak{A})$ is of the form $\pi^{(k)}$ up to equivalence. So any positive functional on $\mathrm{M}_{k}(\mathfrak{A})$ is associated to such a representation by a cyclic vector that can be written in the form in Lemma 4.7. It follows that the map in (4.8) restricts to a bijection between the closed cones $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ and $\mathrm{M}_{k}(\mathfrak{A})_{+}^{\ast}$. Because of positivity, the restriction of the dual norm to these cones satisfies
$\|\langle\varphi,\cdot\rangle\|_{\mathrm{M}_{k}(\mathfrak{A})^{\ast}}=\langle\varphi,1\otimes I_{k}\rangle=\overline{\mathrm{tr}}_{k}\varphi(1)\qquad(\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+})$ (4.9)
(see [38, Corollary 7.3]).
The weak topology on $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$ is generally not metrizable, or even first countable. However, because $\mathfrak{A}$ is separable, this topology becomes metrizable upon restriction to any subset of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$ that is bounded in the dual norm [37, Theorem V.5.1]. The next lemma lies between these cases. It must surely be classical, but I have not found a reference for it.
Lemma 4.8.
The restriction of the weak topology to $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ is locally compact and second countable.
Proof.
By considering pairing functionals on $\mathrm{M}_{k}(\mathfrak{A})$ as in (4.7) instead of positive definite maps on $\mathfrak{A}$, we may reduce to the case $k=1$. Having done so, let
$U_{r}:=\{\varphi\in\mathfrak{A}^{\ast}_{+}:\ \varphi(1)<r\}\qquad(r>0).$
Each of these sets is relatively weak-open in $\mathfrak{A}^{\ast}_{+}$. On the other hand, by (4.9), the weak-closure $\overline{U_{r}}$ is equal to the intersection of $\mathfrak{A}^{\ast}_{+}$ with a closed ball of radius $r$ in $\mathfrak{A}^{\ast}$. Therefore $\overline{U_{r}}$ is metrizable and also compact, the latter by the Banach–Alaoglu theorem, and hence it is second countable. So $\mathfrak{A}^{\ast}_{+}$ is covered by the sequence $U_{1}$, $U_{2}$, …of open subsets, each of which is precompact and second countable in the weak topology. ∎
We can transport the ‘Lebesgue’ and ‘Radon–Nikodym’ theorems for positive functionals across the pairing isomorphism (4.8) to obtain their generalizations for completely positive maps, rather than proving these generalizations afresh.
First, we extend the definitions of $\pi$-normal and $\pi$-singular to completely positive maps by simply referring to minimal dilations rather than GNS representations. Then, given a representation $\pi$ of $\mathfrak{A}$ and a map $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, we can apply Proposition 3.3 to the map $\langle\varphi,\cdot\rangle$ and representation $\pi^{(k)}$. Inspecting the resulting decomposition of $\langle\varphi,\cdot\rangle$ in the light of the isomorphism (4.8) and Lemmas 4.6 and 4.7, we obtain the following.
Proposition 4.9.
There is a unique decomposition $\varphi=\varphi_{\mathrm{ac}}+\varphi_{\mathrm{sing}}$ into completely positive summands such that $\varphi_{\mathrm{ac}}$ is $\pi$-normal and $\varphi_{\mathrm{sing}}$ is $\pi$-singular. This decomposition is determined by
$\langle\varphi_{\mathrm{ac}},\cdot\rangle=\langle\varphi,\cdot\rangle_{\mathrm{ac}}\qquad\hbox{and}\qquad\langle\varphi_{\mathrm{sing}},\cdot\rangle=\langle\varphi,\cdot\rangle_{\mathrm{sing}},$
where the right-hand terms refer to the Lebesgue decomposition of $\langle\varphi,\cdot\rangle$ with respect to $\pi$. The summands are linear in $\varphi$.
In particular, $\varphi$ is $\pi$-normal (respectively, $\pi$-singular) if and only if $\langle\varphi,\cdot\rangle$ is $\pi^{(k)}$-normal (respectively, $\pi^{(k)}$-singular). ∎
Now we consider the generalization of Proposition 3.5. Suppose that $\tau$ is a tracial positive functional on $\mathfrak{A}$, and let it be associated to its GNS representation $\lambda$ by the cyclic vector $\xi$. Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, and assume that it is associated to $\lambda^{\oplus\infty}$. By Lemma 4.7 and Lemma 4.6(d), $\langle\varphi,\cdot\rangle$ is associated to $(\lambda^{(k)})^{\oplus\infty}$. However, at this point we must be careful. The representation $\lambda^{(k)}$ has a cyclic $k$-tuple given by
$\xi_{i}=[0,\dots,0,\xi,0,\dots,0]^{\mathrm{T}}\qquad(i=1,2,\dots,k),$ (4.10)
where only the $i^{\mathrm{th}}$ coordinate of $\xi_{i}$ is nonzero. But $\lambda^{(k)}$ need not have a single cyclic tracial vector in general, and we need such a vector in order to apply Proposition 3.5. To fix this, let
$\mathrm{mult}_{k}:\mathrm{M}_{k}\to\mathfrak{B}(\mathrm{M}_{k})$
be the representation of $\mathrm{M}_{k}$ by left multiplication on itself. By writing elements of $\mathrm{M}_{k}$ as $k$-tuples of column vectors, $\mathrm{mult}_{k}$ is equivalent to ${\mathrm{canon}_{k}\otimes I_{k}}$, where $\mathrm{canon}_{k}$ is as in (4.6). Now form the larger tensor product representation
$\lambda\otimes\mathrm{mult}_{k}\simeq(\lambda^{(k)})^{\oplus k}$
of the algebra $\mathrm{M}_{k}(\mathfrak{A})=\mathfrak{A}\otimes\mathrm{M}_{k}$ on $H_{\lambda}\otimes\mathrm{M}_{k}$. This has the single cyclic vector
$\Xi:=\xi\otimes I_{k}=[\xi_{1},\dots,\xi_{k}],$
and the type of $\Xi$ in $(\lambda^{(k)})^{\oplus k}$ is $\tau\otimes\overline{\mathrm{tr}}_{k}$, so it is tracial. We can therefore apply Proposition 3.5 to $\langle\varphi,\cdot\rangle$ and $(\lambda^{(k)})^{\oplus k}$. The result is an operator $T$ affiliated to the algebra
$(\lambda^{(k)})^{\oplus k}\big{(}\mathrm{M}_{k}(\mathfrak{A})\big{)}^{\prime}\cong\mathrm{M}_{k}(\lambda(\mathfrak{A})^{\prime})\cong\lambda^{\oplus k}(\mathfrak{A})^{\prime}$
such that $\langle\varphi,\cdot\rangle$ is associated to $(\lambda^{(k)})^{\oplus k}$ by $T\Xi$. Writing $T\Xi$ as a $k$-tuple of column vectors, and then unpacking the pairing formula (4.7), we arrive at the following.
Proposition 4.10.
If $\varphi$ is $\lambda$-normal, then it is associated to $\lambda^{\oplus k}$ by the tuple of vectors $T\xi_{1}$, …, $T\xi_{k}$ for a unique non-negative square-integrable operator $T$ affiliated to $\lambda^{\oplus k}(\mathfrak{A})^{\prime}$. ∎
Remark 4.11.
We have reached a source of possible confusion. Suppose that $\varphi$ is associated to $\lambda^{\oplus k}$ by the vectors $x_{i}=T\xi_{i}$, $i=1,\dots,k$, as in Proposition 4.10. Suppose further that $a=[a_{ij}]\in\mathrm{M}_{k}(\mathfrak{A})$ and that we define a new tuple $y_{1}$, …, $y_{k}$ by (4.2), so that the type $\psi$ of this new tuple is given by (4.4). In this case, both sides of (4.2) are column vectors of height $k$ with individual entries that are elements of $H_{\lambda}^{\oplus k}$ – that is, they are column $k$-vectors of column $k$-vectors. Correspondingly, $[\lambda^{\oplus k}(a_{ij})]$ is a $k$-by-$k$ matrix of operators with individual entries that are $k$-fold inflations.
So there are two $k$-fold inflations at work here, and it is important to keep their roles separate. The first inflation provides enough ‘room’ in the representation $\lambda^{\oplus k}$ to find a whole $k$-tuple $x_{1}$, …, $x_{k}$ that associates $\varphi$ to $\lambda^{\oplus k}$. The second inflation is needed to express how a $k$-by-$k$ matrix of elements of $\mathfrak{A}$ acts to turn $\varphi$ into $\psi$.
[…]
With Proposition 4.10 in hand, we finally reach the definition that makes full sense of Theorems AB, and C. To explain it, recall that $\lambda^{\oplus k}(\mathfrak{A})^{\prime}$ has a natural identification with $\mathrm{M}_{k}(\lambda(\mathfrak{A})^{\prime})$, which allows us to regard $\widetilde{\tau}\otimes\mathrm{tr}_{k}$ as a normal tracial positive functional on either algebra.
Definition 4.12.
Let $\lambda$ and $\tau$ be as above. If $\varphi$ is $\lambda$-normal, then its Fuglede–Kadison determinant with respect to $\tau$ is
$\Delta_{\tau}\varphi:=(\Delta_{\widetilde{\tau}\otimes\mathrm{tr}_{k}}T)^{2},$ (4.11)
where $T$ is the operator given by Proposition 4.10.
As before, we often write just $\Delta\varphi$ if the correct choice of $\tau$ is clear.
Notice the un-normalized trace on the right-hand side of (4.11). In terms of the normalized trace $\overline{\mathrm{tr}}_{k}$, we have instead
$\Delta_{\tau}\varphi=(\Delta_{\widetilde{\tau}\otimes\overline{\mathrm{tr}}_{k}}T)^{2k}=\big{(}\Delta_{\tau\otimes\overline{\mathrm{tr}}_{k}}(\langle\varphi,\cdot\rangle)\big{)}^{k}.$ (4.12)
Finally, here is the version of Proposition 3.9 for matrix-valued completely positive maps that results from these conventions.
Proposition 4.13.
Let $\lambda$ and $\tau$ be as above, and let $A$ be a dense $\ast$-subalgebra of $\mathrm{M}_{k}(\mathfrak{A})$. Then any $\varphi\in\mathfrak{B}(\mathfrak{A};\mathrm{M}_{k})_{+}$ satisfies
$(\Delta\varphi_{\mathrm{ac}})^{1/k}=\inf\big{\{}\langle\varphi,a\rangle:\ a\in A\ \hbox{positive and invertible and}\ \Delta_{\tau\otimes\mathrm{tr}_{k}}(a)\geq 1\big{\}}.$ (4.13)
The $k^{\mathrm{th}}$ root appears on the left-hand side of (4.13) because we define $\Delta\varphi$ in (4.11) using the un-normalized tracial functional $\tau\otimes\mathrm{tr}_{k}$, but the application of Proposition 3.9 during the proof requires the normalized functional $\tau\otimes\overline{\mathrm{tr}}_{k}$. On the right-hand side of (4.13), the corresponding $k^{\mathrm{th}}$ power makes no difference to the condition that $\Delta_{\tau\otimes\mathrm{tr}_{k}}(a)\geq 1$.
For matrix-valued completely positive maps, Proposition 4.13 shortens the proof of the following.
Corollary 4.14.
Let $\lambda$ and $\tau$ be as above, let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ and $\psi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{\ell})_{+}$, and let $\theta$ be any joining of $\varphi$ and $\psi$ (recall Definition 4.4). Then
$\Delta\theta_{\mathrm{ac}}\leq\Delta\varphi_{\mathrm{ac}}\cdot\Delta\psi_{\mathrm{ac}}.$
Equality holds if $\theta_{\mathrm{ac}}=\mathrm{diag}(\varphi_{\mathrm{ac}},\psi_{\mathrm{ac}})$ (and so, in particular, if $\theta=\mathrm{diag}(\varphi,\psi)$).
Proof.
Step 1. Suppose that $T_{1}$ and $T_{2}$ are non-negative square-integrable operators affiliated to $\lambda^{\oplus k}(\mathfrak{A})^{\prime}$ and $\lambda^{\oplus\ell}(\mathfrak{A})^{\prime}$, respectively. Let $T:=T_{1}\oplus T_{2}$, so this is a non-negative square-integrable operator affiliated to $\lambda^{\oplus(k+\ell)}(\mathfrak{A})^{\prime}$. If $E_{1}$ and $E_{2}$ are the respective spectral resolutions of $T_{1}$ and $T_{2}$, then the spectral resolution of $T$ is given by
$E(A):=E_{1}(A)\oplus E_{2}(A)\qquad\hbox{for any Borel}\ A\subset[0,\infty).$
This gives the calculation of the Fuglede–Kadison determinant of $T$:
$\log\Delta_{\widetilde{\tau}\otimes\mathrm{tr}_{k+\ell}}T$ $=\int_{[0,\infty)}\log t\ (\widetilde{\tau}\otimes\mathrm{tr}_{k+\ell})(E(dt))$
$=\int_{[0,\infty)}\log t\ \big{(}(\widetilde{\tau}\otimes\mathrm{tr}_{k})(E_{1}(dt))+(\widetilde{\tau}\otimes\mathrm{tr}_{\ell})(E_{2}(dt))\big{)}$
$=\log\Delta_{\widetilde{\tau}\otimes\mathrm{tr}_{k}}T_{1}+\log\Delta_{\widetilde{\tau}\otimes\mathrm{tr}_{\ell}}T_{2}.$
Exponentiating, and then applying this identity to the operators that represent $\varphi_{\mathrm{ac}}$ and $\psi_{\mathrm{ac}}$ according to Proposition 4.10, we arrive at the desired equality when $\theta_{\mathrm{ac}}$ equals $\mathrm{diag}(\varphi_{\mathrm{ac}},\psi_{\mathrm{ac}})$.
Step 2. Now consider an arbitrary joining $\theta$. Pick any positive and invertible elements $a_{1}\in\mathrm{M}_{k}(\mathfrak{A})$ and $a_{2}\in\mathrm{M}_{\ell}(\mathfrak{A})$ that satisfy $\Delta_{\tau\otimes\mathrm{tr}_{k}}(a_{1})\geq 1$ and $\Delta_{\tau\otimes\mathrm{tr}_{\ell}}(a_{2})\geq 1$. Pick also a positive real value $r$, and let $a:=\mathrm{diag}(r^{\ell}a_{1},r^{-k}a_{2})$, so this lies in $\mathrm{M}_{k+\ell}(\mathfrak{A})$. By the calculation in Step 1, we have
$\Delta_{\tau\otimes\mathrm{tr}_{k+\ell}}(a)=\Delta_{\tau\otimes\mathrm{tr}_{k}}(r^{\ell}a_{1})\cdot\Delta_{\tau\otimes\mathrm{tr}_{\ell}}(r^{-k}a_{2})=\frac{r^{k\ell}}{r^{k\ell}}\Delta_{\tau\otimes\mathrm{tr}_{k}}(a_{1})\cdot\Delta_{\tau\otimes\mathrm{tr}_{\ell}}(a_{2})\geq 1.$
On the other hand,
$\langle\theta,a\rangle=\frac{kr^{\ell}}{k+\ell}\langle\varphi,a_{1}\rangle+\frac{\ell r^{-k}}{k+\ell}\langle\psi,a_{2}\rangle.$
Taking the infimum over $a_{1}$ and $a_{2}$, this expression becomes
$\frac{kr^{\ell}}{k+\ell}(\Delta\varphi_{\mathrm{ac}})^{1/k}+\frac{\ell r^{-k}}{k+\ell}(\Delta\psi_{\mathrm{ac}})^{1/\ell}$
by Proposition 4.13. Finally, by the inequality of arithmetic and geometric means, the infimum of this last expression over $r$ is equal to $(\Delta\varphi_{\mathrm{ac}}\cdot\Delta\psi_{\mathrm{ac}})^{1/(k+\ell)}$. By Proposition 4.13 again, this gives the desired upper bound on $\Delta\theta_{\mathrm{ac}}$. ∎
Further references.
The textbook [117] is dedicated to completely positive and completely bounded maps between operator algebras, and includes a more complete guide to original references. The basic results that we need are mostly covered in Chapters 4 and 6 of that book. Some more recent uses of completely positive maps in the study of C-algebras, including group C-algebras, can be found in [30, Sections 1.2 and 2.5 and Appendix D].
[…]
5 Group algebras and positive definite functions
This section specializes some results from the previous two sections to the case when $\mathfrak{A}=C^{\ast}\Gamma$ for a countable discrete group $\Gamma$. We write $e$ for the identity in $\Gamma$.
We write $\mathbf{C}[\Gamma]$ for the complex group algebra of $\Gamma$, and regard it concretely as the space of finitely supported functions from $\Gamma$ to $\mathbf{C}$. It is an algebra over $\mathbf{C}$ with convolution as multiplication, and it becomes a $\ast$-algebra under its standard involution: see [62, Section 2.5] or [48, Chapter 13] (which both allow the greater generality of locally compact groups). Given $g\in\Gamma$, we write $\delta_{g}$ for its canonical image in $\mathbf{C}[\Gamma]$, which is the indicator function of the singleton $\{g\}$. The unit of $\mathbf{C}[\Gamma]$ is $\delta_{e}$.
More generally, we may identity $\mathrm{M}_{k}(\mathbf{C}[\Gamma])$ with the vector space $\mathrm{M}_{k}[\Gamma]$ of finitely supported maps from $\Gamma$ to $\mathrm{M}_{k}$. Given $\varphi,\psi:\Gamma\to\mathrm{M}_{k}$, at least one of them finitely supported, we extend the usual definition of convolution by writing
$(\varphi\ast\psi)(g):=\sum_{h,k:\ hk=g}\varphi(h)\psi(k)=\sum_{h}\varphi(h)\psi(h^{-1}g)\qquad(g\in\Gamma).$ (5.1)
The resulting map $\varphi\ast\psi$ takes values in $\mathrm{M}_{k}$. The individual summands here are matrix products, so the order of factors in each summand matters unless $k=1$. If both $\varphi$ and $\psi$ are finitely supported then so is $\varphi\ast\psi$, and then (5.1) defines the structure of $\mathrm{M}_{k}[\Gamma]$ as a group algebra with matrix coefficients. Given three or more functions on $\Gamma$, all but one of them finitely supported, we can multiply them by repeating (5.1). The resulting product is associative, by the usual proof for the scalar-valued case.
[…]
$\|\varphi\|_{\infty}:=\sup_{g}\|\varphi(g)\|\qquad\hbox{and}\qquad\|\psi\|_{1}:=\sum_{g}\|\psi(g)\|.$
Combining (5.1) with the triangle inequality gives the inequalities
$\|\psi\ast\varphi\|_{\infty},\ \|\varphi\ast\psi\|_{\infty}\leq\|\varphi\|_{\infty}\|\psi\|_{1}$ (5.2)
(see [62, Proposition 2.40(a)] for the scalar-valued case).
The group C*-algebra $C^{\ast}\Gamma$ is the maximal C*-completion of $\mathbf{C}[\Gamma]$. Representations of $C^{\ast}\Gamma$ are in one-to-one correspondence with unitary representations of $\Gamma$ itself, and we generally use the same notation for a representation of $C^{\ast}\Gamma$ and for its restriction to $\Gamma$. See [62, Section 7.1] or [48, Section 13.9] for full accounts of this relationship. For each $g\in\Gamma$, we continue to write $\delta_{g}$ for its image in $C^{\ast}\Gamma$. We can identify $\mathrm{M}_{k}(C^{\ast}\Gamma)$ as the corresponding completion of $\mathrm{M}_{k}[\Gamma]$ in the same way.
Now consider a completely positive map $\varphi:C^{\ast}\Gamma\to\mathrm{M}_{k}$. By restriction, it defines an $\mathrm{M}_{k}$-valued map on $\Gamma$ itself:
$g\mapsto\varphi(\delta_{g})\qquad(g\in\Gamma).$
In the sequel we often write simply $\varphi(g)$ instead of $\varphi(\delta_{g})$. If $\varphi=\Phi^{\pi}_{v_{1},\dots,v_{k}}$ for some representation $\pi$ and tuple $v_{1}$, …, $v_{k}$, then (4.1) becomes
$\Phi^{\pi}_{v_{1},\dots,v_{k}}(g)=[\langle\pi(g)v_{j},v_{i}\rangle]\qquad(g\in\Gamma).$ (5.3)
In representation theory, the function on $\Gamma$ given by $\langle\pi(\cdot)v,u\rangle$ is called the $(u,v)$-matrix element of the representation $\pi$.
If $\varphi:\Gamma\to\mathrm{M}_{k}$ is the restriction of a completely positive map as above, then it is an $\mathrm{M}_{k}$-valued positive definite map. This means that it is bounded and satisfies
[…]
for any $a_{1}$, …, $a_{k}\in\mathbf{C}[\Gamma]$. For instance, for the map in (5.3), this holds because the sum in (5.4) is just equal to the squared length of the vector $\sum_{i}\pi(a_{i})v_{i}$. On the other hand, if $\varphi$ is any map on $\Gamma$ satisfying (5.4), then another variant of the GNS construction produces a unitary representation of $\Gamma$ to which $\varphi$ is associated as in (5.3). Using this, $\varphi$ then extends to a completely positive map on the whole of $C^{\ast}\Gamma$.
In case $k=1$, the scalar-valued functions on $\Gamma$ that arise this way are also sometimes called ‘functions of positive type’. This case appears first in work of Gelfand and Raikov [69], and can also be found in [62, Section 3.3] or [48, Section 13.4]. It is an analog for groups of applying the GNS construction to a state on a C*-algebra. Similarly, the generalization to $\mathrm{M}_{k}$-valued positive definite functions on a group is a cousin of Stinespring’s theorem. For matrix-valued positive definite functions on groups, the essential idea of the construction appears in [116], although it is presented there only for Abelian groups. It is fully covered by the generality in [99], as are various further extensions.
So there is a bijection between the set of $\mathrm{M}_{k}$-valued positive definite maps on $\Gamma$ and the space $\mathfrak{B}(C^{\ast}\Gamma,\mathrm{M}_{k})_{+}$. Under this bijection, the weak topology on $\mathfrak{B}(C^{\ast}\Gamma,\mathrm{M}_{k})_{+}$ corresponds to the usual weak topology restricted from $\ell^{\infty}(\Gamma;\mathrm{M}_{k})$, and when restricted further to any uniformly bounded subset it coincides with the topology of pointwise convergence.
Let $\varphi:\Gamma\to\mathrm{M}_{k}$ be positive definite, let $a\in\mathrm{M}_{k}[\Gamma]$, and regard $a$ as an element of $\mathrm{M}_{k}(C^{\ast}\Gamma)$. Having defined convolution in the generality of (5.1), we may use it to express the pairing from (4.7) like this:
$\langle\varphi,a\rangle=\frac{1}{k}\sum_{ij}\sum_{g}a_{ij}(g)\varphi_{ij}(g)=\sum_{g}\overline{\mathrm{tr}}_{k}(\widehat{\varphi}(g^{-1})a(g))=\overline{\mathrm{tr}}_{k}((\widehat{\varphi}\ast a)(e)),$ (5.5)
where we define
$\widehat{\varphi}(g):=\varphi(g^{-1})^{\mathrm{T}}\qquad(g\in\Gamma)$ (5.6)
(that is, we apply inversion in $\Gamma$ and transpose to elements of $\mathrm{M}_{k}$). The conventions in (5.6) loosely generalize the conjugate that appears in the usual definition of the Fourier–Stieltjes transform of a finite measure on a locally compact Abelian group [62, Section 4.2].
Finally, a state on $C^{\ast}\Gamma$ is tracial if and only if it arises from a character of $\Gamma$, meaning a normalized positive definite function $\chi$ that satisfies
$\chi(g^{-1}hg)=\chi(h)\qquad(g,h\in\Gamma).$ (5.7)
For example, the function $1_{\{e\}}$ is the regular character. It is associated to the left regular representation on $\ell^{2}(\Gamma)$ by the function $\delta_{e}$. More generally, if $H$ is a subgroup of $\Gamma$, then the function $1_{H}$ is positive definite. It is associated to the quasi-regular representation of $\Gamma$ on $\ell^{2}(\Gamma/H)$ by the function $\delta_{eH}$. It is a character if and only if $H$ is normal in $G$.
Further references.
Thorough accounts of unitary group representations and their relation to group C*-algebras can be found in [15, Appendices A–C][14, Chapters 1 and 2], or [61, Chapter VI].
The construction of a representation from a matrix-valued positive definite function on a group can be found as [117, Theorem 4.8]. The relationship between positive definite functions on a group and completely positive maps on the group C*-algebra is discussed further at the end of [117, Chapter 4].
[…]
6 A Szegő-like theorem over amenable groups
This section proves Theorem A. The proof is split across Subsections 6.1 and 6.2. Subsection 6.3 offers comparisons with previous work and some other discussion.
Throughout this section, $\Gamma$ is a countable group and $\lambda$ is its left regular representation with the usual cyclic vector $\xi$. This vector associates the regular character to $\lambda$. Unless stated otherwise, we write $\tau$ for the resulting tracial state on $C^{\ast}\Gamma$, and $\widetilde{\tau}$ for the normal tracial state on either $\lambda(\Gamma)^{\prime\prime}$ or $\lambda(\Gamma)^{\prime}$ defined in (3.5). In all three cases the associated Fuglede–Kadison determinant is denoted by $\Delta$.
6.1 Lower bound
In this subsection we prove the inequality $\geq$” in Theorem A. This direction does not require the amenability of $\Gamma$: see Corollary 6.3 below.
The proof of this inequality can be reduced quickly to the case when $\varphi$ is $\lambda$-normal. For that case, the work is done by a more abstract inequality for von Neumann algebras, given in the next proposition.
Proposition 6.1.
Let $\mathfrak{M}$ be a von Neumann subalgebra of $\mathfrak{B}(H)$. Let $V=[x_{1},\dots,x_{k}]$ be an orthonormal tuple in $H$ such that the functional
$\widetilde{\tau}(A):=\frac{1}{k}\sum_{i=1}^{k}\langle Ax_{i},x_{i}\rangle\qquad(A\in\mathfrak{M})$
is tracial, and let $\Delta$ be the Fuglede–Kadison determinant associated to $\widetilde{\tau}$. Finally, let $T$ be a positive semi-definite operator affiliated to $\mathfrak{M}$ whose domain contains $x_{1}$, …, $x_{k}$. Then
$\det((TV)^{\ast}(TV))\geq(\Delta T)^{2k}.$
For example, if $k=1$, then $x_{1}$ is a tracial vector for $\widetilde{\tau}$, and the desired inequality is simply (3.11) for $T^{2}$. Put roughly, we prove the general case by choosing carefully a single vector in the tensor product $H^{\otimes k}$ that reduces the desired inequality to this special case.
More precisely, we first prove Proposition 6.1 when $T$ is bounded and invertible, and then extend to the general case using Lemma 3.6. For any bounded operator $A$ on $H$, let
$A_{i}:=I_{H}\otimes\cdots\otimes I_{H}\otimes A\otimes I_{H}\otimes\cdots\otimes I_{H}\in\mathfrak{B}(H^{\otimes k}),$
where $A$ is in the $i^{\mathrm{th}}$ position. The operators $A_{1}$, …, $A_{k}$ are all still bounded; they commute; and if $A$ is self-adjoint then so is every $A_{i}$. For another bounded operator $T$ on $H$, the tensor product $T^{\otimes k}\in\mathfrak{B}(H^{\otimes k})$ is equal to the product $T_{1}T_{2}\cdots T_{k}$.
In the coming proof, we apply such tensor products to vectors of the form
$x_{1}\wedge\cdots\wedge x_{k}:=\frac{1}{\sqrt{k!}}\sum_{\pi}\mathrm{sgn}(\pi)x_{\pi(1)}\otimes\cdots\otimes x_{\pi(k)}\qquad(x_{1},\dots,x_{k}\in H),$ (6.1)
where the sum runs over all permutations of $\{1,2,\dots,k\}$. This vector is called the alternating product of $x_{1}$, …, $x_{k}$. Since $H$ is a Hilbert space, the closed span of all alternating product vectors can be identified with the alternating product space $H^{\wedge k}$: see, for instance, [138, Section 1.5]. If $x_{1}$, …, $x_{k}$ are orthonormal then $x_{1}\wedge\cdots\wedge x_{k}$ is a unit vector, and more generally alternating products satisfy
$\langle x_{1}\wedge\cdots\wedge x_{k},y_{1}\wedge\cdots\wedge y_{k}\rangle=\det[\langle x_{i},y_{j}\rangle]$ (6.2)
(see, for instance, [138, equation (1.10)]).
Proof of Proposition 6.1.
Step 1. Assume first that $T$ lies in $\mathfrak{M}$ and has a bounded inverse. Since it is also positive definite, we can define the new self-adjoint operator $S:=\log T$ by the functional calculus. Then $S_{i}$ is equal to $\log T_{i}$, because this is a self-adjoint operator whose exponential equals $T_{i}$, and such an operator is unique. By the functional calculus for the commuting self-adjoint operators $T_{1}$, …, $T_{k}$, it follows that
$\log T^{\otimes k}=\log(T_{1}\cdots T_{k})=S_{1}+\cdots+S_{k}.$ (6.3)
Recall that $V=[x_{1},\dots,x_{k}]$, and let $z:={x_{1}\wedge\cdots\wedge x_{k}}$. This is a unit vector in $H^{\otimes k}$ because $x_{1}$, …, $x_{k}$ are orthonormal. Substituting from (6.1), it satisfies
$\langle S_{1}z,z\rangle=\frac{1}{k!}\sum_{\sigma,\pi}\mathrm{sgn}(\sigma\pi)\langle Sx_{\sigma(1)},x_{\pi(1)}\rangle\langle x_{\sigma(2)},x_{\pi(2)}\rangle\cdots\langle x_{\sigma(k)},x_{\pi(k)}\rangle.$
Since $x_{1}$, …, $x_{k}$ are orthogonal, the summand on the right vanishes unless $\sigma(2)=\pi(2)$, …, $\sigma(k)=\pi(k)$, and hence actually $\sigma=\pi$. For these summands, we have $\mathrm{sgn}(\sigma\pi)=1$, and every factor of the form $\langle x_{\sigma(i)},x_{\pi(i)}\rangle$ also equals $1$. As a result, the equation above simplifies to
$\langle S_{1}z,z\rangle=\frac{1}{k!}\sum_{\pi}\langle Sx_{\pi(1)},x_{\pi(1)}\rangle=\frac{1}{k}\sum_{i=1}^{k}\langle Sx_{i},x_{i}\rangle=\widetilde{\tau}(S)=\log\Delta T.$
By symmetry, the analogous formula also holds for $S_{2}$, …, $S_{k}$. Adding these together and substituting from (6.3), we arrive at
$k\cdot\log\Delta T=\langle S_{1}z,z\rangle+\cdots+\langle S_{k}z,z\rangle=\langle\log T^{\otimes k}z,z\rangle.$ (6.4)
On the other hand, if $E$ is the spectral resolution of $T^{\otimes k}$ on $[0,\infty)$, then
$\langle\log T^{\otimes k}z,z\rangle$ $=\int_{[0,\infty)}\log t\ \langle E(dt)z,z\rangle$
$\leq\frac{1}{2}\log\int_{[0,\infty)}t^{2}\ \langle E(dt)z,z\rangle$
$=\frac{1}{2}\log\langle T^{\otimes k}z,T^{\otimes k}z\rangle$
$=\frac{1}{2}\log\det[\langle Tx_{i},Tx_{j}\rangle],$
where we use Jensen’s inequality [126, Theorem 3.3] on the second line and (6.2) on the last line. Combining this calculation with (6.4) completes the proof.
Step 2. Now let $T$ be any positive semi-definite unbounded operator affiliated to $T$ whose domain contains $x_{1}$, …, $x_{k}$. Apply Step 1 to the operators $T_{\delta}$ from Lemma 3.6:
$(\Delta T_{\delta})^{2k}\leq\det[\langle T_{\delta}x_{i},T_{\delta}x_{j}\rangle].$
As $\delta\downarrow 0$, this inequality converges to the desired conclusion, by applying parts (ii) and (i) of Lemma 3.6 to the left- and right-hand sides, respectively. ∎
Remarks 6.2.
  1. 1.
    In the notation above, let $\omega$ be the pure state on $\mathfrak{B}(H^{\otimes k})$ defined by the vector $z$. The calculations above show that $\omega(S_{1})=\widetilde{\tau}(S)$ for any $S\in\mathfrak{M}$, or equivalently that
    $\omega|\mathfrak{M}\otimes I_{H}\otimes\cdots\otimes I_{H}=\widetilde{\tau}\otimes 1\otimes\cdots\otimes 1.$
[…]
  2. 2.
[…]
Now let $\Gamma$ be any countable group and let $\varphi:\Gamma\to\mathrm{M}_{k}$ be positive definite. As in Theorem A, for any finite subset $F$ of $\Gamma$, we consider the $F$-by-$F$ block matrix
$\varphi[F]:=[\varphi(g^{-1}h):\ g,h\in F].$
Corollary 6.3.
If $F$ is finite and nonempty, then
$\det\varphi[F]\geq(\Delta\varphi_{\mathrm{ac}})^{|F|},$
where $\Delta$ is the Fuglede–Kadison determinant associated to the regular character.
Proof.
The Lebesgue decomposition from Proposition 4.9 gives $\varphi[F_{n}]\geq\varphi_{\mathrm{ac}}[F_{n}]$ in the positive definite ordering for every $n$, and so their determinants are ordered the same way [88, Corollary 7.7.4(e)]. We may therefore discard $\varphi_{\mathrm{sing}}$ and assume that $\varphi$ is $\lambda$-normal.
Let $\lambda$ be the left regular representation of $\Gamma$, and let $H:=\ell^{2}(\Gamma)$ and $\xi:=\delta_{e}$. Denote by $\widetilde{\tau}$ the resulting normal tracial state from (3.5) on either $\lambda(\Gamma)^{\prime\prime}$ or $\lambda(\Gamma)^{\prime}$. Let $k$ be a positive integer, let $\xi_{1}$, …, $\xi_{k}$ be the cyclic $k$-tuple for $\lambda^{\oplus k}$ as in (4.10), and let $\mathfrak{M}:=\lambda^{\oplus k}(\Gamma)^{\prime}$, which we identify with $\mathrm{M}_{k}(\lambda(\Gamma)^{\prime})$. Then this $k$-tuple satisfies
$(\widetilde{\tau}\otimes\overline{\mathrm{tr}}_{k})(A)=\frac{1}{k}\sum_{i=1}^{k}\langle A\xi_{i},\xi_{i}\rangle\qquad(A\in\mathfrak{M}).$ (6.5)
Since $\varphi$ is $\lambda$-normal, Proposition 4.10 gives a positive semi-definite operator $T$ affiliated to $\mathfrak{M}$ such that $\xi_{1},\dots,\xi_{k}\in\mathrm{dom}\,T$ and
$\varphi(g)=[\langle\lambda^{\oplus k}(g)T\xi_{j},T\xi_{i}\rangle]_{i,j}\qquad(g\in\Gamma).$ (6.6)
Now define an orthonormal $k|F|$-tuple in $H$ by
$V:=[\lambda^{\oplus k}(g)\xi_{i}:\ i=1,\dots,k,\ g\in F].$
Regarded as a unitary embedding from $\mathbf{C}^{\oplus k|F|}$ into $H$, this tuple satisfies
$\frac{1}{k|F|}\mathrm{tr}(V^{\ast}AV)=\frac{1}{|F|}\sum_{g\in F}\frac{1}{k}\sum_{i=1}^{k}\langle A\lambda^{\oplus k}(g)\xi_{i},\lambda^{\oplus k}(g)\xi_{i}\rangle\qquad(A\in\mathfrak{M}).$
Since $A$ commutes with $\lambda^{\oplus k}$, this simplifies to the expression in (6.5). On the other hand, the definition of $V$ and the fact that $T$ commutes with $\lambda^{\oplus k}$ give
$\varphi[F]$ $=[\varphi(g^{-1}h):\ g,h\in F]$
$=[\langle\lambda^{\oplus k}(h)T\xi_{j},\lambda^{\oplus k}(g)T\xi_{i}\rangle:\ i,j=1,\dots,k,\ g,h\in F]$
$=[\langle T\lambda^{\oplus k}(h)\xi_{j},T\lambda^{\oplus k}(g)\xi_{i}\rangle:\ i,j=i,\dots,k,\ g,h\in F]$
$=(TV)^{\ast}(TV).$
Because of this calculation and (6.5), we can now apply Proposition 6.1 to obtain
$\det\varphi[F]=\det((TV)^{\ast}(TV))\geq(\Delta_{\widetilde{\tau}\otimes\overline{\mathrm{tr}}_{k}}T)^{2k|F|}=(\Delta\varphi_{\mathrm{ac}})^{|F|},$
recalling (4.12) for the final equality. ∎
6.2 Upper bound and completed proof of Theorem A
Our proof of the inequality $\leq$” in Theorem A uses the variational principle from Proposition 4.13. This saves us from having to handle $\varphi_{\mathrm{sing}}$ explicitly: it has already been controlled inside the proof of that principle.
This direction does require the right Følner property of $(F_{n})_{n\geq 1}$. We apply it through the next lemma and its corollary.
Lemma 6.4.
Let $\varphi:\Gamma\to\mathrm{M}_{k}$, let $a:\Gamma\to\mathrm{M}_{k}$ be finitely supported, and let $(F_{n})_{n\geq 1}$ be a right Følner sequence. There are subsets $E_{n}$ of $F_{n}$ such that $|F_{n}\setminus E_{n}|=o(|F_{n}|)$ and
$(\varphi[F_{n}]\cdot a[F_{n}])(g,h)=(\varphi\ast a)[F_{n}](g,h)\qquad\hbox{whenever}\ (g,h)\in F_{n}\times E_{n}.$
Proof.
Let $S:=\{h:\ a(h)\neq 0\}$, so this is finite by assumption, and now let
$E_{n}:=\{h\in F_{n}:\ hS^{-1}\subset F_{n}\}=F_{n}\cap\bigcap_{s\in S}(F_{n}s)\qquad(n=1,2,\dots).$
This satisfies $|F_{n}\setminus E_{n}|=o(|F_{n}|)$ by the right Følner property of $(F_{n})_{n\geq 1}$.
For any $g,h\in F_{n}$, the definition (5.1) gives
$(\varphi\ast a)[F_{n}](g,h)=(\varphi\ast a)(g^{-1}h)=\sum_{k}\varphi(g^{-1}k)a(k^{-1}h).$
In this sum, the factor $a(k^{-1}h)$ is nonzero only if $k^{-1}h\in S$, or equivalently $k\in hS^{-1}$. If $h\in E_{n}$, then $hS^{-1}\subset F_{n}$, so for these $h$ the sum above agrees with
$\sum_{k\in F_{n}}\varphi(g^{-1}k)a(k^{-1}h)=(\varphi[F_{n}]\cdot a[F_{n}])(g,h).$
Corollary 6.5.
Let $(F_{n})_{n\geq 1}$ be a right Følner sequence. If $\varphi:\Gamma\to\mathrm{M}_{k}$ is bounded and $a:\Gamma\to\mathrm{M}_{k}$ is finitely supported, then
$\overline{\mathrm{tr}}_{k|F_{n}|}\big{(}\varphi[F_{n}]\cdot a[F_{n}]\big{)}\to\overline{\mathrm{tr}}_{k}((\varphi\ast a)(e))\qquad\hbox{as}\ n\to\infty.$
Proof.
By Lemma 6.4, there are subsets $E_{n}$ of $F_{n}$ such that $|F_{n}\setminus E_{n}|=o(|F_{n}|)$ and such that the matrix
$\varphi[F_{n}]\cdot a[F_{n}]-(\varphi\ast a)[F_{n}]$
vanishes in all columns indexed by $E_{n}$. On the other hand, all of its entries can be bounded using (5.2). The result is that
$\big{|}\overline{\mathrm{tr}}_{k|F_{n}|}\big{(}\varphi[F_{n}]\cdot a[F_{n}]\big{)}-\overline{\mathrm{tr}}_{k|F_{n}|}\big{(}(\varphi\ast a)[F_{n}])\big{|}\\ \leq 2\cdot\|a\|_{1}\cdot\|\varphi\|_{\infty}\cdot\frac{|F_{n}\setminus E_{n}|}{|F_{n}|}\to 0.$
Finally, every diagonal block of $(\varphi\ast a)[F_{n}]$ is simply equal to $(\varphi\ast a)(e)$, so
$\overline{\mathrm{tr}}_{k|F_{n}|}\big{(}(\varphi\ast a)[F_{n}])=\overline{\mathrm{tr}}_{k}((\varphi\ast a)(e)).$
The other ingredient we need to prove Theorem A is a special case of that theorem which already appears in the literature.
Proposition 6.6.
Let $(F_{n})_{n\geq 1}$ be a right Følner sequence. Suppose that ${a:\Gamma\to\mathrm{M}_{k}}$ is finitely supported, positive definite, and invertible in $\mathrm{M}_{k}(C^{\ast}\Gamma)$. Then
$(\det a[F_{n}])^{1/k|F_{n}|}\to\Delta_{\tau\otimes\overline{\mathrm{tr}}_{k}}a\qquad\hbox{as}\ n\to\infty.$
The earliest reference I know that includes Proposition 6.6 for general amenable groups is [43, Theorem 3.2]; see also the theorem attributed to Lück and Schick that precedes it in that paper. The proof starts with estimates similar to Lemma 6.4, but in which all functions on $\Gamma$ are finitely supported. Using these, one shows by induction on $d$ that
$\overline{\mathrm{tr}}_{k|F_{n}|}\big{(}a[F_{n}]^{d}\big{)}\to\overline{\mathrm{tr}}_{k}(a^{\ast d}(e))\qquad\hbox{as}\ n\to\infty,$
and then by taking linear combinations that
$\overline{\mathrm{tr}}_{k|F_{n}|}(p(a[F_{n}]))\to\overline{\mathrm{tr}}_{k}(p(a)(e))\qquad\hbox{as}\ n\to\infty$
for any polynomial $p$. Finally, the convergence of determinants follows by approximating the function $\log$ uniformly by polynomials on a compact subinterval of $(0,\infty)$.
Proof of Theorem A.
We prove (1.3) as a pair of inequalities.
Step 1. The inequality $\geq$” holds for every $n$ individually by Corollary 6.3.
Step 2. Let $\varphi:\Gamma\to\mathrm{M}_{k}$ be positive definite, and let $a\in\mathrm{M}_{k}(\mathbf{C}[\Gamma])$ be positive definite, invertible in $\mathrm{M}_{k}(C^{\ast}\Gamma)$, and satisfy $\Delta_{\tau\otimes\overline{\mathrm{tr}}_{k}}a\geq 1$. Then Proposition 6.6 gives
$(\det a[F_{n}])^{1/k|F_{n}|}\to\Delta_{\tau\otimes\overline{\mathrm{tr}}_{k}}a\geq 1\qquad\hbox{as}\ n\to\infty.$ (6.7)
Define $\widehat{\varphi}$ from $\varphi$ as in equation (5.6). This satisfies
$\widehat{\varphi}[F_{n}]=[\widehat{\varphi}(g^{-1}h):\ g,h\in F_{n}]=[\varphi(h^{-1}g)^{\mathrm{T}}:\ g,h\in F_{n}]=(\varphi[F_{n}])^{\mathrm{T}},$
where the right-hand side is the transpose of $\varphi[F]$ as a $k|F|$-by-$k|F|$ matrix. Consequently, $\varphi[F_{n}]$ and $\widehat{\varphi}[F_{n}]$ have the same determinant.
Now the multiplicativity of determinants, the determinant-trace inequality (2.1), and the trace property give
$(\det\varphi[F_{n}])^{1/k|F_{n}|}\cdot(\det a[F_{n}])^{1/k|F_{n}|}$ $=\big{(}\det(\sqrt{a[F_{n}]}\cdot\widehat{\varphi}[F_{n}]\cdot\sqrt{a[F_{n}]})\big{)}^{1/k|F_{n}|}$
$\leq\frac{1}{k|F_{n}|}\mathrm{tr}(\sqrt{a[F_{n}]}\cdot\widehat{\varphi}[F_{n}]\cdot\sqrt{a[F_{n}]})$
$=\frac{1}{k|F_{n}|}\mathrm{tr}(\widehat{\varphi}[F_{n}]\cdot a[F_{n}]).$
By Corollary 6.5 and the calculation (5.5), the last normalized trace converges to
$\overline{\mathrm{tr}}_{k}((\widehat{\varphi}\ast a)(e))=\langle\varphi,a\rangle$
as $n\to\infty$. Combining this with (6.7), we have shown that
$\limsup_{n\to\infty}(\det\varphi[F_{n}])^{1/k|F_{n}|}\leq\langle\varphi,a\rangle.$
Taking the infimum over $a$, Proposition 4.13 completes the proof of $\leq$”. ∎
6.3 Further remarks
Comparison with previous work
Our proof of Theorem A has elements in common with various proofs of Theorem S itself. Here are two examples:
  • The proof of Theorem S presented in [136, Section 2.3], which is modeled on Verblunsky’s from [152, 151], uses a classical cousin of the variational principle from Proposition 3.9. However, in [136, Section 2.3] that variational principle is used only to establish weak upper semicontinuity of the right-hand side of (1.1) as a function of $\mu$: see [136, Theorem 2.3.4]. This is a preparation for that proof of Theorem S, but not really an application inside the proof itself. In other respects, that proof of Theorem S seems essentially disjoint from our proof of Theorem A.
  • In case $\mu\ll m$ in Theorem S, the inequality $\geq$” is proved by an application of Jensen’s inequality. This step essentially corresponds to our use of Proposition 6.1 above.
On the other hand, as discussed in the Introduction, most traditional proofs of Theorem S make rather explicit use of a notion of the ‘past’ of a vector under a unitary operator. In the first place, this refers to the subspace $M$ in (1.4). In the setting of Theorem A we must do without this structure. A meaning of ‘past’ reappears when we turn to Theorem B in the next section.
In this respect, we take inspiration from research in ergodic theory that studies entropy without conditioning on the past as in (1.5). This program is discussed more fully in [71]. Regarding Fuglede–Kadison determinants and generalizations of Theorem S itself, some other recent precedents for our work in ergodic theory also have this flavour.
For a polynomial $f$ in $d$ variables, a log-integral much like the right-hand side of (1.1) defines its ‘Mahler measure’. If $f$ has integer coefficients, then it can be used to construct an action of $\mathbf{Z}^{d}$ by automorphisms of a compact Abelian group. This action necessarily preserves the Haar measure of that compact group, and the Kolmogorov–Sinai entropy of this measure-preserving action turns out to equal the logarithm of the Mahler measure of $f$: see [130, Chapter V] and the references given there.
Starting from the observation that a Mahler measure is a Fuglede–Kadison determinant, Deninger generalized this construction to allow $f\in\mathrm{M}_{k}(\mathbf{Z}[\Gamma])$ for any positive integer $k$ and countable group $\Gamma$. He conjectured that the Kolmogorov–Sinai entropy of the resulting system $\boldsymbol{X}_{f}$ should be the Fuglede–Kadison determinant of $f$ whenever $\Gamma$ is amenable and $f$ is non-singular as an operator on $\ell^{2}(\Gamma)^{\oplus k}$. He proved this under various extra hypotheses: see [43, 44, 45] and also the joint work [46].
Deninger proved [43, Theorem 3.2] in order to apply it in this context. This theorem contains our Proposition 6.6, and it actually allows any positive invertible element of $\mathrm{M}_{k}(\lambda(\Gamma)^{\prime\prime})$, although Deninger required additional hypotheses before reaching his main result in ergodic theory. As Deninger describes, several aspects of this proof already appear in earlier works on $L^{2}$-invariants in algebraic topology such as [106, 50, 129].
Li made further progress in [103], where he proved Deninger’s conjectured entropy formula whenever $\Gamma$ is amenable and $f$ is invertible in $\mathrm{M}_{k}(\lambda(\Gamma)^{\prime\prime})$. Among Li’s technical ingredients, [103, Corollary 7.2] strengthens Proposition 6.6 by allowing certain additional perturbations to each of the matrices $\varphi[F_{n}]$.
Finally, Li and Thom proved Deninger’s full conjecture in [104], and generalized it further still to systems constructed by Pontrjagin duality from any $\mathbf{Z}[\Gamma]$-module of type FL. This work needed another strengthening of Proposition 6.6 as an ingredient: see [104, Theorem 1.4], which allows positive elements of $\mathrm{M}_{k}(\lambda(\Gamma)^{\prime\prime})$ that are not necessarily invertible. This is equivalent to the case of Theorem A when $\varphi$ is $\lambda$-normal and the operator $T$ from Proposition 4.10 is bounded.
Apart from these points of contact around [43, Theorem 3.2], the other details of our proof of Theorem A are largely disjoint from those previous works. The most obvious difference is our use of the variational principle from Proposition 4.13. In addition, not all aspects of Szegő’s theorem make an appearance in those previous papers: for example, they all assume that $f$ already lies in $\mathrm{M}_{k}(\lambda(\Gamma)^{\prime\prime})$, so there is no singular part to take care of.
Difficulties of extension beyond amenable groups
It is natural to ask about generalizations of Theorem A to non-amenable groups or other C*-algebras. Various directions suggest themselves, but none of them seems entirely straightforward. We quickly discuss four of these here. For simplicity, we now restrict our attention to scalar-valued positive definite functions and positive functionals.
First, Corollary 6.3 immediately gives the following: in the notation of Theorem A, if $\Gamma$ is any countable group and $\varphi:\Gamma\to\mathbf{C}$ is positive definite, then
$\Delta\varphi_{\mathrm{ac}}\leq\inf\big{\{}(\det\varphi[F])^{1/|F|}:\ F\subset\Gamma\ \hbox{finite and nonempty}\big{\}}.$ (6.8)
Theorem A shows that this is an equality if $\Gamma$ is amenable. I suspect that this implication can be reversed, even if we allow only certain positive definite functions:
Problem 6.7.
Suppose that equality holds in (6.8) whenever $\varphi=\tau(a^{\ast}(\cdot)a)$ for some $a\in\mathbf{C}[\Gamma]$. Must $\Gamma$ be amenable?
If the answer here is positive and $\Gamma$ is not amenable, then it might be worth looking more closely at the ‘gap’ in (6.8), perhaps for special elements of $\mathbf{C}[\Gamma]$ such as the Laplacians corresponding to finite symmetric subsets of $\Gamma$.
Problem 6.8.
For particular choices of $a$ as in Problem 6.7, how does the gap in (6.8) relate to other measures of non-amenability such as the isoperimetric profile of finite subsets of $\Gamma$?
Remark 6.9.
[…]
For our second direction, let us consider the possibility of replacing finite subsets of $\Gamma$ in Theorem A with finite quotients. This moves us to the class of residually finite groups. By allowing quotients with small defects, one could extend this further to sofic groups [76, 159], but we leave these aside here.
If $\Gamma$ is residually finite, then it has a sequence of permutation representations $\sigma_{n}:\Gamma\to\mathrm{Sym}(V_{n})$ on finite sets such that the kernels $\ker\sigma_{n}$ have trivial intersection. In fact, by replacing $\sigma_{n}$ with the diagonal action of many copies of $\sigma_{n}$ if necessary, we may assume further that this sequence is asymptotically free, meaning that
|$|\{v\in V_{n}:\ \sigma_{n}(g)v=v\}|=o(|V_{n}|)\qquad\hbox{as}\ n\to\infty\ \hbox{for every}\ g\in\Gamma\setminus e.$ (6.9)
Let us consider a sequence $(\sigma_{n})_{n\geq 1}$ with this stronger property for convenience.
Let $\pi_{n}$ be the unitary representation on $\mathbf{C}^{\oplus V_{n}}$ induced by $\sigma_{n}$. Then (6.9) implies the convergence
$\overline{\mathrm{tr}}_{|V_{n}|}\pi_{n}(a)\to\tau(a)\qquad\hbox{as}\ n\to\infty,$
first for any $a\in\mathbf{C}[\Gamma]$, and then for any $a\in C^{\ast}\Gamma$ by approximation in norm.
Starting from here, one can reach an analog of Proposition 6.6 before encountering any problems: if $a\in C^{\ast}\Gamma$ is positive and invertible, then
$(\det\pi_{n}(a))^{1/|V_{n}|}\to\Delta a\qquad\hbox{as}\qquad n\to\infty.$ (6.10)
The proof is closely analogous to that of Proposition 6.6. See [46, Theorem 6.1] (which makes the slightly stronger assumption that $a$ is invertible in the Banach algebra $\ell^{1}(\Gamma)$) and [95, Lemma 7.2 and Theorem 7.3].
However, beyond (6.10) for positive and invertible elements $a$, a couple of serious difficulties quickly present themselves.
  1. 1.
    Suppose that $a$ is non-negative but not necessarily invertible. The proof of (6.10) depends on approximating $\log$ by polynomials, and such approximations break down near the origin. On the other hand, our proof of Proposition 6.1 considers alternating products in the fixed Hilbert space of the regular representation, and has no obvious modification for the sequence of spaces $\mathbf{C}^{\oplus V_{n}}$. Absent either of these arguments, we now obtain only the inequality
    $\limsup_{n\to\infty}(\det\pi_{n}(a))^{1/d_{n}}\leq\Delta a\qquad\hbox{as}\ n\to\infty.$
    It turns out that this inequality really can be strict [107, Example 13.69]. This happens if $a$ is non-singular and satisfies $\Delta a>-\infty$, but the finitary matrices $\pi_{n}(a)$ have a few extremely small eigenvalues that drag their determinants down far below $\Delta a$.
    Despite this possibility, the inequality above is still enough for some valuable applications. An example is Elek and Szabó’s proof of Lück’s determinant conjecture for sofic groups in [53]; see also [13].
  2. 2.
    The convergence in (6.10) can be phrased as a fact about the positive definite function $\varphi=\tau(a^{\ast}(\cdot)a)$. The square of the left-hand side of (6.10) is the determinant of a finite-dimensional matrix that approximates $\Delta\varphi$. But for more general positive definite functions, it is not even clear how to choose finite-dimensional matrices that could serve in such an approximation. For a finite subset $F$ of $\Gamma$, we can always form $\varphi[F]$ simply by restricting, but no obvious analog of ‘restriction’ applies to give matrices over the sets $V_{n}$. This is already a problem even if $\varphi$ is associated to the regular representation $\lambda$, but not by a vector of the form $\lambda(a)\xi$ for some $a\in\mathfrak{A}$.
Hayes discusses the first of these difficulties further near the end of [80, Introduction]. That paper concerns an analog of Deninger’s entropy calculation for certain dynamical systems of algebraic origin over sofic groups. For Hayes, the change of focus to sofic entropy meant that he could sidestep the first difficulty above, and the second did not arise because sofic entropy is not defined directly as a limit of normalized Shannon entropy values. We meet a similar situation in our study of almost periodic entropy and Theorem C below. See also the discussion in Subsection 9.7.
In a third direction, we could ask about generalizations of Theorem A to other C*-algebras besides group algebras, but retaining an assumption to play the role of ‘amenability’.
A simple motivating result can be obtained if $\mathfrak{M}$ is a finite von Neumann subalgebra of $\mathfrak{B}(H)$ and $\widetilde{\tau}$ is a faithful normal tracial state on $\mathfrak{M}$. Let $\Delta$ be the Fuglede–Kadison determinant associated to $\widetilde{\tau}$. Let $k_{1}$, $k_{2}$, …be a divergent sequence of positive integers, let $V_{n}$ be an orthonormal $k_{n}$-tuple in $H$ for each $n$, and let $P_{n}:=V_{n}V_{n}^{\ast}$ (the orthogonal projection from $H$ onto $\mathrm{img}\,V_{n}$). Finally, assume the following two properties:
  • $\overline{\mathrm{tr}}_{k_{n}}V_{n}^{\ast}AV_{n}=\widetilde{\tau}(A)$ for every $A\in\mathfrak{M}$, and
  • $\|P_{n}A-AP_{n}\|_{2}=o(\sqrt{k_{n}})$ as $n\to\infty$ for every $A\in\mathfrak{M}$, where $\|\cdot\|_{2}$ is the Hilbert–Schmidt norm.
Then
$(\det V_{n}^{\ast}AV_{n})^{1/k_{n}}\to\Delta A\qquad\hbox{as}\ n\to\infty\ \hbox{for every non-negative}\ A\in\mathfrak{M}.$
Indeed, the inequality $\geq$” holds for every $n$ by another application of Proposition 6.1, and the reverse inequality holds in the limit by a routine variation on the proof of Proposition 6.6.
Sequences $(V_{n})_{n\geq 1}$ satisfying the two properties above have long-standing importance in the study of hyperfiniteness [49, Chapter III.7]. This property is an abstraction of amenability for von Neumann algebras. Indeed, among factors of type II1, the existence of such a sequence identifies uniquely the hyperfinite one: see [36, Theorem 5.1] and also the related discussion in [8, Section 1].
Beyond this simple result for von Neumann algebras, one could try to push Theorem A towards a class of abstract C*-algebras that satisfy some generalization of amenability. Two possibilities are nuclearity [101] or quasidiagonality [29], although both of these properties capture the amenability of a group $\Gamma$ through the reduced C*-algebra $\lambda(C^{\ast}\Gamma)$ rather than $C^{\ast}\Gamma$. However, I do not know precisely what statement one should try to prove, and I suspect that versions of difficulties (1) and (2) above could re-emerge.
Problem 6.10.
Formulate and prove a generalization of Theorem A for positive functionals on an abstract class of C*-algebras that generalizes the class of group C*-algebras of countable amenable groups.
Finally, one could look for a version of Theorem A in a setting that generalizes amenable groups from a more combinatorial point of view. For example, amenability is defined for unimodular random networks in [3, Section 8]. With the right choice of C*-algebra to consider in place of the group C*-algebra, I expect a generalization of Theorem A in this direction could be fairly straightforward.
Further references.
[…]
7 Subdiagonal subalgebras and random orders
This section proves Theorem B. Let $\Gamma$ be a countable group, $\lambda$ be its left regular representation, and $\xi$ be the usual cyclic tracial unit vector for $\lambda$. Let $\varphi$ be an $\mathrm{M}_{k}$-valued positive definite function on $\Gamma$.
In large part, Theorem B is a special case of the Arveson–Labuschagne generalized Szegő theorem for subdiagonal subalgebras. We recall this theorem in Subsection 7.1.
However, some work is required to make contact with that theorem. In the first place, Arveson’s subdiagonal subalgebras are utraweakly closed subalgebras of finite von Neumann algebras, not subalgebras of abstract C*-algebras, so we need to put our objects into this form.
If $\varphi$ is $\lambda$-normal, then we do this by enlarging $\lambda(\Gamma)^{\prime\prime}$ to a von Neumann crossed product algebra that ‘knows’ the law of some translation-invariant random total order. In general, we must first prove directly that $\varphi_{\mathrm{sing}}$ makes no contribution to the left-hand side of (1.7), thereby reducing our work to the $\lambda$-normal case.
The construction of crossed products that we need is reviewed in Subsection 7.2, and its application to a random total order is explained in Subsection 7.3. Subsection 7.4 proves that the singular part of $\varphi$ can be ignored in Theorem B. We finally construct our subdiagonal subalgebra in Subsection 7.5, and then use it to complete the proof of Theorem B in Subsection 7.6. Subsection 7.7 collects some problems for possible future investigation.
Similar uses of crossed product von Neumann algebras have appeared previously in the study of random processes on Cayley graphs or unimodular random networks. See, for example, [66, 3] and the references given there. However, as far as I know, the connection to subdiagonal subalgebras is new.
The study of subdiagonal subalgebras belongs firmly within the abstract theory of von Neumann algebras. As a result, this section cites some more advanced results from the literature of that field than the rest of this paper.
7.1 Background on subdiagonal subalgebras
Within $L^{\infty}(\mathbf{T})$, the Hardy space $H^{\infty}(\mathbf{T})$ contains those functions whose Fourier transform is one-sided. These can be interpreted as the functions that ‘depend only on the past’ in Fourier space. In [9], Arveson defined a subdiagonal subalgebra of a finite von Neumann algebra $\mathfrak{M}$ by axioms abstracted from this example. It can be seen as providing $\mathfrak{M}$ with an abstract notion of ‘past’.
To formulate Arveson’s definition, let $\Phi$ be a faithful normal conditional expectation from $\mathfrak{M}$ onto a von Neumann subalgebra of $\mathfrak{M}$, in the sense studied by Dixmier [47], Umegaki [147, 148, 149] and Tomiyama [144, 145, 146]; see also [9, Definition 2.1.3 and Chapter 6].
The following combines [9, Definitions 2.1.1 and 2.2.2] (and uses slightly different notation). We also restrict our attention to unital subalgerbas for simplicity, although [9] does not.
Definition 7.1.
A unital subalgebra $\mathfrak{H}$ of $\mathfrak{M}$ is subdiagonal (with respect to $\Phi$) if it has the following properties:
  • i.
    $\mathfrak{H}+\mathfrak{H}^{\ast}$ is ultraweakly dense in $\mathfrak{M}$;
  • ii.
    $\Phi(AB)=\Phi(A)\Phi(B)$ whenever $A,B\in\mathfrak{H}$;
  • iii.
    $\Phi[\mathfrak{H}]\subset\mathfrak{H}\cap\mathfrak{H}^{\ast}$.
It is maximal if it is not properly contained in any other subdiagonal subalgebra.
Note that $\mathfrak{H}$ need not be closed, and in non-trivial cases it is not invariant under taking adjoints. By [9, Theorem 2.2.1], any subdiagonal subalgebra with respect to $\Phi$ is contained in exactly one maximal subdiagonal subalgebra.
The intersection $\mathfrak{D}:=\mathfrak{H}\cap\mathfrak{H}^{\ast}$ is called the diagonal of $\mathfrak{H}$. The subset ${\mathfrak{J}:=(\ker\Phi)\cap\mathfrak{H}}$ is an ideal of $\mathfrak{H}$ because of property (ii), and $\mathfrak{H}=\mathfrak{D}+\mathfrak{J}$. We refer to $\mathfrak{J}$ as the off-diagonal ideal of $\mathfrak{H}$.
A maximal subdiagonal subalgebra is necessarily ultraweakly closed, but the reverse implication is not known to hold in general: see [9, Remark 2.2.3]. However, it does hold under hypotheses that cover our present needs.
Theorem 7.2.
Suppose that $\mathfrak{M}$ has a faithful normal tracial state $\widetilde{\tau}$ satisfying $\widetilde{\tau}\circ\Phi=\widetilde{\tau}$. Let $\mathfrak{H}_{0}$ be a subdiagonal subalgebra with respect to $\Phi$, let $\mathfrak{H}$ be the maximal subdiagonal subalgebra containing $\mathfrak{H}_{0}$, and let $\mathfrak{J}_{0}$ and $\mathfrak{J}$ be the off-diagonal ideals of $\mathfrak{H}_{0}$ and $\mathfrak{H}_{0}$, respectively.
If $\mathfrak{H}_{0}$ is ultraweakly closed, then it is equal to $\mathfrak{H}$. In general, $\mathfrak{H}_{0}$ (respectively, $\mathfrak{J}_{0}$) is ultraweakly dense in $\mathfrak{H}$ (respectively, $\mathfrak{J}$). ∎
The first conclusion of Theorem 7.2 is a theorem of Exel [55, Theorem 7], subsequently generalized in [89, 160]. The conclusion about $\mathfrak{H}_{0}$ and $\mathfrak{H}$ follows from this because it implies that the ultraweak closure of $\mathfrak{H}_{0}$ is already maximal. Finally, the conclusion about $\mathfrak{J}_{0}$ and $\mathfrak{J}$ follows because
$A\in\mathfrak{H}_{0}\qquad\Rightarrow\qquad A-\Phi(A)\in\mathfrak{J}_{0}$
and $\Phi$ is normal, hence ultraweakly continuous.
Here is the abstract Szegő theorem for subdiagonal subaglebras:
Theorem 7.3.
In the setting above, assume that $\mathfrak{M}$ has a faithful normal tracial state $\widetilde{\tau}$ such that $\widetilde{\tau}\circ\Phi=\widetilde{\tau}$, and that $\mathfrak{H}$ is maximal. Let $\Delta$ be the Fuglede–Kadison determinant associated to $\widetilde{\tau}$. If $\varphi$ is a normal positive functional on $\mathfrak{M}$, then
$\Delta\varphi$ $=\inf\big{\{}\varphi(|D+A|^{2}):\ D\in\mathfrak{D}_{+},\ A\in\mathfrak{J},\ \Delta D\geq 1\big{\}}$ (7.1)
$=\inf\big{\{}\varphi((1+A^{\ast})B(1+A)):\ B\in\mathfrak{D}_{+},\ A\in\mathfrak{J},\ \Delta B\geq 1\big{\}}.$ (7.2)
In the form of (7.1), this theorem was proposed in [9], and proved there in special cases. The general case of (7.1) was proved much more recently in [100] in the equivalent form of an abstract ‘Jensen formula’. Note that those references define the Fuglede–Kadison determinant of a functional via the variational principle, but this is equivalent to our formulation because of Proposition 3.9.
The expression in (7.2) is a re-arrangement of the right-hand side of (7.1). This depends on the fact that $\mathfrak{H}$ is unital. Indeed, if $D$ is invertible, then we can write
$\varphi(|D+A|^{2})=\varphi((1+D^{-1}A)^{\ast}D^{2}(1+D^{-1}A)),$
and we have $D^{-1}\mathfrak{J}=\mathfrak{J}$ because $\mathfrak{J}$ is an ideal. In general, any $D\in\mathfrak{D}$ satisfying $\Delta D\geq 1$ may be approximated in norm by the invertible elements $D+\varepsilon$, which also have determinant at least $1$. Setting $B:=D^{2}$, this implies that the infima in (7.1) and (7.2) are equal.
Now let $k$ be a positive integer. To allow $\mathrm{M}_{k}$-valued completely positive maps, we can form $k$-by-$k$ matrices in a subdiagonal subalgebra. We begin by noting that the map $\Phi^{(k)}$ is a conditional expectation on $\mathfrak{M}^{(k)}$. Its multiplicative property may be checked directly, and it is still positive because one of Tomiyama’s results implies that conditional expectations are actually completely positive. Now a direct check of the axioms shows that $\mathfrak{H}^{(k)}$ is subdiagonal in $\mathfrak{M}^{(k)}$ with respect to $\Phi^{(k)}$, and that its diagonal and off-diagonal ideal are $\mathfrak{D}^{(k)}$ and $\mathfrak{J}^{(k)}$, respectively.
As a result, if $\varphi:\mathfrak{M}\to\mathrm{M}_{k}$ is normal and completely positive, then we may apply Theorem 7.3 to the pairing functional $\langle\varphi,\cdot\rangle$ to conclude the following.
Theorem 7.4.
Let $\mathfrak{M}$, $\mathfrak{H}$, $\widetilde{\tau}$ and $\Delta$ be as in Theorem 7.3. Then
$(\Delta\varphi)^{1/k}$ $=\inf\big{\{}\langle\varphi,|D+A|^{2}\rangle:\ D\in\mathfrak{D}_{+}^{(k)},\ A\in\mathfrak{J}^{(k)},\ \Delta D\geq 1\big{\}}$
$=\inf\big{\{}\langle\varphi,(1+A)^{\ast}B(1+A)\rangle:\ B\in\mathfrak{D}_{+}^{(k)},\ A\in\mathfrak{J}^{(k)},\ \Delta B\geq 1\big{\}}.$ (7.3)
The $k^{\mathrm{th}}$ root on the left-hand side of (7.4) appears for the same reason as in Proposition 4.13.
Further references.
A modern introduction to conditional expectations between C*-algebras can be found in [30, Section 1.5]. These maps find several applications later in that book.
[…]
7.2 Crossed product actions
Let $\pi$ be a separable representation of $\Gamma$. In addition, let $(\Omega,\mu)$ be a standard probability space with a measure-preserving left $\Gamma$-action, which we denote by
$(g,\omega)\mapsto g\cdot\omega\qquad(g\in\Gamma,\ \omega\in\Omega).$
Let $\kappa$ be the resulting Koopman representation on $L^{2}(\mu)$:
$[\kappa(g)f](\omega):=f(g^{-1}\cdot\omega)\qquad(g\in\Gamma,\ \omega\in\Omega,\ f\in L^{2}(\mu))$
(see [93, Section II.10]).
From these ingredients, form the Hilbert space $L^{2}(\mu;H_{\pi})$ of Bochner square-integrable $H_{\pi}$-valued functions (see [38, Section 52] or [124, Appendix to Section IV.5]). It is the closed span of the functions that have the form
$(a\otimes v)(\omega):=a(\omega)v\qquad(\omega\in\Omega)$
for some $a\in L^{2}(\mu)$ and $v\in H_{\pi}$. By checking that they have the correct inner products, we find that these functions identify $L^{2}(\mu;H_{\pi})$ with the Hilbert-space tensor product $L^{2}(\mu)\otimes H_{\pi}$. We sometimes use these two spaces interchangeably in the rest of this section. For example, for any $S\in\mathfrak{B}(L^{2}(\mu))$ and $T\in\mathfrak{B}(H_{\pi})$, we write $S\otimes T$ for the resulting element of $\mathfrak{B}(L^{2}(\mu;H_{\pi}))$.
The space $L^{2}(\mu;H_{\pi})$ carries the following two actions:
  • I.
    The group $\Gamma$ acts diagonally:
    $[(\kappa\otimes\pi)(g)F](\omega)=\pi(g)F(g^{-1}\cdot\omega)\qquad(g\in\Gamma,\ \omega\in\Omega,\ F\in L^{2}(\mu;H_{\pi})).$
    (See [62, Section 7.3] or [48, Section 13.1] for tensor products of unitary group representations in general.)
  • II.
    The Abelian C*-algebra $L^{\infty}(\mu)$ acts by multiplication:
    $[m_{a}F](\omega):=a(\omega)F(\omega)\qquad(\omega\in\Omega,\ a\in L^{\infty}(\mu),\ F\in L^{2}(\mu;H_{\pi}))$
    (see [38, Theorem 52.7]). In anticipation of its role as a diagonal subalgebra in Subsection 7.6, let us also write $\mathfrak{D}$ for $\{m_{a}:\ a\in L^{\infty}(\mu)\}$.
These actions are related according to
$[m_{a}(\kappa\otimes\pi)(g)F](\omega)$ $=a(\omega)\pi(g)(F(g^{-1}\cdot\omega))$
$=\pi(g)\big{(}a(g\cdot g^{-1}\cdot\omega)F(g^{-1}\cdot\omega)\big{)}$
$=[(\kappa\otimes\pi)(g)m_{a\circ g}F](\omega).$ (7.4)
Let $\mathfrak{M}$ be the von Neumann algebra generated by these two actions:
$\mathfrak{M}:=((\kappa\otimes\pi)(\Gamma)\cup\mathfrak{D})^{\prime\prime}.$ (7.5)
Equivalently, this is the weak operator closure of the linear span of all operators of the form $m_{a}(\kappa\otimes\pi)(g)$ for $g\in\Gamma$ and $a\in L^{\infty}(\mu)$; it is also the ultraweak closure by [49, Theorem I.3.2]. The product of any two operators of this form can be put back into this form using (7.2). This relation identifies that linear span as a representation of a crossed product algebra of $L^{\infty}(\mu)$ and $\Gamma$.
Consider the problem of describing an element of the commutant $\mathfrak{M}^{\prime}$: that is, a bounded operator $T$ on $L^{2}(\mu;H_{\pi})$ that commutes with both $\kappa\otimes\pi$ and $\mathfrak{D}$. Because $T$ commutes with $\mathfrak{D}$, it must have the direct-integral form
[…]
for some strongly measurable map $(T_{\omega}:\ \omega\in\Omega)$ taking values in $\mathfrak{B}(H_{\pi})$ [38, Theorem 52.8]. Then the condition that $T$ commutes with $\kappa\otimes\pi$ turns into this equivariance relation:
$T_{g\cdot\omega}\pi(g)=\pi(g)T_{\omega}\qquad\hbox{for}\ \mu\hbox{-a.e.}\ \omega\qquad(g\in\Gamma).$ (7.6)
For example, this holds if $T_{\omega}$ is constant and equal to a member of $\pi(\Gamma)^{\prime}$, so $\mathfrak{M}^{\prime}$ contains $I_{L^{2}(\mu)}\otimes\pi(\Gamma)^{\prime}$.
Without additional assumptions on $\pi$ or the action on $\Omega$, determining all other solutions to (7.6) can be difficult. As a consequence, it can also be difficult to uncover the structure of $\mathfrak{M}$, or how well this structure ‘remembers’ the equivalence class of the representation $\pi$. However, one case in which a finer analysis is possible is when $\pi$ is the left regular representation (or an inflation of it). That case is the setting in which we eventually find a subdiagonal subalgebra as defined by Arveson. We return to this in Subsection 7.6.
7.3 Random total orders and local past subspaces
Fix $(\Omega,\mu)$ as in the previous subsection. A random total order on this space is a measurable function $(<_{\omega}:\ \omega\in\Omega)$ from $\Omega$ to the compact metrizable space of all relations on $\Gamma$ such that $<_{\omega}$ is almost surely a total order and satisfies
$gh<_{g\cdot\omega}gk\qquad\Leftrightarrow\qquad h<_{\omega}k\qquad(g,h,k\in\Gamma,\ \omega\in\Omega).$ (7.7)
For example, if $<$ is a left-invariant total order on $\Gamma$, then the constant map $<_{\omega}:=<$ is a random total order for any choice of $(\Omega,\mu)$.
An arbitrary countable group $\Gamma$ may not have a left-invariant total order, but it does always have a random total order via the following construction.
Example 7.5.
Let $\Omega:=[0,1]^{\Gamma}$, and let $\mu$ be the product measure under which each coordinate is uniformly distributed. Let $\Gamma$ act by left-translation: if $\omega=(\omega_{h})_{h\in\Gamma}$, then $g\cdot\omega=(\omega_{g^{-1}h})_{h\in\Gamma}$. Now define
$g<_{\omega}h\qquad\Leftrightarrow\qquad\omega_{g}<\omega_{h}.$
This is almost surely a total order because the event $\{\omega_{g}=\omega_{h}\}$ is negligible for any distinct $g$ and $h$. It is called the Bernoulli random order on $\Gamma$, and is discussed more fully in [133, Section 7]. ∎
Now suppose that $<_{\bullet}$ is a random total order on $(\Omega,\mu)$ and that $\pi$ is an arbitrary representation of $\Gamma$. Consider again the actions (I) and (II) on $L^{2}(\mu;H_{\pi})$ from the previous subsection. Fix $x_{1},\dots,x_{k}\in H_{\pi}$, and define
$N_{\omega}:=\overline{\mathrm{span}}\{\pi(g)x_{i}:\ g<_{\omega}e,\ i=1,\dots,k\}\qquad(\omega\in\Omega).$ (7.8)
This is a measurable field of closed subspace of $H_{\pi}$ (see [62, Section 7.4] or [49, Chapter II.1]). These are the measurably-varying subspaces that appear in the statement of Theorem B.
We refer to $N_{\omega}$ as the local past subspace of $x_{1}$, …, $x_{k}$ at $\omega$. The equivariance in (7.7) gives the following identities:
$\pi(g)[N_{g^{-1}\cdot\omega}]$ $=\overline{\mathrm{span}}\{\pi(gh)x_{i}:\ h<_{g^{-1}\cdot\omega}e,\ i=1,\dots,k\}$
$=\overline{\mathrm{span}}\{\pi(gh)x_{i}:\ gh<_{\omega}g,\ i=1,\dots,k\}$
$=\overline{\mathrm{span}}\{\pi(h^{\prime})x_{i}:\ h^{\prime}<_{\omega}g,\ i=1,\dots,k\}.$
For $g\neq e$, this has the following consequences:
$\hbox{if}\qquad g<_{\omega}e\qquad$ $\hbox{then}\qquad\pi(g)[N_{g^{-1}\cdot\omega}+\mathrm{span}\{x_{1},\dots,x_{k}\}]\subset N_{\omega};$ (7.9)
$\hbox{if}\qquad g>_{\omega}e\qquad$ $\hbox{then}\qquad\pi(g)[N_{g^{-1}\cdot\omega}]\supset N_{\omega}+\mathrm{span}\{x_{1},\dots,x_{k}\}.$ (7.10)
7.4 Removing the singular part
In this section we show that the singular part of $\varphi$ can be removed without affecting the conclusion of Theorem B.
Until we complete this removal, I do not see a way to introduce a true subdiagonal subalgebra into our von Neumann algebra $\mathfrak{M}$. Nevertheless, the proofs in this section overlap extensively with how the singular and normal parts of a functional are separated in [17, Section 3], where a subdiagonal subalgebra is already part of the context.
The key to this subsection is the following classical idea. If $\pi$ is any representation of $\Gamma$, then a closed subspace $K$ of $H_{\pi}$ is called wandering under $\pi$ if its images $\pi(g)K$ for $g\in\Gamma$ are pairwise orthogonal. In this case, if $L$ is the $\pi$-invariant subspace of $H_{\pi}$ generated by $K$, then the uniqueness of the GNS construction shows that $\pi^{L}\simeq\lambda\otimes I_{K}$.
Lemma 7.6.
Let $\pi^{M}$ be the $\lambda$-normal part of $\pi$. Let ${(K_{\omega}:\ \omega\in\Omega)}$ be a measurable field of closed subspace of $H_{\pi}$, and let
[…]
so $K$ is a closed subspace of $L^{2}(\mu;H_{\pi})$. If $K$ is wandering under $\kappa\otimes\pi$, then $K$ is contained in $L^{2}(\mu;M)$.
Proof.
The Hilbert space $L^{2}(\mu)$ is the orthogonal sum of $\mathbf{C}1_{\Omega}$ and $L^{2}_{0}(\mu)$, where the latter is the subspace of functions whose integral is zero. The representation $\kappa$ respects these two subspaces, and is trivial on the first one. Let $\kappa=1\oplus\kappa_{0}$ be the resulting decomposition of $\kappa$. Forming the tensor product with $\pi$, this becomes
$\kappa\otimes\pi\simeq(1\otimes\pi)\ \oplus\ (\kappa_{0}\otimes\pi).$
By the wandering assumption, the subrepresentation $(\kappa\otimes\pi)^{K}$ is equivalent to an inflation of $\lambda$. Is it therefore disjoint from $\pi^{M^{\perp}}$, and hence orthogonal to the subspace $1\otimes M^{\perp}$ in $L^{2}(\mu;H_{\pi})$. In view of the direct-integral form of $K$, this is possible only if $K_{\omega}\subset M$ for almost every $\omega$. ∎
The next lemma helps us to argue about a wandering subspace once we have found it.
Lemma 7.7.
Let $H$ be a Hilbert space, $M$ be a closed subspace of it, $S\subset H$, and $x\in H$. Let $P$ be the orthogonal projection of $H$ onto $M$, and let
$L:=\overline{\mathrm{span}}\,S\qquad\hbox{and}\qquad L_{0}:=\overline{\mathrm{span}\,}P[S]$
(beware that $L_{0}$ need not be contained in $L$). Finally, let $R$ and $R_{0}$ be the orthogonal projections from $H$ onto $L$ and $L_{0}$, respectively.
If $R^{\perp}x\in M$, then $R^{\perp}x=R_{0}^{\perp}Px$.
Proof.
First observe that $P[S]$ has dense span in $P[L]$, and hence $P[L]$ is a dense (but not necessarily closed) subspace of $L_{0}$.
Let $y:=R^{\perp}x$. To show that $y=R_{0}^{\perp}Px$, we show separately that $y$ is orthogonal to $L_{0}$ and that $Px-y$ lies in $L_{0}$.
First, $y$ is orthogonal to $L$ by construction, and $y=Py$ by assumption, so any $z\in L$ satisfies
$\langle y,Pz\rangle=\langle Py,z\rangle=\langle y,z\rangle=0.$
That is, $y$ is orthogonal to $P[L]$, and so it is orthogonal to $\overline{P[L]}=L_{0}$.
Secondly, $Px-y=Px-Py=PRx$, so this lies in $P[L]\subset L_{0}$. ∎
To explain the consequences of these results for Theorem B, consider now a general representation $\pi$ and a positive definite function $\varphi:\Gamma\to\mathrm{M}_{k}$ that is associated to $\pi$ by the tuple $x_{1}$, …, $x_{k}$. Let $\pi^{M}$ be the $\lambda$-normal part of $\pi$. Let $P$ be the orthogonal projection of $H$ onto $M$, so $\varphi_{\mathrm{ac}}$ is associated to $\pi^{M}$ by the tuple $Px_{1}$, …, $Px_{k}$. Finally, for each $\omega$, let $R_{\omega}$ (respectively, $R_{0\omega}$) be the orthogonal projection from $H$ onto the local past subspace $N_{\omega}$ of $x_{1}$, …, $x_{k}$ (respectively, $N_{0\omega}$ of $Px_{1}$, …, $Px_{k}$) at $\omega$.
Corollary 7.8.
In the setting above, the left-hand side of (1.7) is equal to
$\int\log\det[\langle R^{\perp}_{0\omega}Px_{j},R^{\perp}_{0\omega}Px_{i}\rangle]\ d\mu(\omega).$
In particular, the left-hand side of (1.7) is the same for $\varphi$ and for $\varphi_{\mathrm{ac}}$.
Proof.
Let $K_{\omega}$ be the span of $\{R_{\omega}^{\perp}x_{i}:\ i=1,2,\dots,k\}$ for each $\omega\in\Omega$, and define $K$ as in (7.11). The main step is to show that $K$ is wandering under $\pi\otimes\kappa$.
Consider two elements of $K$ that have the forms
$U(\omega)=a(\omega)R_{\omega}^{\perp}x_{i}\qquad\hbox{and}\qquad V(\omega)=b(\omega)R_{\omega}^{\perp}x_{j}\qquad(\omega\in\Omega)$
for some $a,b\in L^{2}(\mu)$ and indices $i,j$. Functions of this form have dense span in $K$, so it suffices to show that $U$ is orthogonal to $(\pi\otimes\kappa)(g)V$ for every $g\in\Gamma\setminus e$. More fully, we have
$[(\pi\otimes\kappa)(g)V](\omega)$ $=b(g^{-1}\cdot\omega)\pi(g)(R^{\perp}_{g^{-1}\cdot\omega}x_{j}),$
and so we need to show that
[…]
For each $\omega$, there are two possibilities:
  • If $g<_{\omega}e$, then relation (7.9) gives
    $\pi(g)R^{\perp}_{g^{-1}\cdot\omega}x_{j}\in\pi(g)[\mathbf{C}x_{j}+N_{g^{-1}\cdot\omega}]\subset N_{\omega}\ \perp\ R_{\omega}^{\perp}x_{i}.$
  • If $g>_{\omega}e$, then relation (7.10) gives
    $\pi(g)R^{\perp}_{g^{-1}\cdot\omega}x_{j}\ \perp\ \pi(g)[N_{g^{-1}\cdot\omega}]\supset N_{\omega}+\mathbf{C}x_{i}\ni R^{\perp}_{\omega}x_{i}.$
Partitioning $\Omega$ into these two events, we see that the integrand in (7.12) is zero in either event, and so that identity is proved.
Having proved this wandering property, Lemma 7.6 shows that $K$ is contained in $L^{2}(\mu;M)$. In particular, the vectors $R_{\omega}^{\perp}x_{i}$ for $i=1,2,\dots,k$ all lie in $M$ for almost every $\omega$. In this event, Lemma 7.7 shows that $R_{\omega}^{\perp}x_{i}$ is equal to $R^{\perp}_{0\omega}Px_{i}$. Now substitute this equality into the left-hand side of (1.7). ∎
7.5 Construction of a subdiagonal subalgebra
We have almost reached the proof of Theorem B itself, but we still need to find a subdiagonal subalgebra of $\mathfrak{M}$. This requires the extra assumption that $\pi$ be the left regular representation $\lambda$. We explain the consequences of this assumption next.
Consider again the actions (I) and (II) and the von Neumann algebra $\mathfrak{M}$ defined in (7.5). Since $H_{\lambda}=\ell^{2}(\Gamma)$, we can regard $L^{2}(\mu;H_{\lambda})$ as a direct sum of copies of $L^{2}(\mu)$ indexed by $\Gamma$, with action (I) being given by $\kappa$ on $L^{2}(\mu)$ together with permuting the summands from the left. This identifies $\mathfrak{M}$ as a crossed product von Neumann algebra in the sense introduced by Murray and von Neumann; see [49, Sections I.9.2–3].
Let $\xi:=\delta_{e}$, so the associated positive definite function is the regular character $\chi$ on $\Gamma$. In this subsection we also write $\chi$ for the tracial state defined by this character on $C^{\ast}\Gamma$.
The vector $1\otimes\xi$ is still cyclic and tracial for $\mathfrak{M}$, and the resulting tracial state $\widetilde{\tau}$ on $\mathfrak{M}$ is given by linear extension of the formula
[…]
this is essentially [49, Proposition I.9.1]. So in this case $\mathfrak{M}$ is a finite von Neumann algebra, and one can determine $\mathfrak{M}^{\prime}$ via the properties of the resulting canonical involution from Lemma 3.4. The algebra $\mathfrak{M}$ also has a conditional expectation operator $\Phi$ onto $\mathfrak{D}$ given by linear extension of this formula:
$\Phi(m_{a}(\kappa\otimes\pi)(g)):=\left\{\begin{array}[]{ll}m_{a}&\quad g=e\\ 0&\quad g\neq e.\end{array}\right.$ (7.14)
In the notation of [49, Section I.9.2], this is the map “$T\mapsto T_{e}$” for $T\in\mathfrak{M}$; it also falls within [9, Example 6.1.3(2)]. Comparing with (7.13) shows that $\widetilde{\tau}\circ\Phi=\widetilde{\tau}$. For our work in this section, the maps $\widetilde{\tau}$ and $\Phi$ are the key extra features of $\mathfrak{M}$ that appear when $\pi=\lambda$.
Now suppose we have a random total order $(<_{\omega}:\ \omega\in\Omega)$. We use it to distinguish certain special operators in $\mathfrak{M}$ as follows. First, we call the operator $m_{a}(\kappa\otimes\pi)(g)$ pre-subdiagonal if
$a(\omega)=0\qquad\hbox{whenever}\qquad g>_{\omega}e.$ (7.15)
We call it strictly pre-subdiagonal if also $g\neq e$. Now let $\mathfrak{H}_{0}$ be the linear span of all pre-subdiagonal operators and $\mathfrak{J}_{0}$ be the linear span of all strictly pre-subdiagonal operators, in both cases without taking closures. Note that (7.15) allows any function $a$ in case $g=e$, so
$\mathfrak{H}_{0}=\mathfrak{D}+\mathfrak{J}_{0}.$
Proposition 7.9.
If $\pi=\lambda$, then $\mathfrak{H}_{0}$ is a unital subdiagonal subalgebra of $\mathfrak{M}$ with respect to $\Phi$, its diagonal is equal to $\mathfrak{D}$, and its off-diagonal ideal is equal to $\mathfrak{J}_{0}$.
The proof generalizes [9, Subsection 3.2] ‘relative to $\Omega$’.
Proof.
The collection $\mathfrak{H}_{0}$ is a linear space by construction, and it contains $\mathfrak{D}$ so it is unital. To show that $\mathfrak{H}_{0}$ is an algebra, consider two pre-subdiagonal operators ${m_{a}(\kappa\otimes\pi)(g)}$ and ${m_{b}(\kappa\otimes\pi)(h)}$. We must show that their product is still pre-subdiagonal. The relation (7.2) gives
$m_{a}(\kappa\otimes\pi)(g)m_{b}(\kappa\otimes\pi)(h)=m_{a\cdot(b\circ g^{-1})}(\kappa\otimes\pi)(gh).$ (7.16)
We verify that the function $a\cdot(b\circ g^{-1})$ satisfies (7.15) for the group element $gh$. Assume this function is nonzero at some point $\omega$. Since both operators satisfy (7.15), we must have $g\leq_{\omega}e$ and $h\leq_{g^{-1}\cdot\omega}e$. By (7.7), the second of these implies that $gh\leq_{\omega}g$, and now the transitivity of $\leq_{\omega}$ implies that $gh\leq_{\omega}e$. So the right-hand side of (7.16) is still pre-subdiagonal.
We now verify the three axioms from Definition 7.1.
  • i.
    The relation (7.2) gives
    $(m_{a}(\kappa\otimes\pi)(g))^{\ast}=(\kappa\otimes\pi)(g^{-1})m_{\overline{a}}=m_{\overline{a\circ g}}(\kappa\otimes\pi)(g^{-1}).$
    Combined with (7.15) and (7.7), this shows that $\mathfrak{H}_{0}^{\ast}$ is the linear span of all operators of the form $m_{a}(\kappa\otimes\pi)(g)$ that satisfy the additional condition
    $a(\omega)=0\qquad\hbox{whenever}\qquad g<_{\omega}e.$
    Any function $a$ on $\Omega$ may be decomposed almost everywhere according to
    $a(\omega)=a(\omega)1_{\{g\geq_{\omega}e\}}+a(\omega)1_{\{g<_{\omega}e\}},$
    so $\mathfrak{H}_{0}+\mathfrak{H}_{0}^{\ast}$ contains all finite sums of operators of the form $m_{a}(\kappa\otimes\pi)(g)$, and is therefore ultraweakly dense in $\mathfrak{M}$.
  • ii.
    Let $m_{a}(\kappa\otimes\pi)(g)$ and $m_{b}(\kappa\otimes\pi)(h)$ be pre-subdiagonal. Using (7.14) and (7.16) again, we have
    $\Phi\big{(}m_{a}(\kappa\otimes\pi)(g)m_{b}(\kappa\otimes\pi)(h)\big{)}=\left\{\begin{array}[]{ll}m_{a\cdot(b\circ g^{-1})}&\quad h=g^{-1}\\ 0&\quad g\neq h^{-1}.\end{array}\right.$ (7.17)
    We must show that this equals the product
    $\Phi(m_{a}(\kappa\otimes\pi)(g))\Phi(m_{b}(\kappa\otimes\pi)(h))=\left\{\begin{array}[]{ll}m_{a\cdot b}&\quad h=g=e\\ 0&\quad\hbox{otherwise}.\end{array}\right.$
    This follows directly if either $g\neq h^{-1}$ or $g=h=e$, so it remains to show that the first right-hand possibility in (7.17) vanishes also in the case $g\neq e$. This holds by our assumptions of pre-subdiagonality:
    $\left\{\begin{array}[]{c}g>_{\omega}e\\ e>_{\omega}g\ \Rightarrow\ h=g^{-1}>_{g^{-1}\cdot\omega}e\end{array}\right\}\qquad\Rightarrow\qquad\left\{\begin{array}[]{c}a(\omega)=0\\ b(g^{-1}\cdot\omega)=0\end{array}\right\}.$
    By linearity, this verifies condition (ii).
  • iii.
    This holds because $\mathfrak{H}\cap\mathfrak{H}^{\ast}=\mathfrak{D}$, which is already the entire image of $\Phi$.
Finally, the assertions about $\mathfrak{D}$ and $\mathfrak{J}_{0}$ follow from the definitions and the formula for $\Phi$. ∎
Now let $\mathfrak{H}$ be the unique maximal subdiagonal subalgebra that contains $\mathfrak{H}_{0}$, as provided by [9, Theorem 2.2.1]. Its diagonal must contain the diagonal of $\mathfrak{H}_{0}$ and must be contained in the image of $\Phi$, so the diagonal of $\mathfrak{H}$ is still equal to $\mathfrak{D}$. Let $\mathfrak{J}$ be the off-diagonal ideal of $\mathfrak{H}$. Because our setting includes the faithful tracial state $\widetilde{\tau}$, Theorem 7.2 tells us that $\mathfrak{H}_{0}$ (respectively, $\mathfrak{J}_{0}$) is ultraweakly (and hence also weak operator) dense in $\mathfrak{H}$ (respectively, $\mathfrak{J}$).
7.6 Completed proof of Theorem B
To prove Theorem B, it mostly remains to apply Theorem 7.4 to $\mathfrak{H}$ and interpret the result. However, we need some final preparations to allow for the possibility that $k>1$. In that case, we need the constructions above starting from $\lambda^{\oplus k}$ rather than $\lambda$. In the notation of the previous subsection, they then lead us to $\mathfrak{M}\otimes I_{k}$, $\mathfrak{H}\otimes I_{k}$, $\mathfrak{J}\otimes I_{k}$ and $\mathfrak{D}\otimes I_{k}$. These are all images under the inflation isomorphism $\mathfrak{M}\to\mathfrak{M}\otimes I_{k}$, so we can carry $\widetilde{\tau}$ and $\Phi$ over directly to a tracial functional $\widetilde{\tau}_{k}$ and a conditional expectation $\Phi_{k}$ on $\mathfrak{M}\otimes I_{k}$. All the other von Neumann algebra properties proved so far are preserved, including the conclusions of Proposition 7.9.
This inflation step should not be confused with the role of $k$-by-$k$ matrices in the next lemma; compare with Remark 4.11.
Lemma 7.10.
Let $\varphi:\Gamma\to\mathrm{M}_{k}$ be positive definite, and suppose it is associated to $\lambda^{\oplus k}$ by the vectors $x_{1}$, …, $x_{k}$. Define a normal completely positive map $\psi:\mathfrak{M}\otimes I_{k}\to\mathrm{M}_{k}$ by the pairing formula
$\langle\psi,A\rangle:=\left\langle A\left[\begin{array}[]{c}1\otimes x_{1}\\ \vdots\\ 1\otimes x_{k}\end{array}\right],\left[\begin{array}[]{c}1\otimes x_{1}\\ \vdots\\ 1\otimes x_{k}\end{array}\right]\right\rangle\qquad(A\in\mathrm{M}_{k}(\mathfrak{M}\otimes I_{k})).$
Then
$\Delta_{\widetilde{\tau}_{k}}\psi=\Delta_{\chi}\varphi.$
Proof.
According to Proposition 4.10, there is a unique non-negative square-integrable operator $T$ affiliated to $\lambda^{\oplus k}(\Gamma)^{\prime}\cong\mathrm{M}_{k}(\lambda(\Gamma)^{\prime})$ such that $x_{i}=T\xi_{i}$ for each $i$. Then the tensor product $I_{L^{2}(\mu)}\otimes T$ is a non-negative operator affiliated to $(\mathfrak{M}\otimes I_{k})^{\prime}$, and $1\otimes x_{i}=(I_{L^{2}(\mu)}\otimes T)(1\otimes\xi_{i})$ for each $i$. Finally, if $E$ is the spectral resolution of $T$ on $[0,\infty)$, then $I_{L^{2}(\mu)}\otimes E$ is the spectral resolution of $I_{L^{2}(\mu)}\otimes T$. Now two appeals to Definition 4.12 give
$\log\Delta_{\widetilde{\tau}_{k}}\psi$ $=2\log\Delta_{\widetilde{\tau}_{k}\otimes\mathrm{tr}_{k}}(I_{L^{2}(\mu)}\otimes T)$
$=2\int_{[0,\infty)}\log s\ (\widetilde{\tau}_{k}\otimes\mathrm{tr}_{k})(I_{L^{2}(\mu)}\otimes E(ds))$
$=2\int_{[0,\infty)}\log s\ (\chi\otimes\mathrm{tr}_{k})E(ds)$
$=\log\Delta_{\chi}\varphi.$
Now let $\pi:=\lambda^{\oplus k}$, consider a $k$-tuple $x_{1}$, …, $x_{k}\in H_{\pi}$, and consider the local past subspaces $N_{\omega}$ of this tuple for $\omega\in\Omega$. Define
[…]
understanding this direct integral as a closed subspace of $L^{2}(\mu;H_{\pi}^{\oplus k})$ (see [62, Section 7.4] or [49, Chapter II.1]).
Lemma 7.11.
The set
$N_{0}:=\left\{A\left[\begin{array}[]{c}1\otimes x_{1}\\ \vdots\\ 1\otimes x_{k}\end{array}\right]:\ A\in\mathrm{M}_{k}(\mathfrak{J}_{0}\otimes I_{k})\right\}$
is a dense subspace of $N$, and the same conclusion holds if $\mathfrak{J}_{0}$ is replaced by $\mathfrak{J}$.
Proof.
Step 1. First suppose that $A=[m_{a_{ij}}(\kappa\otimes\pi)(g)]$, where $[a_{ij}]$ is a $k$-by-$k$ matrix of elements of $L^{\infty}(\mu)$ and every entry is strictly pre-subdiagonal. Set
$[y_{1},\dots y_{k}]^{\mathrm{T}}:=A[1\otimes x_{1},\dots,1\otimes x_{k}]^{\mathrm{T}}.$
Written out in full, this reads
$y_{i}(\omega)=\sum_{j}a_{ij}(\omega)\pi(g)x_{j}\qquad(\omega\in\Omega,\ i=1,2,\dots,k).$
Since every entry of $A$ is strictly pre-subdiagonal, the sum above is zero unless $g<_{\omega}e$. So, for each fixed $\omega$, this vector $y_{i}(\omega)$ is either zero or a linear combination of the shifts of $x_{1}$, …, $x_{k}$ that appear in (7.8). This shows that $y_{i}(\omega)\in N_{\omega}$ for almost every $\omega$, and hence that $N_{0}\subset N$ (when defined using $\mathfrak{J}_{0}$).
Step 2. On the other hand, the allowed choices for $g$ and $a_{i1}$, …, $a_{ik}$ above give vectors whose span is dense in $N_{\omega}$. Since we may make separate choices for each $i$, the tuples $[w_{1}(\omega),\dots,w_{k}(\omega)]^{\mathrm{T}}$ that we obtain from all possible $A\in\mathrm{M}_{k}(\mathfrak{J}_{0}\otimes I_{k})$ are dense in the whole of $N_{\omega}^{\oplus k}$. By taking measurable selections in $\omega$, this proves that $N_{0}$ is dense in $N$.
Step 3. Finally, if we replace $\mathfrak{J}_{0}$ by $\mathfrak{J}$, then this can only enlarge the set $N_{0}$, so we need only show that it is still contained in $N$. This follows from Theorem 7.2, because $N$ is a closed subspace of $L^{2}(\mu;H_{\pi})$, hence also weakly closed, and the map
$A\mapsto A[1\otimes x_{1},\dots,1\otimes x_{k}]^{\mathrm{T}}$
is continuous from the weak operator topology to the weak topology. ∎
Proof of Theorem B.
By Corollary 7.8, the left-hand side of (1.7) is unchanged if we replace $\varphi$ with $\varphi_{\mathrm{ac}}$, so we may assume that $\varphi$ is $\lambda$-normal. Having done so, by Proposition 4.10 it is associated to the representation $\pi:=\lambda^{\oplus k}$ by a tuple $x_{1}$, …, $x_{k}$ in $H_{\pi}=\ell^{2}(\Gamma)^{\oplus k}$. Let $N_{\omega}$ be the local past subspace of $x_{1}$, …, $x_{k}$ at $\omega$.
Use $\varphi$ to construct a completely positive map $\psi:\mathfrak{M}\otimes I_{k}\to\mathrm{M}_{k}$ as in Lemma 7.10. By that lemma, the right-hand side of (1.7) is equal to $\log\Delta_{\widetilde{\tau}_{k}}\psi$. On the other hand, because of Proposition 7.9 and its version for the inflation $\mathfrak{H}\otimes I_{k}$, Theorem 7.4 tells us that $\Delta_{\widetilde{\tau}_{k}}\psi$ is given by the second infimum in (7.4).
It remains to prove that this infimum agrees with the left-hand side of (1.7). We do this by first evaluating the infimum over $B$ with $A$ fixed, and then evaluating the infimum over $A$.
First infimum. Let $B\in\mathrm{M}_{k}(\mathfrak{D}\otimes I_{k})$ and $A\in\mathrm{M}_{k}(\mathfrak{J}\otimes I_{k})$, and write $B$ as $[m_{b_{ij}}\otimes I_{k}]$. Define a new $k$-tuple $w_{1}$, …, $w_{k}$ by
$[w_{1},\dots,w_{k}]^{\mathrm{T}}=(I+A)[1\otimes x_{1},\dots,1\otimes x_{k}]^{\mathrm{T}}$
(so each $w_{i}$ individually is an element of $L^{2}(\mu;\ell^{2}(\Gamma)^{\oplus k})$). Then we can write the expression inside the second infimum in (7.4) like this:
$\langle\psi,(I+A)^{\ast}B(I+A)\rangle$ $=\frac{1}{k}\left\langle B\left[\begin{array}[]{c}w_{1}\\ \vdots\\ w_{k}\end{array}\right],\left[\begin{array}[]{c}w_{1}\\ \vdots\\ w_{k}\end{array}\right]\right\rangle$ (7.24)
$=\frac{1}{k}\sum_{i,j}\langle(m_{b_{ij}}\otimes I_{k})w_{j},w_{i}\rangle$
[…]
For fixed $A$, we now recognize the infimum over $B$ in (7.4) as a special case of the variational principle from Proposition 3.9: see Example 3.10. For fixed $A$, and hence fixed $w_{1}$, …, $w_{k}$, we may define $q(\omega)$ to be $\langle w_{j}(\omega),w_{i}(\omega)\rangle$, and then (3.10) shows that the infimum of (7.24) over all permitted choices of $B$ is equal to
$\exp\frac{1}{k}\int\log\det[\langle w_{j}(\omega),w_{i}(\omega)\rangle]\ d\mu(\omega).$ (7.26)
Second infimum. To finish, consider the integrand in (7.26). For each $\omega$ and $i$, the vector $w_{i}(\omega)$ is a perturbation of $x_{i}$ within the subspace $N_{\omega}$, by Lemma 7.11. Among all such perturbations, the Gram determinant is minimized by the tuple $R_{\omega}^{\perp}x_{1}$, …, $R_{\omega}^{\perp}x_{k}$ (see the discussion in Subsection 2.2). Moreover, using elements of $A$ we can bring each $w_{i}(\omega)$ arbitrarily close to $R_{\omega}^{\perp}x_{i}$, again because of Lemma 7.11. Therefore the infimum of (7.26) over all $A\in\mathrm{M}_{k}(\mathfrak{J}\otimes I_{k})$ is equal to
$\exp\frac{1}{k}\int\log\det[\langle R_{\omega}^{\perp}x_{j},R_{\omega}^{\perp}x_{i}\rangle]\ d\mu(\omega).$
7.7 Further remarks
Fuglede–Kadison determinants can be hard to compute in general (see [70] for some recent examples).
Problem 7.12.
Can Theorem B be turned into a concrete evaluation of the Fuglede–Kadison determinants of any specific functionals or operators?
Fuglede–Kadison determinants of matrices over group von Neumann algebras appear in several connections with other parts of mathematics. A succinct account of some of these is given in [42]. A much more thorough exposition emphasizing $L^{2}$-invariants in topology is given in [107, Section 3.2]. See also [107, Chapter 13][97], and the references cited there for an overview of outstanding open problems in this direction, such as Lück’s approximation and determinant conjectures.
In addition, discrete group von Neumann algebras have been generalized to algebras constructed from data such as probability-preserving equivalence relations [56, 57] or unimodular random networks [3]. See the Introduction and Section 5 of the latter reference for an overview. Lück’s determinant conjecture is generalized accordingly to the ‘measure-theoretic determinant conjecture’ in [108], and shown to have consequences for uniform measure equivalence of discrete groups. I expect that Theorem B generalizes in this way as well, for instance by using a version of Example 7.5 over a unimodular random network.
Problem 7.13.
Does the construction of a subdiagonal subalgebra from an equivariant random total order shed any light on specific examples of processes on unimodular random networks, or on any of the open questions discussed in [97]?
Hayes’ paper [83] studies certain measure-preserving systems of algebraic origin for a finitely generated group $\Gamma$ with a left invariant total order. They are constructed from elements of $\mathbf{Z}[\Gamma]$ that he calls ‘lopsided’, which are defined in terms of the order. His main result is that a nondegenerate lopsided group-ring element always gives a factor of a Bernoulli shift.
Problem 7.14.
Can coupling to a random total order generalize Hayes’ construction and result to other finitely generated groups?
8 Approximate association and almost periodic sequences
This section recalls more background in preparation for our introduction of ‘almost periodic entropy’, the new notion that appears in Theorem C.
8.1 Typical vectors and approximate association
Consider again a general separable, unital C*-algebra $\mathfrak{A}$. We can classify tuples of vectors in a representation $\pi$ of $\mathfrak{A}$ according to their type.
For each $k$, let $\Sigma_{k}(\mathfrak{A})$ be the set of normalized elements of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$. In particular, $\Sigma_{1}(\mathfrak{A})$ is just the state space $\Sigma(\mathfrak{A})$. By the identity (4.9), the pairing isomorphism from (4.8) identifies $\Sigma_{k}(\mathfrak{A})$ with $\Sigma(\mathrm{M}_{k}(\mathfrak{A}))$, and this is a norm-bounded subset of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$. It is also compact and convex by the Banach–Alaoglu theorem.
Definition 8.1.
For any positive integer $k$ and subset $O$ of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$, let
$\mathcal{X}(\pi,O):=\big{\{}[v_{1},\dots,v_{k}]^{\mathrm{T}}\in H_{\pi}^{\oplus k}:\ \Phi^{\pi}_{v_{1},\dots,v_{k}}\in O\big{\}}.$
The elements of $\mathcal{X}(\pi,O)$ are the $O$-typical tuples of the representation $\pi$. In addition, let
$\Sigma_{k}(\pi):=\big{\{}\Phi^{\pi}_{v_{1},\dots,v_{k}}:\ v_{1},\dots,v_{k}\in H_{\pi}\ \hbox{and}\ \|v_{1}\|^{2}+\cdots+\|v_{k}\|^{2}=k\big{\}}.$
This is the subset of elements of $\Sigma_{k}(\mathfrak{A})$ that are associated to $\pi$.
The transpose in the definition of $\mathcal{X}(\pi,O)$ is not conceptually significant, but it simplifies some manipulations later.
Observe that
$O\cap\Sigma_{k}(\pi)\neq\emptyset\quad\Leftrightarrow\quad\mathcal{X}(\pi,O)\neq\emptyset$ (8.1)
for any representation $\pi$ and any $O\subset\Sigma_{k}(\mathfrak{A})$.
We often use Definition 8.1 when $O$ is a small neighbourhood around a given ‘target’ completely positive map $\varphi$. In this case we may informally describe elements of $\mathcal{X}(\pi,O)$ as ‘approximately $\varphi$-typical tuples’. This resembles the use of terms such as ‘microstate’ in free probability or ‘good model’ in the study of sofic entropy in ergodic theory: compare [26, Section 2.3], for example.
Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, let $O$ be a neighbourhood of $\varphi$, and let $\widetilde{O}$ be the corresponding neighbourhood of $\langle\varphi,\cdot\rangle$ under the pairing isomorphism (4.8). Then Lemma 4.7 tells us that
$\mathcal{X}(\pi^{(k)},\widetilde{O})=k^{-1/2}\mathcal{X}(\pi,O).$ (8.2)
If $\pi$ is a representation and $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, then $\varphi$ is approximately associated to $\pi$ if it lies in the closure of the set of completely positive maps associated to $\pi$. Because of the second countability from Lemma 4.8, this holds if and only if there is a sequence of $k$-tuples $v_{n,1}$, …, $v_{n,k}$ in $H_{\pi}$ such that
$\Phi^{\pi}_{v_{n,1},\dots,v_{n,k}}(a)\to\varphi(a)\qquad\hbox{for every}\ a\in\mathfrak{A}\ \hbox{as}\ n\to\infty.$ (8.3)
If $\varphi$ is approximately associated to $\pi$ and also normalized, and if the tuples $v_{n,1}$, …, $v_{n,k}$ witness the convergence in (8.3), then we can normalize those tuples slightly to show that $\pi$ actually lies in the weak closure $\overline{\Sigma_{k}(\pi)}$.
The term ‘approximate association’ is not standard. The literature makes more use of a weaker notion: $\varphi$ is weakly associated to $\pi$ if $\varphi$ can be approximated by finite sums of maps associated to $\pi$. This is the basis for Godemont and Fell’s relation of weak containment for representations: see [73][59] or [48, Section 3.4].
Approximate association is finer than weak association. In particular, approximate association can sometimes detect multiplicities in the GNS representation of a positive functional, whereas weak association cannot.
Example 8.2.
[…]
Nevertheless, for some representations the sets of approximately and weakly associated functionals coincide. These include all irreducible representations [48, Theorem 3.4.10], and also regular representations of locally compact groups by a result of Takenouchi (see [140] or [48, Proposition 18.3.5]). For a discrete group, the latter result can actually be deduced from Proposition 3.5. In that case, if $\lambda$ is the regular representation, then it has a tracial cyclic vector, and so Proposition 3.5 shows that $\Sigma_{1}(\lambda^{\oplus\infty})$ is actually equal to $\Sigma_{1}(\lambda)$. Therefore all the intermediate spaces $\Sigma_{1}(\lambda^{\oplus k})$ are equal to $\Sigma_{1}(\lambda)$ as well. On the other hand, given an element of $\Sigma_{1}(\lambda^{\oplus m})$ and an element of $\Sigma_{1}(\lambda^{\oplus n})$, any convex combination of them lies in $\Sigma_{1}(\lambda^{\oplus(m+n)})$.
Our notion of almost periodic entropy (Definition 9.4 below) involves approximations to a positive functional, but it does not allow for taking finite sums. For this reason, approximate association plays a larger role than weak association in the present paper.
Remark 8.3.
[…]
In the next lemma and corollary, we compare typical tuples for two completely positive maps if they are related by association or approximate association. First, fix an $\ell$-by-$k$ matrix $[a_{ij}]$ of elements of $\mathfrak{A}$. If $\pi$ is a representation and ${v_{1},\dots,v_{k}}\in H_{\pi}$, recall that we can define a new $\ell$-tuple ${y_{1},\dots,y_{\ell}}$ in $H_{\pi}$ using $[a_{ij}]$ and ${v_{1},\dots,v_{k}}$ as in formula (4.2). The resulting type of ${y_{1},\dots,y_{\ell}}$ is then related to the type of ${v_{1},\dots,v_{k}}$ by Lemma 4.3. Because of the weak continuity given by that lemma, we obtain the following.
Lemma 8.4.
Let $\psi$ be obtained from $\varphi$ and $[a_{ij}]$ as in Lemma 4.3. For any neighbourhood $O$ of $\psi$, there is a neighbourhood $U$ of $\varphi$ such that
$\big{\{}[\pi(a_{ij})]\cdot[v_{1},\dots,v_{k}]^{\mathrm{T}}:\ [v_{1},\dots,v_{k}]^{\mathrm{T}}\in\mathcal{X}(\pi,U)\big{\}}\subset\mathcal{X}(\pi,O)$
for any representation $\pi$.
In particular, suppose that
$\psi(b)=(Q^{\mathrm{T}})^{\ast}\varphi(b)Q^{\mathrm{T}}\qquad(b\in\mathfrak{A})$
for some $Q\in\mathrm{M}_{\ell,k}$, as in (4.5). Then, for any neighbourhood $O$ of $\psi$, there is a neighbourhood $U$ of $\varphi$ such that
$(I_{H_{\pi}}\otimes Q)[\mathcal{X}(\pi,U)]\subset\mathcal{X}(\pi,O)$
for any representation $\pi$ (identifying $H_{\pi}^{\oplus k}$ with $H_{\pi}\otimes\mathbf{C}^{\oplus k}$ as in (4.5)). ∎
Corollary 8.5.
Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ and $\psi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{\ell})_{+}$, and assume that $\psi$ is approximately associated to $\pi_{\varphi}$. Then for any neighbourhood $O$ of $\psi$ there is a neighbourhood $U$ of $\varphi$ such that
$\mathcal{X}(\pi,U)\neq\emptyset\quad\Rightarrow\quad\mathcal{X}(\pi,O)\neq\emptyset.$
for any representation $\pi$.
Proof.
Let $\varphi$ be associated to $\pi_{\varphi}$ by the cyclic tuple $x_{1}$, …, $x_{k}$. By cyclicity and Lemma 4.2, there is some $\ell$-by-$k$ matrix $[a_{ij}]$ of elements of $\mathfrak{A}$ such that the tuple defined by
$[y_{1},\dots,y_{\ell}]^{\mathrm{T}}:=[\pi(a_{ij})]\cdot[x_{1},\dots,x_{k}]^{\mathrm{T}}$
satisfies $\psi^{\prime}:=\Phi^{\pi}_{y_{1},\dots,y_{\ell}}\in O$. Now apply Lemma 8.4 to $\varphi$, $\psi^{\prime}$ and $[a_{ij}]$. ∎
8.2 Pairs and sums of typical tuples
Let $k$ and $\ell$ be positive integers, and let
$K\>=\{1,\dots,k\}\qquad\hbox{and}\qquad L:=\{k+1,\dots,k+\ell\}.$
Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ and $\psi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{\ell})_{+}$. The next lemma is a robust form of Lemma 4.5.
Lemma 8.6.
If $\varphi$ and $\psi$ are disjoint, then for every neighbourhood $O$ of $\mathrm{diag}(\varphi,\psi)$ there are neighbourhoods $U$ of $\varphi$ and $V$ of $\psi$ such that the following holds:
If $\theta\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k+\ell})_{+}$ satisfies $\theta[K]\in U$ and $\theta[L]\in V$, then $\theta\in O$.
These neighbourhoods satisfy
$\mathcal{X}(\pi,O)\supset\mathcal{X}(\pi,U)\times\mathcal{X}(\pi,V)$
for any representation $\pi$.
Proof.
We prove the first conclusion by contraposition. Assume that $\mathrm{diag}(\varphi,\psi)$ has a neighbourhood $O$ for which no pair of neighbourhoods $U$ and $V$ gives the desired implication. By the second countability from Lemma 4.8, we may therefore choose a sequence $(\theta_{n})_{n\geq 1}$ in $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}\setminus O$ such that
$\theta_{n}[K]\to\varphi\qquad\hbox{and}\qquad\theta_{n}[L]\to\psi.$
This sequence $(\theta_{n})_{n\geq 1}$ must be uniformly bounded in the dual norm because of the identity (4.9) and the relation
$\overline{\mathrm{tr}}_{k+\ell}\theta_{n}(1)=\frac{k}{k+\ell}\overline{\mathrm{tr}}_{k}\big{(}\theta_{n}(1)[K]\big{)}+\frac{\ell}{k+\ell}\overline{\mathrm{tr}}_{\ell}\big{(}\theta_{n}(1)[L]\big{)}.$
Therefore, by the Banach–Alaoglu theorem, $(\theta_{n})_{n\geq 1}$ has a subsequential limit in the weak topology. This limit must be a joining of $\varphi$ and $\psi$, but also it cannot lie in $O$ and so it must be different from $\mathrm{diag}(\varphi,\psi)$. Therefore $\varphi$ and $\psi$ are not disjoint.
The second conclusion follows from the first one and the definition of type. ∎
The assumption of disjointness in Lemma 8.6 is necessary. Indeed, if $\pi_{\varphi}$ is finite-dimensional and irreducible, then $\mathcal{X}(\pi,U)$ is nonempty for any neighourhood $U$ of $\varphi$, but $\mathrm{diag}(\varphi,\varphi)$ is not necessarily approximately associated to $\pi$, only to $\pi^{\oplus 2}$, by reasoning similar to Example 8.2.
Now we consider typical vectors for the sum $\gamma:=\varphi+\psi$. Here we restrict our attention to the case $k=\ell=1$. The case $k=\ell>1$ would not involve any new ideas, but would require heavier notation. Later in the paper we use the constructions of Subsection 4.3 to avoid needing that case.
Lemma 8.7.
Assume that $\varphi$ and $\psi$ are disjoint. For any neighbourhoods $U$ of $\varphi$ and $V$ of $\psi$ there are a neighbourhood $W$ of $\gamma$ and an element $a$ of $\mathfrak{A}$ such that
$\{(\pi(a)x,\pi(1-a)x):\ x\in\mathcal{X}(\pi,W)\}\subset\mathcal{X}(\pi,U)\times\mathcal{X}(\pi,V)$
for any representation $\pi$.
Proof.
The representations $\pi_{\varphi}$ and $\pi_{\psi}$ are disjoint by assumption, and they are both contained in $\pi_{\gamma}$ by Proposition 3.2. Therefore, by the uniqueness of GNS representations, we may identify $\pi_{\gamma}$ with ${\pi_{\varphi}\oplus\pi_{\psi}}$. The resulting orthogonal projection $P$ from $H_{\gamma}$ to $H_{\varphi}$ lies in the centre of $\pi_{\gamma}(\mathfrak{A})^{\prime\prime}$ by Lemma 3.1.
Now suppose that $\gamma$ is associated to $\pi_{\gamma}$ by the cyclic vector $v$. The Kaplansky density theorem [49, Section I.3.5] applied to $\pi_{\gamma}(\mathfrak{A})$ gives an element $a\in\mathfrak{A}$ such that $0\leq a\leq 1$ and such that $\pi_{\gamma}(a)v$ lies as close as we wish to $Pv$. In particular, we may choose $a$ so that
$\Phi^{\pi_{\gamma}}_{\pi_{\gamma}(a)v}=\gamma(a^{\ast}(\cdot)a)\in U\qquad\hbox{and}\qquad\Phi^{\pi_{\gamma}}_{\pi_{\gamma}(1-a)v}=\gamma((1-a)^{\ast}(\cdot)(1-a))\in V.$
Now two applications of Lemma 8.4 produce the required neighbourhood $W$. ∎
Corollary 8.8.
Assume that $\varphi$ and $\psi$ are disjoint.
  1. a.
    For every neighbourhood $W$ of $\gamma$ there are neighbourhoods $U$ of $\varphi$ and $V$ of $\psi$ such that
    $\mathcal{X}(\pi,W)\supset\mathcal{X}(\pi,U)+\mathcal{X}(\pi,V)$
    for any representation $\pi$.
  2. b.
    For any neighbourhoods $U$ of $\varphi$ and $V$ of $\psi$ there is a neighbourhood $W$ of $\gamma$ such that
    $\mathcal{X}(\pi,W)\subset\mathcal{X}(\pi,U)+\mathcal{X}(\pi,V)$
    for any representation $\pi$.
Proof.
First let $W$ be a neighbourhood of $\gamma$. Given a representation $\pi$ and $x$, $y\in H_{\pi}$, observe that
$\Phi^{\pi}_{x+y}(a)=\Phi^{\pi}_{x}(a)+\Phi^{\pi}_{y}(a)+\langle\pi(a)x,y\rangle+\langle\pi(a)y,x\rangle\qquad(a\in\mathfrak{A}).$
We may therefore choose a neighbourhood $O$ of $\mathrm{diag}(\varphi,\psi)$ such that
$\{x+y:\ [x,y]^{\mathrm{T}}\in\mathcal{X}(\pi,O)\}\subset\mathcal{X}(\pi,W)$
for any representation $\pi$. Part (a) follows by concatenating this with the last inclusion from Lemma 8.6.
On the other hand, given the neighbourhoods $U$ of $\varphi$ and $V$ of $\psi$, choose $W$ and $a$ as in Lemma 8.7. Since
$x=\pi(a)x+\pi(1-a)x$
for any $\pi$ and $x\in H_{\pi}$, the conclusion of Lemma 8.7 represents any element of $\mathcal{X}(\pi,W)$ as an element of $\mathcal{X}(\pi,U)+\mathcal{X}(\pi,V)$. ∎
8.3 The strong-quotient topology on representations
Approximate association leads to a natural equivalence relation on representations, and to a topology on equivalence classes.
First, if $X$ is any topological space, then its hyperspace $\mathcal{K}(X)$ is the set of its nonempty compact subsets.
Given now a separable representation $\pi$, let us call $\overline{\Sigma_{k}(\pi)}$ the $k$-summary of $\pi$. It is an element of $\mathcal{K}(\Sigma_{k}(\mathfrak{A}))$. We refer to the whole sequence
$\overline{\Sigma_{\bullet}(\pi)}:=\big{(}\overline{\Sigma_{1}(\pi)},\overline{\Sigma_{2}(\pi)},\dots\big{)}\in\prod_{k=1}^{\infty}\mathcal{K}(\Sigma_{k}(\mathfrak{A}))$ (8.4)
as the summary of $\pi$, and we call $\pi\mapsto\overline{\Sigma_{\bullet}(\pi)}$ the summary map.
Two separable representations $\pi$ and $\rho$ are approximately equivalent if they satisfy $\overline{\Sigma_{\bullet}(\pi)}=\overline{\Sigma_{\bullet}(\rho)}$. This is a classic weakening of equivalence for representations. It is the reason I propose the term ‘approximate association’ in Subsection 8.1. From penetrating works by Glimm [72] and Voiculescu [153], approximate equivalence is known to have several different characterizations, many of them appearing much stronger than the definition above; see also [8] or [38, Section 41]. In those works, the definition we take above is often called ‘weak approximate equivalence’, and it is then proved to imply the other characterizations. However, we do not need this theory below. Approximate equivalence is also the same as ‘weak equivalence in the sense of Zimmer’: see Remark 8.3. Our use of the term ‘summary’ is new.
For any topological space $X$, we endow $\mathcal{K}(X)$ with the classic Vietoris topology: see [54, Exercises 2.7.20, 3.12.27 and 4.5.23] or [92, Subsection I.4.F]. We assume a few standard properties from those references in the sequel. In particular, if $X$ is compact and metrizable, then so is $\mathcal{K}(X)$, for instance using a Hausdorff metric. As a result, in this case we can describe the Vietoris topology in terms of convergence of sequences. More specifically, for any sequence $(K_{n})_{n\geq 1}$ in $\mathcal{K}(X)$, define its topological upper and lower limits to be the sets
$\mathrm{T}\limsup_{n}K_{n}:=\{x\in X:\ \hbox{every neighbourhood of $x$}\\ \hbox{meets $K_{n}$ for infinitely many $n$}\}$
and
$\mathrm{T}\liminf_{n}K_{n}:=\{x\in X:\ \hbox{every neighbourhood of $x$}\\ \hbox{ meets $K_{n}$ for all sufficiently large $n$}\},$
respectively. If $X$ is compact and metrizable, then the sequence $(K_{n})_{n\geq 1}$ Vietoris converges if and only if its topological upper and lower limits are equal, and in this case that common set is the limit of the sequence [92, Subsection I.4.F].
We now endow the product space on the right-hand side of (8.4) with the product of the Vietoris topologies. Each $\Sigma_{k}(\mathfrak{A})$ is compact by the Banach–Alaoglu theorem and metrizable because $\mathfrak{A}$ is separable. Therefore each of those Vietoris topologies is again compact and metrizable, and hence so is that whole countable product topology.
Finally, if $R$ is any set of separable representations (or of equivalence classes of such), then the strong-quotient topology on $R$ is the pullback of that product of Vietoris topologies under the restriction of the summary map to $R$.
As far as I know, this topology was first studied by Abért and Elek in [1]. They were prompted by an analogous idea in graph theory from [20] (see also Remark 8.13 below). Abért and Elek focused on the analogous definition in ergodic theory, but their paper indicates the story for unitary representations of groups as well. However, the term ‘strong-quotient’ is new. We comment on this further after presenting some basic properties.
One key conclusion of the representation theoretic analog of [1, Theorem 1] is this: if one allows all separable representations, then the image of the summary map is a compact subset of the right-hand side of (8.4), and hence closed. The proof adapts with only cosmetic changes to a general separable unital C*-algebra $\mathfrak{A}$, and the result has the following consequence.
Proposition 8.9.
Let $(\pi_{n})_{n\geq 1}$ be a sequence of separable representations of $\mathfrak{A}$, and assume that
$\mathrm{T}\limsup_{n}\overline{\Sigma_{k}(\pi_{n})}=\mathrm{T}\liminf_{n}\overline{\Sigma_{k}(\pi_{n})}\qquad\hbox{for every}\ k=1,2,\dots.$ (8.5)
Then there is a separable representation $\pi$ such that $\pi_{n}\to\pi$ in the strong-quotient topology, and this $\pi$ is unique up to approximate equivalence. ∎
We meet this kind of convergence several times below, so let us write more succinctly that the sequence $(\pi_{n})_{n\geq 1}$ strong-quotient converges if (8.5) holds. For later convenience we also make the following definition, which is a variant of approximate association.
Definition 8.10.
If $(\pi_{n})_{n\geq 1}$ is a sequence of separable representations and $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, then $\varphi$ is asymptotically associated to $(\pi_{n})_{n\geq 1}$ if every neighbourhood $O$ of $\varphi$ satisfies ${\mathcal{X}(\pi_{n},O)\neq\emptyset}$ for infinitely many $n$.
Lemma 8.11.
Let $(\pi_{n})_{n\geq 1}$ be a sequence of separable representations, and let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$. Then $\varphi$ is asymptotically associated to $(\pi_{n})_{n\geq 1}$ if and only if $\langle\varphi,\cdot\rangle$ is asymptotically associated to $(\pi_{n}^{(k)})_{n\geq 1}$.
Proof.
For each neighbourhood $O$ of $\varphi$, let $\widetilde{O}$ be the corresponding neighbhourhood of $\langle\varphi,\cdot\rangle$ under the pairing isomorphism. As $O$ runs over all neighbourhoods of $\varphi$, likewise $\widetilde{O}$ runs over all neighbourhoods of $\langle\varphi,\cdot\rangle$. The result follows from this and the relation (8.2). ∎
By the equivalence (8.1), if $\varphi\in\Sigma_{k}(\mathfrak{A})$, then it is asymptotically associated to $(\pi_{n})_{n\geq 1}$ if and only if it lies in $\mathrm{T}\limsup_{n}\overline{\Sigma_{k}(\pi_{n})}$. The choice of Tlim $\mathrm{T}\limsup$ rather Tlim $\mathrm{T}\liminf$ here is a matter of convention. If $(\pi_{n})_{n\geq 1}$ strong-quotient converges then this choice makes no difference, but otherwise it can be consequential. The next lemma illustrates this.
Lemma 8.12.
Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ and $\psi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{\ell})_{+}$. Assume that $\boldsymbol{\pi}$ strong-quotient converges. If $\varphi$ and $\psi$ are disjoint and they are both asymptotically associated to $\boldsymbol{\pi}$, then so is $\mathrm{diag}(\varphi,\psi)$, and if $k=\ell$ then so is $\varphi+\psi$.
Proof.
Let $O$ be any neighbourhood of $\mathrm{diag}(\varphi,\psi)$. Lemma 8.6 gives neighbourhoods $U$ of $\varphi$ and $V$ of $\psi$ such that
$\mathcal{X}(\pi_{n},O)\supset\mathcal{X}(\pi_{n},U)\times\mathcal{X}(\pi_{n},V)$
for all $n$. By asymptotic association and strong-quotient convergence, both factors in this Cartesian product are nonempty for all sufficiently large $n$, and so $\mathcal{X}(\pi_{n},O)$ is also nonempty for these $n$.
If $k=\ell=1$, then the sum $\varphi+\psi$ is handled in the same way, using Corollary 8.8(a) in place of Lemma 8.6. Finally, the case when $k=\ell>1$ follows by Lemma 8.11. ∎
Without assuming strong-quotient convergence, the two factors in the Cartesian product above might be nonempty only along disjoint subsequences of indices $n$, and then the conclusion of the lemma may fail.
The term ‘strong-quotient’ is chosen to locate this topology relative to some other standard notions for representations. Let us quickly review some of these in order to explain this choice.
The most common topology on the spectrum $\widehat{\mathfrak{A}}$ is the Fell topology, which is the pullback of the Jacobson topology on $\mathrm{Prim}\,\mathfrak{A}$. This has several equivalent definitions, originally explored by Godemont [73] and Fell [59]: see [48, Chapter 3] or [62, Section 7.2] for textbook accounts.
One can try to extend this topology to larger spaces of representations, not just the irreducible ones. But then the different definitions no longer coincide in general. Two possibilities were defined and studied by Fell in [58, Section 3] and [60]. He called them the ‘quotient’ and ‘inner hull-kernel’ topologies. Modern practice usually assigns Fell’s name to his ‘inner hull-kernel topology’, but it is actually the quotient topology that is closer to our present needs.
These topologies can be defined by approximately comparing the sets of completely positive maps associated to representations. The difference boils down to whether one allows approximation by finite sums (giving the Fell topology) or not (giving the quotient topology). This sits alongside our discussion of the difference between weak and approximate association in Subsection 8.1.
Focusing on the quotient topology, it can be defined by pulling back a product topology on the space in (8.4) under the summary map, much like our approach to the strong-quotient topology. But the quotient topology uses the ‘lower Vietoris’ topology on each of the factors $\mathcal{K}(\Sigma_{k}(\mathfrak{A}))$ (see [114, Appendix]), rather than the Vietoris topology. The lower Vietoris topologies are much weaker: indeed, they are usually not even T1, and so limits of sequences need not be unique. Concretely, a sequence $(\pi_{n})_{n\geq 1}$ converges to another representation $\pi$ in the quotient topology if
$\mathrm{T}\liminf_{n}\overline{\Sigma_{k}(\pi_{n})}\supset\overline{\Sigma_{k}(\pi)}\qquad\hbox{for every}\ k=1,2,\dots.$
So, in this topology, any representation whose $k$-summaries are all sufficiently small is a limit of $(\pi_{n})_{n\geq 1}$.
Another manifestation of the weakness of the quotient topology is the following. For a set $R$ of representations (or equivalence classes of them) and an element $a\in\mathfrak{A}$, the norm map
$\pi\mapsto\|\pi(a)\|$ (8.6)
is lower semicontinuous for the quotient topology on $R$, but not necessarily continuous. This is already true for $R=\widehat{\mathfrak{A}}$: see, for instance, the longer discussion in [59, Chapter II] or [48, Section 3.3].
To avoid that weakness, recent research has seen a swell of interest in ‘strong convergence’ of representations. This mode of convergence is defined precisely by requiring that the norm maps in (8.6) converge to their values on a limiting representation. See, for example, the survey [113].
Intuitively, one can understand why strong convergence is not always implied by quotient-topology convergence from the following description. Consider a sequence $(\pi_{n})_{n\geq 1}$ and a desired limit $\pi$.
  • For quotient-topology convergence: for every open subset $O$ of $\Sigma_{k}(\mathfrak{A})$ that meets $\overline{\Sigma_{k}(\pi)}$, and for all sufficiently large $n$, there must be some $k$-tuples $V_{n}$ in $H_{\pi_{n}}$ such that $\Phi^{\pi_{n}}_{V_{n}}\in O$.
  • For strong convergence: for every $a\in\mathfrak{A}$, every $\varepsilon>0$, and all sufficiently large $n$, all unit vectors $v\in H_{\pi_{n}}$ must satisfy $\|\pi_{n}(a)v\|\leq\|\pi(a)\|+\varepsilon$.
On the other hand, control of the norm maps from (8.6) does not require controlling multiplicities, or tuples of more than one vector. In fact, the strong convergence of an AP sequence $(\pi_{n})_{n\geq 1}$ is equivalent to Vietoris convergence of the sequence of closed convex hulls $\overline{\mathrm{conv}}\,\Sigma_{1}(\pi_{n})$.
With all this in mind, I have chosen the term ‘strong-quotient topology’ because strong-quotient convergence is a simultaneous strengthening of (i) convergence in the quotient topology and (ii) strong convergence.
Remark 8.13.
[…]
Further references.
[…]
8.4 Almost periodic sequences
In the next section, the almost periodic entropy of a completely positive map $\varphi$ is defined in terms of the volumes of the $O$-typical sets for certain representations as $O$ ranges over neighbourhoods of $\varphi$. The representations that we use belong to the following sequences.
Definition 8.14.
An almost periodic (‘AP’) sequence for $\mathfrak{A}$ is a sequence of finite-dimensional representations of $\mathfrak{A}$ whose dimensions diverge to $\infty$.
Fix an AP sequence $\boldsymbol{\pi}=(\pi_{n})_{n\geq 1}$, and let $d_{n}$ be the dimension of $\pi_{n}$ for each $n$. The strong-quotient topology provides a natural mode of convergence for such a sequence. Since each $d_{n}$ is finite, by Lemma 4.2 the sets $\Sigma_{k}(\pi_{n})$ are continuous images of finite-dimensional spheres, and hence already closed. Therefore we do not need to take their closures in this case.
Later we also need another mode of convergence for an AP sequence $\boldsymbol{\pi}$. This is the convergence of the tracial states $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}$ to some limit tracial state $\tau$ in the weak topology of $\mathfrak{A}_{+}^{\ast}$. In the terminology of free probability theory, this asserts that, for any finite subset $F$ of $\mathfrak{A}$, the tuples $(\pi_{n}(a):\ a\in F)$ form a sequence of ‘microstates’ for $F$ in the ‘non-commutative probability space’ $(\mathfrak{A},\tau)$ (see [156], for example). This convergence of tracial states is one of the hypotheses of Theorem C. It determines which limiting tracial state should be used to define the Fuglede–Kadison determinant in the conclusion of that theorem.
The next lemma is a companion to Lemma 8.11.
Lemma 8.15.
Let $(\pi_{n})_{n\geq 1}$ be an AP sequence for $\mathfrak{A}$, and consider the AP sequence $(\pi^{(k)}_{n})_{n\geq 1}$ for $\mathrm{M}_{k}(\mathfrak{A})$. If $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}\to\tau$, then $\overline{\mathrm{tr}}_{kd_{n}}\circ\pi_{n}^{(k)}\to\tau\otimes\overline{\mathrm{tr}}_{k}$.
Proof.
After identifying $\overline{\mathrm{tr}}_{kd_{n}}$ with $\overline{\mathrm{tr}}_{d_{n}}\otimes\overline{\mathrm{tr}}_{k}$, this follows from the continuity of the operation $(\cdot)\otimes\overline{\mathrm{tr}}_{k}$ on tracial states. That continuity can be checked entry-wise against elements of $\mathrm{M}_{k}(\mathfrak{A})$. ∎
In general, neither of strong-quotient convergence nor convergence of tracial states implies the other. But convergence of tracial states does give a lower bound on strong-quotient limits: see Corollary 9.15 below.
9 Almost periodic entropy
In this section we define the almost periodic entropy of an $\mathrm{M}_{k}$-valued completely positive map on a separable, unital C*-algebra $\mathfrak{A}$. We then build up its properties towards the proof of Theorem C, and follow that with a few consequences.
9.1 Preliminary results from high-dimensional probability
For the rest of this section, let
$v(d):=\frac{\pi^{d}}{d!}\qquad(d=1,2,\dots).$ (9.1)
Then $v(d)$ is equal to the volume of the unit ball in $\mathbf{C}^{d}$ [127, Subsection 1.4.9]. By Stirling’s approximation, this function satisfies
$k^{kd}\cdot v(kd)=e^{o(d)}\cdot v(d)^{k}$ (9.2)
for any fixed $k$ as $d\to\infty$.
Recall that we write $\sigma_{2d-1}$ for the normalized spherical measure on the unit sphere $\mathrm{S}^{2d-1}$ in $\mathbf{C}^{d}$ or in any other space with a standing identification with $\mathbf{R}^{2d}$. A key feature of these measures in high dimensions is the phenomenon of measure concentration. We need the following special case.
Lemma 9.1.
There is an absolute positive constant $c$ such that, for any positive integer $d$ and linear transformation $A$ of $\mathbf{C}^{d}$, we have
$\sigma_{2d-1}\{v:\ |\langle Av,v\rangle-\overline{\mathrm{tr}}_{d}A|\geq t\}\leq 4e^{-ct^{2}d/\|A\|^{2}}\qquad(t>0).$
Proof.
If $f(v)=\langle Av,v\rangle$ for $v\in\mathrm{S}^{2d-1}$, then $f$ is $2\|A\|$-Lipschitz, and its spherical average is $\overline{\mathrm{tr}}_{d}A$ by the invariance of trace under conjugation. Therefore the desired inequality follows from the concentration of general Lipschitz functions on high-dimensional spheres. See, for instance, [115, Corollary V.2] for a proof of this concentration result from its Gaussian analog, or [102] for a more modern proof by interpolation along the heat semigroup. ∎
For any positive integers $d$ and $k$ and any $O\subset\mathrm{M}_{k+}$, let
$T(d,O):=\{X\in\mathrm{M}_{d,k}:\ X^{\ast}X\in O\}.$ (9.3)
This is the set of $k$-tuples in $\mathbf{C}^{\oplus d}$ whose Gram matrices lie in $O$.
Corollary 9.2.
Let $k$ be a positive integer and let $O$ be any neighbourhood of $I_{k}$ in $\mathrm{M}_{k+}$. Then there are positive constants $C$ and $c$ (depending on $k$ and $O$) such that
$\sigma_{2dk-1}\big{(}k^{-1/2}T(d,O)\big{)}>1-Ce^{-cd}\qquad(d=1,2,\dots).$
Proof.
By shrinking $O$ if necessary, we may assume it is equal to
$\bigcap_{i=1}^{k}\{Q\in\mathrm{M}_{k+}:\ |q_{ii}-1|<\varepsilon\}\cap\bigcap_{1\leq i<j\leq k}\{Q\in\mathrm{M}_{k+}:\ |q_{ij}|<\varepsilon\}$
for some $\varepsilon>0$. In this case, we have
$k^{-1/2}T(n,O)=\bigcap_{i=1}^{k}\{[x_{1},\dots,x_{k}]\in\mathrm{M}_{d,k}:\ |\langle x_{i},x_{i}\rangle-1/k|<\varepsilon/k\}\\ \cap\bigcap_{1\leq i<j\leq k}\{[x_{1},\dots,x_{k}]\in\mathrm{M}_{d,k}:\ |\langle x_{i},x_{j}\rangle|<\varepsilon/k\}.$
This is an intersection of at most $k^{2}$ sets, and for each of them its complement has $\sigma_{2kd-1}$-measure controlled by Lemma 9.1. ∎
Lemma 9.14 below is a generalization of Corollary 9.2 that depends on Lemma 9.1 in the same way.
Subsection 1.2 discusses the analogy between the Gram matrix of a tuple of vectors and the joint distribution of a tuple of finite-valued random variables. The next theorem adds another layer to this analogy: a ‘method of types’ interpretation for $\log\det Q$ when $Q$ is a positive semi-definite matrix. It approximates the volumes of the sets $T(d,O)$ when $O$ is a small neighbourhood of $Q$. In information theory, the method of types captures the basic combinatorial meaning of discrete Shannon entropy [40, Section 11.1], and similarly the large-deviations interpretation of Kullback–Leibler divergence [41, Subsection 2.1]. The differential entropy of a jointly Gaussian random vector also has a method-of-types interpretation in terms of volumes [40, Section 8.2].
Theorem 9.3.
Let $Q$ be a $k$-by-$k$ positive semidefinite matrix.
  • a.
    (Lower bound) If $O$ is any neighbourhood of $Q$ in $\mathrm{M}_{k+}$, then
    $\frac{\mathrm{vol}_{2kd}T(d,O)}{v(d)^{k}}\geq(\det Q)^{d-o(d)}.$
  • b.
    (Upper bound) For any $a>\det Q$ there is a neighbourhood $O$ of $Q$ in $\mathrm{M}_{k+}$ such that
    $\frac{\mathrm{vol}_{2kd}T(d,O)}{v(d)^{k}}\leq a^{d+o(d)}.$
Variants of Theorem 9.3 are widely known, for instance in the study of complex Wishart distributions for random Gram matrices [74]. But I have not found a convenient reference for this particular one, so include a full proof.
Proof.
We write a typical element of $\mathrm{M}_{d,k}$ as $X=[x_{1},\dots,x_{k}]$, and write
$\|X\|_{2}^{2}:=\sum_{i=1}^{k}\|x_{i}\|^{2}.$
Step 1. We first prove part (a) for $Q=I_{k}$. By shrinking the neighbourhood $O$ if necessary, we may assume that
$O=\big{\{}Q^{\prime}\in\ \mathrm{M}_{k+}:\ e^{-2\varepsilon}<\overline{\mathrm{tr}}_{k}Q^{\prime}<e^{2\varepsilon}\ \hbox{and}\ (\overline{\mathrm{tr}}_{k}Q^{\prime})^{-1}\cdot Q^{\prime}\in U\big{\}}$
for some $\varepsilon>0$ and some other neighbourhood $U$ of $I_{k}$ in $\mathrm{M}_{k+}$. This turns into
$T(d,O)=\big{\{}X\in\mathrm{M}_{d,k}:\ \sqrt{k}e^{-\varepsilon}<\|X\|_{2}<\sqrt{k}e^{\varepsilon}\ \hbox{and}\ X/\|X\|_{2}\in k^{-1/2}T(d,U)\big{\}}.$
Identifying $\mathrm{M}_{d,k}$ with $\mathbf{C}^{dk}$ and integrating in polar coordinates [127, Subsection 1.4.3], we obtain
[…]
$=k^{kd}\cdot(e^{2kd\varepsilon}-e^{-2kd\varepsilon})\cdot v(kd)\cdot\sigma_{2kd-1}\big{(}k^{-1/2}T(n,U)\big{)}.$
By Corollary 9.2 and the asymptotic (9.2), this is greater than $v(d)^{k}$ for all sufficiently large $d$.
Step 2. On the other hand, for any $\varepsilon>0$, the set
$O:=\{Q^{\prime}\in\mathrm{M}_{k+}:\ \overline{\mathrm{tr}}_{k}Q^{\prime}<e^{2\varepsilon}\}$
is a neighbourhood of $I_{k}$ in $\mathrm{M}_{k+}$, and it satisfies
$\mathrm{vol}_{2kd}T(d,O)=\mathrm{vol}_{2kd}B_{\sqrt{k}e^{\varepsilon}}(0)=k^{kd}\cdot e^{2kd\varepsilon}\cdot v(kd)=e^{2kd\varepsilon+o(d)}\cdot v(d)^{k},$
using (9.2) again. Since $\varepsilon$ is arbitrary, this proves part (b) for $Q=I_{k}$.
Step 3. Now let $Q\in\mathrm{M}_{k+}$, let $R\in\mathrm{M}_{k}$, and let $Q^{\prime}:=R^{\ast}QR$. Then $Q^{\prime}$ also lies in $\mathrm{M}_{k+}$. By the continuity of matrix multiplication, if $O^{\prime}$ is any neighbourhood of $Q^{\prime}$, then $Q$ has a neighbourhood $O$ such that
$O^{\prime}\supset R^{\ast}\cdot O\cdot R.$
In terms of tuples of vectors, this turns into
$T(d,O^{\prime})\supset\{XR:\ X\in T(d,O)\}.$ (9.5)
If we regard a $d$-by-$k$ matrix as the $d$-tuple of its rows, then right-multiplication by $R$ on $\mathrm{M}_{d,k}$ becomes the direct sum of $d$ copies of $R^{\mathrm{T}}$ acting on $\mathbf{C}^{k}$. Regarded as a real linear transformation acting on $2kd$ real linear dimensions, this has Jacobian $|\det R|^{2d}$: see, for instance, [127, Subsection 1.3.5]. Therefore (9.5) gives
$\mathrm{vol}_{2kd}T(d,O^{\prime})\geq|\det R|^{2d}\cdot\mathrm{vol}_{2kd}T(d,O).$
Since $O^{\prime}$ is an arbitrary neighbourhood of $Q^{\prime}$, and
$\det Q^{\prime}=|\det R|^{2}\cdot\det Q,$
this shows that part (a) for $Q^{\prime}$ follows from part (a) for $Q$. Similarly, if $R$ is invertible, then we may reverse the roles of $Q$ and $Q^{\prime}$ above and deduce that part (b) for $Q^{\prime}$ follows from part (b) for $Q$.
In particular, combining this reasoning with Steps 1 and 2 and making the choice $R=Q^{-1/2}$, we conclude parts (a) and (b) whenever $Q$ is nonsingular.
Step 4. Finally, assume that $Q$ is singular. Then part (a) is vacuous.
Applying Step 3 with $R$ a unitary matrix, we may assume that $Qe_{1}=0$ for the standard basis $e_{1}$, …, $e_{k}$ of $\mathbf{C}^{k}$. Having done so, let $r>\max_{i}\sqrt{q_{ii}}$, let $\varepsilon>0$, and let
$O:=\{Q^{\prime}\in\mathrm{M}_{k+}:\ q_{11}^{\prime}<\varepsilon^{2}\ \hbox{and}\ \max_{i}q^{\prime}_{ii}<r^{2}\}.$
Then $O$ is a neighbourhood of $Q$, and
$T(d,O)=\big{\{}X\in\mathrm{M}_{d,k}:\ \|x_{1}\|<\varepsilon\ \hbox{and}\ \max_{i}\|x_{i}\|<r\big{\}}.$
Therefore this neighbourhood satisfies
$\frac{\mathrm{vol}_{2kd}T(d,O)}{v(d)^{k}}\leq\frac{(\varepsilon^{2d}\cdot v(d))\cdot(r^{2d}\cdot v(d))^{k-1}}{v(d)^{k}}=\varepsilon^{2d}\cdot r^{2(k-1)d}.$
Since $\varepsilon$ can be chosen independently of $r$, this completes the proof of part (b). ∎
Theorem 9.3 is a template for Theorem C (and also a special case of that theorem with $\mathfrak{A}=\mathrm{M}_{k}$). The rather soft proof given above, which avoids evaluating any integrals over spaces of matrices exactly, is also a precursor to the proof of Theorem C.
For discrete Shannon entropy, the usual proofs in the method of types involve counting strings that have exactly a given empirical distribution and then applying Stirling’s approximation. By contrast, in Step 3 of the proof of Theorem 9.3, we use a change of variables to transport the desired estimate from $I_{k}$ to most other elements of $\mathrm{M}_{k+}$. This way of using the symmetries of $\mathrm{vol}_{2kn}$ has no obvious analog for probability distributions over finite sets. A related use of symmetry is also essential to our later proof of Theorem C.
9.2 Definition and first properties of almost periodic entropy
Consider again a separable, unital C*-algebra $\mathfrak{A}$. Fix an AP sequence $\boldsymbol{\pi}=(\pi_{n})_{n\geq 1}$, a positive integer $k$, and an element $\varphi$ of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$. Let $d_{n}$ be the dimension of $\pi_{n}$ for each $n$. Recall Definition 8.1 of the sets $\mathcal{X}(\pi_{n},O)$.
Definition 9.4.
The almost periodic (‘AP’) entropy of $\varphi$ along $\boldsymbol{\pi}$ is the quantity
$\mathrm{h}_{\boldsymbol{\pi}}(\varphi):=\inf_{O}\limsup_{n\to\infty}\frac{1}{d_{n}}\log\frac{\mathrm{vol}_{2kd_{n}}\mathcal{X}(\pi_{n},O)}{v(d_{n})^{k}},$ (9.6)
where the infimum runs over all neighbourhoods of $\varphi$.
Remark 9.5.
[…]
The constant outside the logarithm in (9.6) is a normalization choice, and is somewhat arbitrary. One could divide by $2d_{n}$ instead, but I have preferred to emphasize complex rather than real dimensions. Another natural choice would be $kd_{n}$, the complex dimension of the ambient space containing $\mathcal{X}(\pi_{n},O)$. However, the present choice seems to make for fewer explicit factors of $k$ later, for example in Lemma 9.7 below.
[…]
Lemma 9.6.
For any $\boldsymbol{\pi}$ and $k$, the function $\mathrm{h}_{\boldsymbol{\pi}}$ is upper semicontinuous on $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$.
Proof.
This holds because $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ is defined as an infimum of values associated to the neighbourhoods of $\varphi$. ∎
Lemma 9.7.
Any $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ satisfies
$\mathrm{h}_{\boldsymbol{\pi}}(\varphi)\leq\log\det\varphi(1).$
In particular, if $\varphi(1)$ is singular then $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)=-\infty$.
Proof.
If $O$ is any neighbourhood of $\varphi(1)$ in $\mathrm{M}_{k}$, then the set
$U:=\{\psi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}:\ \psi(1)\in O\}$
is a neighbourhood of $\varphi$ in $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$. If $\pi$ is a $d$-dimensional representation, then this $U$ satisfies
$\mathcal{X}(\pi,U)=\{V^{\mathrm{T}}=[x_{1},\dots,x_{k}]^{\mathrm{T}}\in(\mathbf{C}^{\oplus d})^{\oplus k}:\ V^{\ast}V\in O\}.$
Therefore, for any $h>\log\det\varphi(1)$, Theorem 9.3(b) gives
$\log\frac{\mathrm{vol}_{2kd}\mathcal{X}(\pi,U)}{v(k)^{d}}<hd$
for all sufficiently large $d$. Since $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ is an infimum over all neighbourhoods of $\varphi$ in $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, the particular neighbourhoods considered above show that $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)<h$ whenever $h>\log\det\varphi(1)$. ∎
Lemma 9.7 is analogous to the inequality between the entropy of a partition and the entropy rate of the generated process under a measure-preserving transformation: see, for instance, [157, Theorem 4.12(i)]. We determine the cases of equality in Proposition 9.17 below.
For a given AP sequence $\boldsymbol{\pi}=(\pi_{n})_{n\geq 1}$ and completely positive map $\varphi$, the property of $\varphi$ being asymptotically associated to $\boldsymbol{\pi}$ depends only on the equivalence class of $\pi_{\varphi}$. However, the actual value $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ is more sensitive. This is an important point where our story diverges from ergodic theory. In ergodic theory, one of the most essential properties of sofic entropy is its independence of the choice of generating observable [25], and hence its invariance under isomorphism for abstract measure-preserving systems. The analog of this for AP entropy is false. Even if we restrict attention to single cyclic unit vectors in a representation $\pi$, the resulting values of AP entropy can vary greatly.
However, rather than invariance, AP entropy enjoys a general transformation law when one cyclic tuple is exchanged for another: see Proposition 9.10.
Some proofs about AP entropy are easier to digest in the special case $k=1$, if only because the notation is lighter. The next result sometimes lets us make this simplification without losing any generality. It accompanies Lemmas 8.11 and 8.15.
Lemma 9.8.
Let $k$ be a positive integer, let $\boldsymbol{\pi}=(\pi_{n})_{n\geq 1}$ be an AP sequence, and let $\boldsymbol{\pi}^{(k)}:=(\pi_{n}^{(k)})_{n\geq 1}$. Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, and define $\langle\varphi,\cdot\rangle$ as in equation (4.7). Then
$\mathrm{h}_{\boldsymbol{\pi}^{(k)}}(\langle\varphi,\cdot\rangle)=\frac{1}{k}\mathrm{h}_{\boldsymbol{\pi}}(\varphi).$
Proof.
Let $O$ be a neighbourhood of $\varphi$, and let $\widetilde{O}$ be the associated neighbourhood of $\langle\varphi,\cdot\rangle$ under the pairing isomorphism from Subsection 8.1. Then we have
$\frac{\mathrm{vol}_{2kd_{n}}\mathcal{X}(\pi_{n}^{(k)},\widetilde{O})}{v(kd_{n})}=\frac{k^{-kd_{n}}\cdot\mathrm{vol}_{2kd_{n}}\mathcal{X}(\pi_{n},O)}{v(kd_{n})}=\frac{e^{o(d_{n})}\cdot\mathrm{vol}_{2kd_{n}}\mathcal{X}(\pi_{n},O)}{v(d_{n})^{k}},$
using (8.2) for the first equality and (9.2) for the second. Now take logarithms, normalize by $kd_{n}$ (the dimension of $\pi_{n}^{(k)}$), and insert into Definition 9.4. ∎
Remark 9.9.
The Introduction discusses AP entropy as a representation theoretic analog of sofic entropy. Let us comment further on their relationship.
Theorem C applies to an AP sequence whose pulled-back traces converge to some $\tau$. As a result, the Fuglede–Kadison determinant that appears in Theorem C can arise as an AP entropy value whenever $\tau$ is a limit of tracial states pulled back from some AP sequence.
Suppose that $\mathfrak{A}=C^{\ast}\Gamma$ and that $\tau$ comes from the regular charater. Then it is a limit of finite-dimensional characters if and only if $\Gamma$ has an AP sequence $(\pi_{i})_{i\geq 1}$ that separates its elements. Such a group is called ‘maximally almost periodic’. For finitely presented groups this property turns out to be equivalent to residual finiteness. Maximally almost periodic groups were introduced by von Neumann [150] and covered by Weil in [158, Chapter VII] and Dixmier in [48, Sections 16.4–5] (where they are called ‘injectable’ groups).
However, we can also study AP entropy for regular characters on a larger class of groups as follows. If $\Gamma$ is any countable group, then we can write it as $F/N$ for some free group $F$ and normal subgroup $N$. Now we can look for finite-dimensional representations of $F$ whose characters converge to the quasi-regular character $1_{N}$, rather than finite-dimensional representations of $\Gamma$ whose characters converge to $1_{\{e\}}$. This offers more flexibility, because those finite-dimensional representations of $F$ need not have trivial restriction to $N$ until we take their limit. Allowing convergence in this sense, the availability of finite-dimensional approximants to the regular representation of $\Gamma$ is equivalent to $\Gamma$ being ‘hyperlinear’ in the terminology of [125]; in particular, it does not depend on the choice of presentation $F/N$. This class of groups is the ‘linear’ analog of the sofic groups, and includes the sofic groups by a result from [53]. Analogously, soficity itself can be formulated in terms of quotients of a free group and finitary approximations to the resulting Schreier graphs; see [32, Section 4], for example.
[…]
Further references.
See [14, Subsection 4.C.b] for a modern introduction to maximally almost periodic groups, and [30, Section 6.4.2] for a connection to Kirchberg’s ‘factorization property’.
[…]
9.3 First transformation formula
Fix an AP sequence $\boldsymbol{\pi}=(\pi_{n})_{n\geq 1}$, and let $d_{n}$ be the dimension of $\pi_{n}$. At this point, some of our work starts to need the assumption that the tracial states $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}$ converge to a limit in $\mathfrak{A}^{\ast}$.
The next proposition is a basic change-of-variables formula for AP entropy. It can be seen as a cousin of Voiculescu’s change-of-variables formulas for free entropy in free probability theory [154, 155], although the proof in our ‘linear’ setting is simpler.
Proposition 9.10.
Fix $\varphi\in\mathfrak{B}(\mathfrak{A};\mathrm{M}_{k})_{+}$.
  1. a.
    Let $Q\in\mathrm{M}_{k}$ be invertible, and define
    $\psi(b):=(Q^{\mathrm{T}})^{\ast}\varphi(b)Q^{\mathrm{T}}\qquad(b\in\mathfrak{A}).$
    Then
    $\mathrm{h}_{\boldsymbol{\pi}}(\psi)=2\log|\det Q|+\mathrm{h}_{\boldsymbol{\pi}}(\varphi).$
  2. b.
    Assume further that $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}\to\tau$. Let $a\in\mathrm{M}_{k}(\mathfrak{A})$ be invertible, and define $\psi\in\mathfrak{B}(\mathfrak{A};\mathrm{M}_{k})_{+}$ in terms of $\varphi$ and $a$ as in Lemma 4.3. Then $\varphi$ is asymptotically associated to $\boldsymbol{\pi}$ if and only if $\psi$ is, and in that case
    $\mathrm{h}_{\boldsymbol{\pi}}(\psi)=2\log\Delta_{\tau\otimes\mathrm{tr}_{k}}|a|+\mathrm{h}_{\boldsymbol{\pi}}(\varphi).$
Proof.
Part (a). For any neighbourhood $O$ of $\psi$, the second part of Lemma 8.4 gives a neighbourhood $U$ of $\varphi$ such that
$(I_{d}\otimes Q)[\mathcal{X}(\pi,U)]\subset\mathcal{X}(\pi,O)$
for any $d$-dimensional representation $\pi$. For each $n$, it follows that
$\mathrm{vol}_{2kd_{n}}\mathcal{X}(\pi_{n},O)$ $\geq|\det(I_{d_{n}}\otimes Q)|^{2}\cdot\mathrm{vol}_{2kd_{n}}\mathcal{X}(\pi_{n},U)$
$=|\det Q|^{2d_{n}}\cdot\mathrm{vol}_{2kd_{n}}\mathcal{X}(\pi_{n},U).$
The determinants are squared here because $I_{d_{n}}\otimes Q$ is a linear transformation in $kd_{n}$ complex dimensions, but we must treat it as a real linear map in $2kd_{n}$ real dimensions for the purpose of computing volumes (see, for instance, [127, Subsection 1.3.5]). Inserting this inequality into Definition 9.4 and taking the infimum over $O$, it follows that
$\mathrm{h}_{\boldsymbol{\pi}}(\psi)\geq 2\log|\det Q|+\mathrm{h}_{\boldsymbol{\pi}}(\varphi).$
Applying the same reasoning with the roles of $\varphi$ and $\psi$ reversed and with $Q^{-1}$ in place of $Q$, we obtain the reverse inequality as well.
Part (b). This time, if $O$ is a neighbourhood of $\psi$, then Lemma 8.4 gives a neighbourhood $O^{\prime}$ of $\varphi$ such that
$\mathrm{vol}_{2kd_{n}}\mathcal{X}(\pi_{n},O)$ $\geq\mathrm{vol}_{2kd_{n}}\big{(}\pi_{n}^{(k)}(a)[\mathcal{X}(\pi_{n},O^{\prime})]\big{)}$
$=\mathrm{det}\,|\pi_{n}^{(k)}(a)|^{2}\cdot\mathrm{vol}_{2kd_{n}}\mathcal{X}(\pi_{n},O^{\prime})$ (9.7)
for every $n$. Once again, the determinant is squared because we must treat $\pi_{n}^{(k)}(a)$ as a linear map in $2kd_{n}$ real dimensions for the purpose of computing volumes.
Since $a$ is invertible and $\pi_{n}^{(k)}$ is a unital C*-algebra homomorphism, we have
$\mathrm{det}\,|\pi_{n}^{(k)}(a)|=\exp(\mathrm{tr}_{kd_{n}}(\log|\pi_{n}^{(k)}(a)|))=\exp\big{(}\mathrm{tr}_{kd_{n}}(\pi_{n}^{(k)}(\log|a|))\big{)}.$
By our assumption on $\boldsymbol{\pi}$ and Lemma 8.15, this is equal to
$\exp\big{(}d_{n}\big{(}(\tau\otimes\mathrm{tr}_{k})(\log|a|)+o(1)\big{)}\big{)}=(\Delta_{\tau\otimes\mathrm{tr}_{k}}|a|)^{d_{n}+o(d_{n})}\qquad\hbox{as}\ n\to\infty.$
Therefore, normalizing and taking logarithms in (9.3), that inequality becomes
$\frac{1}{d_{n}}\log\frac{\mathrm{vol}_{2kd_{n}}\mathcal{X}(\pi_{n},O)}{v(d_{n})^{k}}\geq 2\log\Delta_{\tau\otimes\mathrm{tr}_{k}}|a|+o(1)+\frac{1}{d_{n}}\log\frac{\mathrm{vol}_{2kd_{n}}\mathcal{X}(\pi_{n},O^{\prime})}{v(d_{n})^{k}}.$
Letting $n\to\infty$ and then taking the infimum over $O$, this gives
$\mathrm{h}_{\boldsymbol{\pi}}(\psi)\geq 2\log\Delta_{\tau\otimes\mathrm{tr}_{k}}|a|+\mathrm{h}_{\boldsymbol{\pi}}(\varphi).$
The reverse of this inequality also holds by swapping the roles of $\varphi$ and $\psi$ and replacing $a$ with $a^{-1}$, which satisfies
$\Delta_{\tau\otimes\mathrm{tr}_{k}}|a^{-1}|=(\Delta_{\tau\otimes\mathrm{tr}_{k}}|a|)^{-1}$
(see, for instance, [49, Theorem I.6.10(iii)]). ∎
Remark 9.11.
[…]
9.4 Spherical measures and concentration
When $k=1$ and $\varphi$ is normalized, the next lemma gives an alternative to using Lebesgue measure in Definition 9.4.
Lemma 9.12.
Let $\varphi\in\Sigma(\mathfrak{A})$ and let $\boldsymbol{\pi}$ be an AP sequence. Let $\mathcal{O}$ be a base of neighbourhoods around $\varphi$ in $\Sigma(\mathfrak{A})$. Then
$\mathrm{h}_{\boldsymbol{\pi}}(\varphi)=\inf_{O\in\mathcal{O}}\limsup_{n\to\infty}\frac{1}{d_{n}}\log\sigma_{2d_{n}-1}\mathcal{X}(\pi_{n},O).$ (9.8)
Proof.
Let $\mathcal{U}$ be the family of sets that have the form
$\big{\{}\psi\in\mathfrak{A}_{+}^{\ast}:\ e^{-2\delta}<\psi(1)<e^{2\delta}\ \hbox{and}\ \psi(1)^{-1}\cdot\psi\in O\big{\}}$ (9.9)
for some $O\in\mathcal{O}$ and $\delta>0$. Then $\mathcal{U}$ is a base of neighbourhoods around $\varphi$ in $\mathfrak{A}_{+}^{\ast}$. We may therefore restrict attention to neighbourhoods from $\mathcal{U}$ when evaluating $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ (see Remark 9.5).
So now let $U$ be the set in (9.9) for some $\delta>0$ and $O\in\mathcal{O}$, and let $\pi$ be a $d$-dimensional representation. Then the special form of $U$ gives
$1_{\mathcal{X}(\pi,U)}(ry)=1_{(e^{-\delta},e^{\delta})}(r)\cdot 1_{\mathcal{X}(\pi,O)}(y)\qquad(r>0,\ y\in\mathrm{S}^{2d-1}).$
As a result, integrating in polar coordinates [127, Subsection 1.4.3] gives
[…]
$=v(d)\cdot(e^{2\delta d}-e^{-2\delta d})\cdot\sigma_{2d-1}\mathcal{X}(\pi,O).$
This implies that
$\sigma_{2d-1}\mathcal{X}(\pi,O)\leq\frac{\mathrm{vol}_{2d}\mathcal{X}(\pi,U)}{v(d)}\leq e^{2\delta d}\cdot\sigma_{2d-1}\mathcal{X}(\pi,O)$
for all sufficiently large $d$. Inserting this into Definition 9.4 and taking the limit supremum over $n$ and then the infimum over elements if $\mathcal{U}$, we obtain (9.8). ∎
Remark 9.13.
[…]
If $\varphi\in\Sigma_{k}(\mathfrak{A})$ for some $k>1$, then we may apply Lemma 9.12 to the positive functional $\langle\varphi,\cdot\rangle$ on $\mathrm{M}_{k}(\mathfrak{A})$, whose sets of approximately typical vectors along $\boldsymbol{\pi}^{(k)}$ are given by equation (8.2). We can then deduce a result for $\varphi$ itself via Lemma 9.8.
For a state, we now have a choice between the original definition of AP entropy and the alternative given by Lemma 9.12. Each has its advantages. A major advantage of the measures $\sigma_{2d-1}$ is contact with the measure concentration from Lemma 9.1. As above, we discuss this only for $k=1$ to lighten notation. Let $\pi$ be a $d$-dimensional representation. Since the trace of a matrix is invariant under unitary conjugation, the average of the type $\Phi^{\pi}_{x}$ with respect to the spherical measure $\sigma_{2d-1}$ is equal to $\overline{\mathrm{tr}}_{d}\circ\pi$. When $d$ is large, Lemma 9.1 improves this conclusion considerably: $\Phi^{\pi}_{x}$ is actually close to $\overline{\mathrm{tr}}_{d}\circ\pi$ for most individual $x\in\mathrm{S}^{2d-1}$.
Let $\tau$ be a tracial state on $\mathfrak{A}$, and let $\lambda$ be its GNS representation.
Lemma 9.14.
If $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}\to\tau$, then for every neighbourhood $U$ of $\tau$ there are positive constants $C$ and $c$ such that
$\sigma_{2d_{n}-1}\mathcal{X}(\pi,U)\geq 1-Ce^{-cd_{n}}.$
Proof.
It suffices to prove this for $U$ belonging to some sub-base of neighbourhoods of $\tau$, since any other neighbourhood contains a finite intersection of these. We may therefore assume that
$U=\{\psi\in\Sigma(\mathfrak{A}):\ |\psi(a)-\tau(a)|<\varepsilon\}$
for some $a\in\mathfrak{A}$ and $\varepsilon>0$.
Since $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}(a)\to\tau(a)$, this $U$ satisfies
$\mathcal{X}(\pi_{n},U)\supset\{v\in\mathrm{S}^{2d_{n}-1}:\ |\langle\pi_{n}(a)v,v\rangle-\overline{\mathrm{tr}}_{d_{n}}\pi_{n}(a)|<\varepsilon/2\}$
for all sufficiently large $n$. Now the result follows Lemma 9.1. ∎
Corollary 9.15.
Assume that $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}\to\tau$. If $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ and $\varphi$ is approximately associated to $\lambda^{\oplus\infty}$, then it is asymptotically associated to $(\pi_{n})_{n\geq 1}$.
Proof.
We may assume that $\varphi$ is normalized. Proposition 4.10 gives that $\Sigma_{k}(\lambda^{\oplus\infty})$ equals $\Sigma_{k}(\lambda^{\oplus k})$, so now our assumption actually says that $\varphi\in\overline{\Sigma_{k}(\lambda^{\oplus k})}$.
Since $\lambda^{\oplus k}$ is the minimal dilation of $\tau\otimes I_{k}$, by Corollary 8.5 it suffices to show that $\tau\otimes I_{k}$ itself is asymptotically associated to $\boldsymbol{\pi}$ (recall Example 4.1 for this tensor-product notation). When $k=1$, this follows from Lemma 9.14: indeed, a random vector drawn from $\sigma_{2d_{n}-1}$ is approximately typical for $\tau$ with high probability once $d_{n}$ is large enough. Finally, if $k>1$, then we can apply the previous case to the sequence $\boldsymbol{\pi}^{(k)}$ using Lemmas 8.15 and 8.11 and the fact that
$\langle\tau\otimes I_{k},\cdot\rangle=\tau\otimes\overline{\mathrm{tr}}_{k}.$
In the notation of Subsection 8.3, Corollary 9.15 gives a lower bound on the set $\mathrm{T}\liminf_{n}\overline{\Sigma_{k}(\pi_{n})}$ for each $k$. This can be an equality: for example, this is so if $\mathfrak{A}=C^{\ast}\Gamma$, the states $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}$ converge to the regular tracial state, and $(\pi_{n})_{n\geq 1}$ also converges strongly to the regular representation. Some important examples satisfying these three conditions are surveyed in [113]. But in general the inclusion can be strict. Nevertheless, we do always obtain the following.
Corollary 9.16.
Assume that $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}\to\tau$. Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, and consider its Lebesgue decomposition relative to $\tau$ as given by Proposition 4.9. Then $\varphi$ is asymptotically associated to $\boldsymbol{\pi}$ if and only if $\varphi_{\mathrm{sing}}$ is asymptotically associated to $\boldsymbol{\pi}$.
Proof.
First, by Proposition 4.9 and Lemmas 8.11 and 8.15, it suffices to prove that $\langle\varphi,\cdot\rangle$ is asymptotically associated to $\boldsymbol{\pi}^{(k)}$ if and only if $\langle\varphi_{\mathrm{sing}},\cdot\rangle$ is asymptotically associated to $\boldsymbol{\pi}^{(k)}$. After adjusting some notation if necessary, this reduces our work to the case when $k=1$. Assume this for the rest of the proof.
By Proposition 3.2, $\varphi_{\mathrm{sing}}$ is associated to $\pi_{\varphi}$. Therefore, if $\varphi$ is asymptotically associated to $\boldsymbol{\pi}$, then $\varphi_{\mathrm{sing}}$ is as well by Corollary 8.5.
On the other hand, Corollary 9.15 tells us that $\varphi_{\mathrm{ac}}$ is always asymptotically associated to $\boldsymbol{\pi}$, and $\varphi_{\mathrm{ac}}$ and $\varphi_{\mathrm{sing}}$ are disjoint by construction. Therefore, if $\varphi_{\mathrm{sing}}$ is asymptotically associated to $\boldsymbol{\pi}$, then so is $\varphi$ itself by Lemma 8.12. ∎
Lemmas 9.12 and 9.14 let us determine the cases of equality in Lemma 9.7.
Proposition 9.17.
Assume that $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}\to\tau$, and let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ with $\varphi(1)$ nonsingular. Then equality holds in Lemma 9.7 if and only if $\varphi=\tau\otimes\varphi(1)$, where this is a tensor-product map as in Example 4.1.
Proof.
Step 1. First assume that $k=1$. For any neighbourhood $U$ of $\tau$, Lemma 9.14 shows that
$\sigma_{2d_{n}-1}\mathcal{X}(\pi_{n},U)\to 1.$
Therefore $\mathrm{h}_{\boldsymbol{\pi}}(\tau)=0$, by Lemma 9.12.
On the other hand, suppose that $\varphi\neq\varphi(1)\cdot\tau$. After normalizing using Proposition 9.10(a), we may assume that $\varphi(1)=1$. Since $\varphi\neq\tau$, they have disjoint neighbourhoods, say $U$ and $U^{\prime}$ respectively. Applying Lemma 9.14 to the neighbourhood $U^{\prime}$, there are positive constants $C$ and $c$ such that
$\sigma_{2d_{n}-1}\mathcal{X}(\pi_{n},U)\leq\sigma_{2d_{n}-1}\big{(}\mathrm{S}^{2d_{n}-1}\setminus\mathcal{X}(\pi_{n},U^{\prime})\big{)}\leq Ce^{-cd_{n}}.$
This turns into $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)\leq-c<0$, again by Lemma 9.12.
Step 2. Now suppose that $k>1$. Then the case $k=1$ and Lemmas 8.15 and 9.8 show that $\mathrm{h}_{\boldsymbol{\pi}}(\tau\otimes I_{k})=0$.
Let $\varphi$ be any other element of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ such that $\varphi(1)$ is non-singular and equality is achieved in Lemma 9.7. Applying Proposition 9.10(a) with $Q:=\varphi(1)^{-1/2}$, we may this time assume that $\varphi(1)=I_{k}$ and $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)=0$.
Having done so, consider the sequence $\boldsymbol{\pi}^{(k)}$ and the pairing functional $\langle\varphi,\cdot\rangle$ on $\mathrm{M}_{k}(\mathfrak{A})$. Since $\varphi$ achieves equality in Lemma 9.7, Lemma 9.8 gives
$\mathrm{h}_{\boldsymbol{\pi}^{(k)}}(\langle\varphi,\cdot\rangle)=\frac{1}{k}\mathrm{h}_{\boldsymbol{\pi}}(\varphi)=0.$
Therefore the case $k=1$ applied to $\boldsymbol{\pi}^{(k)}$ and $\langle\varphi,\cdot\rangle$ shows that
$\langle\varphi,\cdot\rangle=\tau\otimes\overline{\mathrm{tr}}_{k}=\langle\tau\otimes I_{k},\cdot\rangle,$
and hence that $\varphi=\tau\otimes I_{k}$. ∎
Corollary 9.18.
Assume that $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}\to\tau$. Let $\tau$ be associated to $\lambda$ by $\xi$, let $a\in\mathfrak{A}$, and let $\varphi:=\Phi^{\lambda}_{\lambda(a)\xi}$. Then $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)=2\log\Delta_{\tau}a$.
Proof.
The expression for $\varphi$ in terms of $\tau$ and $a$ is a special case of the relationship from Lemma 4.3. Therefore Proposition 9.10(b) gives
$\mathrm{h}_{\boldsymbol{\pi}}(\varphi)=2\log\Delta_{\tau}a+\mathrm{h}_{\boldsymbol{\pi}}(\tau).$
On the right-hand side, the second term vanishes by Proposition 9.17. ∎
Corollary 9.18 is a precursor to Theorem C. In the next subsection we use this corollary in the full proof of that theorem. This is somewhat similar to the use of Proposition 6.6 in the proof of Theorem A.
9.5 Proof of Theorem C
Let $\tau$ be associated to the representation $\lambda$ by the cyclic tracial vector $\xi$, and let $\widetilde{\tau}$ be the normal tracial state on $\lambda(\mathfrak{A})^{\prime}$ defined by $\xi$ as in (3.5). We use the letter $\Delta$ for both (i) the Fuglede–Kadison determinant on $\mathfrak{A}$ defined by $\tau$ and also (ii) the Fuglede–Kadison determinant on log-integrable operators affiliated to $\lambda(\mathfrak{A})^{\prime}$ defined by $\widetilde{\tau}$.
Proof of Theorem C.
We break this into four steps.
Step 1. First we prove the inequality $\geq$” when $k=1$ and $\varphi$ is $\lambda$-normal. In this case, Proposition 3.5 associates $\varphi$ to $\lambda$ by a vector of the form $T\xi$ for some positive semi-definite operator $T$ affiliated to $\lambda(\mathfrak{A})^{\prime}$, and Definition 3.7 gives $\Delta\varphi=(\Delta T)^{2}$. Now Corollary 3.8 gives a sequence of invertible positive elements $a_{1}$, $a_{2}$, …of $\mathfrak{A}$ such that
$\lambda(a_{i})\xi\to T\xi\ \qquad\hbox{and}\qquad\Delta a_{i}\to\Delta T\qquad\hbox{as}\ i\to\infty.$
Let $\varphi_{i}:=\Phi^{\lambda}_{\lambda(a_{i})\xi}$ for each $i$. Then Lemma 4.2 gives $\varphi_{i}\to\varphi$ in $\mathfrak{A}^{\ast}$, and Corollary 9.18 gives $\mathrm{h}_{\boldsymbol{\pi}}(\varphi_{i})=2\log\Delta a_{i}$ for each $i$. Letting $i\to\infty$, the upper semicontinuity from Lemma 9.6 turns this into
$\mathrm{h}_{\boldsymbol{\pi}}(\varphi)\geq 2\lim_{i\to\infty}\log\Delta a_{i}=2\log\Delta T=\log\Delta\varphi.$
Step 2. Next we prove the inequality $\geq$” when $k=1$ but $\varphi$ is otherwise arbitrary. Let $O$ be any neighbourhood of $\varphi$. Since $\varphi_{\mathrm{ac}}$ and $\varphi_{\mathrm{sing}}$ are disjoint, Corollary 8.8(a) gives neighbourhoods $U$ of $\varphi_{\mathrm{sing}}$ and $W$ of $\varphi_{\mathrm{ac}}$ such that
$\mathcal{X}(\pi_{n},O)\supset\mathcal{X}(\pi_{n},U)+\mathcal{X}(\pi_{n},W)$
for every $n$. Since $\varphi$ is asymptotically associated to $\boldsymbol{\pi}$ by assumption, so is $\varphi_{\mathrm{sing}}$ by Corollary 9.16. Therefore $\mathcal{X}(\pi_{n},U)$ is nonempty along an infinite subsequence of values of $n$, say $n_{1}<n_{2}<\dots$. Let $\boldsymbol{\pi}^{\prime}$ be the corresponding AP subsequence of $\boldsymbol{\pi}$. For each $n_{i}$, the set $\mathcal{X}(\pi_{n_{i}},O)$ contains a translate of $\mathcal{X}(\pi_{n_{i}},W)$, and so
$\limsup_{n\to\infty}\frac{1}{d_{n}}\log\frac{\mathrm{vol}_{2d_{n}}\mathcal{X}(\pi_{n},O)}{v(d_{n})}$ $\geq\limsup_{i\to\infty}\frac{1}{d_{n_{i}}}\log\frac{\mathrm{vol}_{2d_{n_{i}}}\mathcal{X}(\pi_{n_{i}},O)}{v(d_{n_{i}})}$
$\geq\limsup_{i\to\infty}\frac{1}{d_{n_{i}}}\log\frac{\mathrm{vol}_{2d_{n_{i}}}\mathcal{X}(\pi_{n_{i}},W)}{v(d_{n_{i}})}$
$\geq\mathrm{h}_{\boldsymbol{\pi}^{\prime}}(\varphi_{\mathrm{ac}}).$
This lower bound is at least $\log\Delta\varphi_{\mathrm{ac}}$ by applying Step 1 along the AP sequence $\boldsymbol{\pi}^{\prime}$. Since $O$ is arbitrary, this proves that $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ is also at least $\log\Delta\varphi_{\mathrm{ac}}$.
Step 3. We now prove the inequality $\leq$” in case $k=1$. This proof is quickest via the variational principle from Proposition 3.9.
Let $a\in\mathfrak{A}$ be positive and invertible and satisfy $\Delta a\geq 1$. Define a new positive functional by $\psi:=\varphi(\sqrt{a}(\cdot)\sqrt{a})$. Then we have
$\log\varphi(a)$ $=\log\psi(1)$
$\geq\mathrm{h}_{\boldsymbol{\pi}}(\psi)$ (Lemma 9.7)
$=2\log\Delta\sqrt{a}+\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ (Proposition 9.10(b))
[…]
Taking the infimum over $a$, Proposition 3.9 turns this into $\log\Delta\varphi_{\mathrm{ac}}\geq\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$.
Step 4. Finally, if $k>1$, then we can apply the previous steps to the functional $\langle\varphi,\cdot\rangle$ on $\mathrm{M}_{k}(\mathfrak{A})$. First, Lemma 8.15 gives that $\overline{\mathrm{tr}}_{kd_{n}}\circ\pi_{n}^{(k)}\to\tau\otimes\overline{\mathrm{tr}}_{k}$. Secondly, Lemma 8.11 gives that $\langle\varphi,\cdot\rangle$ is asymptotically associated to $\boldsymbol{\pi}^{(k)}$. Finally, we have
$\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ $=k\cdot\mathrm{h}_{\boldsymbol{\pi}^{(k)}}(\langle\varphi,\cdot\rangle)$ (Lemma 9.8)
$=k\cdot\log\Delta_{\tau\otimes\overline{\mathrm{tr}}_{k}}(\langle\varphi,\cdot\rangle_{\mathrm{ac}})$ (case $k=1$ of Theorem C)
$=k\cdot\log\Delta_{\tau\otimes\overline{\mathrm{tr}}_{k}}(\langle\varphi_{\mathrm{ac}},\cdot\rangle)$ (Proposition 4.9)
[…]
9.6 Some consequences of Theorem C
Corollary 9.19.
Assume that $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}\to\tau$, and let $\varphi,\psi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$.
  1. a.
    We have $\mathrm{h}_{\boldsymbol{\pi}}(\varphi_{\mathrm{ac}})=\log\Delta\varphi_{\mathrm{ac}}$ and
    $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)=\left\{\begin{array}[]{ll}\mathrm{h}_{\boldsymbol{\pi}}(\varphi_{\mathrm{ac}})&\quad\hbox{if $\varphi_{\mathrm{sing}}$ is asymptotically associated to $\boldsymbol{\pi}$}\\ -\infty&\quad\hbox{otherwise.}\end{array}\right.$
  2. b.
    If $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)>-\infty$, then $\varphi_{\mathrm{ac}}$ is associated to $\lambda^{\oplus k}$ by a $k$-tuple that is cyclic and separating for $\lambda^{\oplus k}(\mathfrak{A})^{\prime\prime}$, and so
    $\pi_{\varphi}\gtrsim\pi_{\varphi_{\mathrm{ac}}}\simeq\lambda^{\oplus k}.$
  3. c.
    If $\varphi\geq\psi$ in the positive definite ordering, and if $\varphi$ is asymptotically associated to $\boldsymbol{\pi}$, then $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)\geq\mathrm{h}_{\boldsymbol{\pi}}(\psi)$.
Proof.
Part (a) follows directly from Corollary 9.16 and Theorem C.
If $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)>-\infty$, then Theorem C tells us that $\Delta\varphi_{\mathrm{ac}}>0$. Expressing this Fuglede–Kadison determinant in terms of the affiliated operator $T$ from Proposition 4.10, it follows that $\Delta T$ is also positive, and hence $T$ is nonsingular. Since $T$ is self-adjoint, it therefore also has dense image. Since $\pi_{\varphi_{\mathrm{ac}}}$ is equivalent to the subrepresentation of $\lambda^{\oplus k}$ defined by $\overline{\mathrm{img}\,T}$, this proves part (b).
Finally, if $\varphi\geq\psi$ and $\varphi$ is asymptotically associated to $\boldsymbol{\pi}$, then so is $\psi$, and Theorem C shows that part (c) is implied by the inequality $\Delta\varphi_{\mathrm{ac}}\geq\Delta\psi_{\mathrm{ac}}$ from Corollary 3.11(a). ∎
Under the conditions of Theorem C, AP entropy along $\boldsymbol{\pi}$ is really just the logarithm of a Fuglede–Kadison determinant, provided one knows which positive definite functions are asymptotically associated to $\boldsymbol{\pi}$. By passing to a subsequence, one can always assume that $\boldsymbol{\pi}$ strong-quotient converges, and then Theorem C shows that the AP entropy function $\mathrm{h}_{\boldsymbol{\pi}}$ depends only on the values of the limits
$\tau=\lim_{n}\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}\qquad\hbox{and}\qquad\lim_{n}\overline{\Sigma_{k}(\pi_{n})}\qquad\hbox{for}\ k=1,2,\dots.$
If $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}\to\tau$ but $\boldsymbol{\pi}$ does not strong-quotient converge, then Theorem C may give different values of AP entropy along different subsequences, but the only two possible values are $\log\Delta_{\tau}\varphi_{\mathrm{ac}}$ or $\infty$- ∞ (which may still happen to be equal).
In ergodic theory, an important open problem asks whether the sofic entropy of a measure-preserving system along any sofic approximation must always equal either one particular value or $\infty$- ∞. This would mean that the sofic entropy of a measure-preserving system is ‘unambiguous’ in that it does not depend on the choice of sofic approximation provided it is finite. Corollary 9.19 answers the analogous question positively for AP entropy. On the other hand, it is known that the answer is negative for the topological variant of sofic entropy [2]. If the answer is negative for sofic entropy itself, then any general method for evaluating or estimating sofic entropy must depend on the particular sofic approximation through more than just its set of asymptotically associated stationary processes.
Remark 9.20.
[…]
Assume again that $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}\to\tau$. We can reverse the discussion above by asking whether the function $\mathrm{h}_{\boldsymbol{\pi}}$ determines the topological upper limit of the sequence $(\Sigma_{k}(\pi_{n}))_{n\geq 1}$ for each $k$. This is slightly subtle, because for $\varphi\in\Sigma_{k}(\mathfrak{A})$ the value $\log\Delta_{\tau}\varphi$ may equal $\infty$- ∞ even if $\varphi$ is $\lambda$-normal. We can evade this issue by instead using $\varphi$ to form the perturbations
$\varphi_{t}:=\tau\otimes I_{k}+t\varphi\qquad(t\geq 0).$
These give us a way to ‘mollify’ $\varphi$. We can control their entropies as follows.
Corollary 9.21.
Under the assumptions above, the following hold.
  1. a.
    If $\varphi_{t}$ is asymptotically associated to $\boldsymbol{\pi}$ for some $t>0$, then this holds for all $t>0$, and we have
    $0\leq\mathrm{h}_{\boldsymbol{\pi}}(\varphi_{t})\leq k\cdot\log(1+t\cdot\overline{\mathrm{tr}}_{k}\varphi(1))\qquad(t\geq 0).$
  2. b.
    If $\varphi_{t}$ is not asymptotically associated to $\boldsymbol{\pi}$ for any $t>0$, then $\mathrm{h}_{\boldsymbol{\pi}}(\varphi_{t})=-\infty$ for every $t>0$.
Proof.
The linearity of the Lebesgue decomposition gives
$(\varphi_{t})_{\mathrm{ac}}=\tau\otimes I_{k}+t\varphi_{\mathrm{ac}}\qquad\hbox{and}\qquad(\varphi_{t})_{\mathrm{sing}}=t\varphi_{\mathrm{sing}}$
for every $t\geq 0$. Therefore, if $\varphi_{t}$ is asymptotically associated to $\boldsymbol{\pi}$ for some $t>0$, then so is $\varphi_{\mathrm{sing}}$, and then so is $\varphi_{s}$ for every other $s$, by two applications of Corollary 9.16.
Now the lower bound on $\mathrm{h}_{\boldsymbol{\pi}}(\varphi_{t})$ in part (a) follows from Corollary 9.19(c) and Proposition 9.17, because $\varphi_{t}\geq\tau\otimes I_{k}$ in the positive definite ordering. On the other hand, since $\Delta_{\tau\otimes\overline{\mathrm{tr}}_{k}}(1\otimes I_{k})=1$, the element $1\otimes I_{k}$ of $\mathrm{M}_{k}(\mathfrak{A})$ is allowed inside the infimum of the variational principle from Proposition 4.13. Combined with Theorem C, this gives
$\mathrm{h}_{\boldsymbol{\pi}}(\varphi_{t})=\log\Delta_{\tau}(\varphi_{t})_{\mathrm{ac}}\leq k\cdot\log\langle\varphi_{t},1\otimes I_{k}\rangle.$
This gives the desired upper bound on $\mathrm{h}_{\boldsymbol{\pi}}(\varphi_{t})$, because
$\langle\varphi_{t},1\otimes I_{k}\rangle=\langle\tau\otimes I_{k},1\otimes I_{k}\rangle+t\langle\varphi,1\otimes I_{k}\rangle=1+t\cdot\overline{\mathrm{tr}}_{k}\varphi(1).$
Finally, part (b) follows from Definition 9.4, applied when the sets $\mathcal{X}(\pi_{n},O)$ are empty for some sufficiently small neighbourhood $O$ of $\varphi_{t}$. ∎
Remark 9.22.
[…]
Now we can see how the function $\mathrm{h}_{\boldsymbol{\pi}}$ determines topological upper limits.
Corollary 9.23.
Assume that $\overline{\mathrm{tr}}_{d_{n}}\circ\pi\to\tau$. Then
$\mathrm{T}\limsup_{n\to\infty}\Sigma_{k}(\pi_{n})$ =$=\{\varphi\in\Sigma_{k}(\mathfrak{A}):\ \mathrm{h}_{\boldsymbol{\pi}}(\varphi_{t})\geq 0\ \forall t\geq 0\}$
=$=\{\varphi\in\Sigma_{k}(\mathfrak{A}):\ \mathrm{h}_{\boldsymbol{\pi}}(\varphi_{t})\to 0\ \hbox{as}\ t\downarrow 0\}$
for each positive integer $k$, and so $\boldsymbol{\pi}$ strong-quotient converges if and only if we have $\mathrm{h}_{\boldsymbol{\pi}^{\prime}}=\mathrm{h}_{\boldsymbol{\pi}}$ whenever $\boldsymbol{\pi}^{\prime}$ is a subsequence of $\boldsymbol{\pi}$. ∎
In ergodic theory, Kolmogorov–Sinai entropy for single transformations is additive under Cartesian products [157, Theorem 4.23]. By contrast, sofic entropy is subadditive for joinings, but it may fail to be additive, even for a product joining [11]. The same phenomena hold for AP entropy, for essentially the same reasons. Let us assume again that that $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}\to\tau$ and that $\boldsymbol{\pi}$ strong-quotient converges; if these conditions fail then we can always pass to a subsequence that satisfies them.
Corollary 9.24.
Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, let $\psi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{\ell})_{+}$, and let $\theta$ be a joining of them (recall Definition 4.4). Then
$\mathrm{h}_{\boldsymbol{\pi}}(\theta)\leq\mathrm{h}_{\boldsymbol{\pi}}(\varphi)+\mathrm{h}_{\boldsymbol{\pi}}(\psi).$ (9.10)
If $\theta_{\mathrm{ac}}=\mathrm{diag}(\varphi_{\mathrm{ac}},\psi_{\mathrm{ac}})$, then equality holds provided $\theta$ is asymptotically associated to $\boldsymbol{\pi}$.
Proof.
If $\theta$ is not asymptotically associated to $\boldsymbol{\pi}$, then $\mathrm{h}_{\boldsymbol{\pi}}(\theta)=-\infty$ and there is nothing else to prove. So assume that $\theta$ is asymptotically associated to $\boldsymbol{\pi}$. Both $\varphi$ and $\psi$ are associated to $\pi_{\theta}$, so these are also asymptotically associated to $\boldsymbol{\pi}$. Now the AP entropies of all three maps are given by Theorem C, and the desired conclusions follow from subadditivity and the case of equality for $\log\Delta$: see Corollary 4.14. ∎
Remark 9.25.
Alternatively, the subadditivity of $\mathrm{h}_{\boldsymbol{\pi}}$ may be proved by observing that, for any neighbourhoods $U$ of $\varphi$ and $V$ of $\psi$, there is a neighbourhood $O$ of $\theta$ such that
$\mathcal{X}(\pi,O)\subset\mathcal{X}(\pi,U)\times\mathcal{X}(\pi,V)$
[…]
It is not enough to assume that $\varphi$ and $\psi$ are both separately asymptotically associated to $\boldsymbol{\pi}$ in the last part of Corollary 9.24; see Example 9.27 below.
Notions of entropy are often given by concave functions, so it is natural to ask whether the restrictions $\mathrm{h}_{\boldsymbol{\pi}}|\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ have this property. In a restricted sense, they do. We obtain the following from Theorem C together with the log-concavity inequality (3.4) (a consequence of Corollary 3.11).
Corollary 9.26.
Let $\varphi,\psi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, and assume moreover that ${t\varphi+(1-t)\psi}$ is asymptotically associated to $\boldsymbol{\pi}$ for every $t\in[0,1]$. Then $\mathrm{h}_{\boldsymbol{\pi}}$ is concave along the line segment from $\varphi$ to $\psi$:
$\mathrm{h}_{\boldsymbol{\pi}}(t\varphi+(1-t)\psi)\geq t\mathrm{h}_{\boldsymbol{\pi}}(\varphi)+(1-t)\mathrm{h}_{\boldsymbol{\pi}}(\psi)\qquad(0\leq t\leq 1).$
However, the function $\mathrm{h}_{\boldsymbol{\pi}}$ need not be globally concave on $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, or even on its subset $\Sigma_{k}(\mathfrak{A})$. The problem is that the set $\lim_{n}\overline{\Sigma_{k}(\pi_{n})}$ itself need not be convex. We illustrate this by an extension of Example 8.2.
Example 9.27.
Let $\Gamma$ be a countable group with left regular representation $\lambda$. Let $\pi$ be a finite-dimensional irreducible representation of $\Gamma$ that (i) has dimension at least $2$, and (ii) is not approximately contained in $\lambda$. Separately, let $(\rho_{n})_{n\geq 1}$ be an AP sequence for $\Gamma$ that strong-quotient converges to $\lambda$. Suitable examples include randomly generated unitary or permutation representations of free groups, as proved in [35] and [21], respectively; see also [113].
Now let $\pi_{n}:=\pi\oplus\rho_{n}$ for each $n$. Checking the definitions shows that
[…]
This is closed for every $k$ and $n$ because it is a continuous image of a finite-dimensional sphere. One can now check that $(\pi_{n})_{n\geq 1}$ strong-quotient converges to $\pi\oplus\lambda$, and so in particular
[…]
in the Vietoris topology. However, since $\pi$ is irreducible and not approximately associated to $\lambda$, it follows that $\lim_{n}\Sigma_{2}(\pi_{n})$ does not contain $\mathrm{diag}(\varphi,\psi)$ when $\varphi$ and $\psi$ are linearly independent elements of $\Sigma_{1}(\pi)$. Moreover, by the same reasoning as for Example 8.2, it also follows that $\lim_{n}\Sigma_{1}(\pi_{n})$ does not contain $(\varphi+\psi)/2$, so $\lim_{n}\Sigma_{1}(\pi_{n})$ is not convex. Finally, Corollary 9.21(a) gives that $\mathrm{h}_{\boldsymbol{\pi}}(\varphi+\tau)$ and $\mathrm{h}_{\boldsymbol{\pi}}(\psi+\tau)$ are both non-negative, but we have
$\mathrm{h}_{\boldsymbol{\pi}}(\mathrm{diag}(\varphi+\tau,\psi+\tau))=\mathrm{h}_{\boldsymbol{\pi}}((\varphi+\tau)/2+(\psi+\tau)/2)=-\infty.$
As suggested by Example 9.27, one can show in general that the parts of $\lim_{n}\Sigma_{k}(\pi_{n})$ that lie outside the lower bound on this set provided by Corollary 9.15 are responsible for the failure of convexity.
Nevertheless, if $(\pi_{n})_{n\geq 1}$ strong-quotient converges, then $\lim_{n}\Sigma_{k}(\pi_{n})$ is at least star-shaped around the element $\tau\otimes I_{k}$ (or any other $\lambda$-normal elements of $\Sigma_{k}(\mathfrak{A})$). This is a consequence of Corollary 9.16, because $\tau\otimes I_{k}$ does not contribute to the singular part in a convex combination.
Since $\mathrm{h}_{\boldsymbol{\pi}}$ may not be concave on the whole of $\Sigma_{k}(\mathfrak{A})$, it may not be exactly recoverable from its Legendre transform. Nevertheless, it might be interesting to investigate how various features of the sequence $(\pi_{n})_{n\geq 1}$ are reflected by that transform, and how it compares with the Legendre transform of $\log\Delta\varphi_{\mathrm{ac}}$ itself. See [135, Section I.6] for a general account of infinite-dimensional Legendre transforms with a view towards statistical mechanics, or [34] for the resulting variational principle in the case of sofic entropy. In the analogous setting of modes of convergence for graphs of uniformly bounded degree (see Remarks 8.13 and 9.20), convergence of these Legendre transforms would correspond to ‘right convergence’ [20, 23].
9.7 Further remarks
Comparison with previous work
Just like Theorem A, Theorem C has a number of predecessors in the literature. Some of the first that lie beyond the discussion in Subsection 6.3 are Lyons’ calculations in [110, 111]. These concern the problem of asymptotically counting spanning trees along sequences of finite connected graphs using their random weak limits. Lyons shows that this ‘tree entropy’ of the random limit graph is given by a Fuglede–Kadison determinant of its Laplacian. His setting does not require a group action, but it yields results for sofic groups as a special case.
More recent examples are continuations of Deninger’s work on determinantal formulas for the entropies of certain algebraically defined measure-preserving systems, but now for sofic groups and sofic entropy. The first of these was due to Kerr and Li [95, Theorem 7.1]. This theme was then taken up by Hayes. In [80], he generalized Deninger’s conjecture to equate the sofic entropy of $\boldsymbol{X}_{f}$ with the Fuglede–Kadison determinant of $f\in\mathbf{Z}[\Gamma]$ whenever $\Gamma$ is sofic. His main result actually allows larger finite matrices over $\mathbf{Z}[\Gamma]$ in place of $f$. In this respect it approaches a generalization of the Li–Thom theorem from [104] to sofic groups, but Hayes also shows that the full generalization of the Li–Thom theorem is false. See also the alternative proof with further refinements in [84].
Alongside those papers, Hayes has developed other connections between sofic entropy and representation theory. In [82] he proved that an arbitrary measure-preserving $\Gamma$-system can have completely positive sofic entropy only if its Koopman representation is contained in the infinite inflation of the regular representation. In [81], he computed the sofic entropy of a stationary Gaussian process over $\Gamma$ in terms of the real orthogonal representation that defines its first chaos, generalizing one of the results from [78] for single transformations. The main theorem in [81] is worth comparing with the way in which $\varphi_{\mathrm{sing}}$ and $\varphi_{\mathrm{ac}}$ determine whether $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ equals $\infty$- ∞ in Corollary 9.19 above.
Other aspects of Szegő’s theorem
The setting of Theorem C has taken us quite far from our original motivation in the form of Szegő’s theorem. AP entropy is not defined as a limit of finite-dimensional determinants, and related data such as Verblunsky coefficients have no obvious meaning in this generality.
However, some of those finer aspects of Szegő’s theorem make a return in a sequel to the present paper [10]. That paper studies random AP sequences, and in particular an annealed version of AP entropy that can be defined using these. In the special case of uniformly random finite-dimensional representations of free groups, this is a representation theoretic analog of Bowen’s annealed sofic entropy (formerly called the ‘f-invariant’) from [27, 24]. Like annealed sofic entropy, this instance of annealed AP entropy admits a precise formula.
In studying this entropy and its formula, many features of the theory of orthogonal polynomials on $\mathbf{T}$ reappear. For example, sequences of ‘generalized Verblunsky coefficients’ can be used to parametrize positive definite functions over free groups. Rather than requiring a total ordering of the group, these generalized Verblunksy coefficients depend on the fact that free groups have tree-like Cayley graphs. One of the first main theorems about annealed AP entropy is a formula for it as an infinite series in terms of those coefficients. This can be regarded as a much more complete analog of Szegő’s limit theorem for positive definite functions on free groups, with one entirely new feature: an additional term called ‘zeroth-order’ entropy that reflects the non-amenability of the groups.
We leave further details to [10], which also develops applications to large deviations for tuples of random matrices.